### v1

# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

Abstract
We study the covariance model belonging to the Gauss hypergeometric ($\mathcal{GH}$) class, a highly flexible and compactly supported correlation model (kernel). This class includes the well-known Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) kernels as special cases. First, we provide necessary and sufficient conditions for the validity of the $\mathcal{GH}$ model. We then demonstrate that this family of models suffers from identifiability issues under both increasing and fixed-domain asymptotics.
To address this problem, we propose a parsimonious version of the model that adheres to an optimality criterion. This approach results in a new class of compactly supported kernels, which can be seen as an improvement over the $\mathcal{GW}$ model. Additionally, we show that two specific compact support reparameterizations allow us to recover the $\mathcal{MT}$ model, highlighting the advantages and disadvantages of each reparameterization.
[…]
Keywords: Fixed domain asymptotics, Gaussian random fields, Generalized Wendland model, Integral range, Sparse matrices.
1 Introduction
Gaussian processes, or Gaussian random fields (RFs), are widely used in various disciplines such as spatial statistics (Cressie,, 1993; Stein,, 1999; Banerjee et al.,, 2004), machine learning (Rasmussen and Williams,, 2006), numerical analysis (Wendland,, 2004), and computer experiments (Santner et al.,, 2003), among others. The Gaussian assumption implies that the finite-dimensional distributions of the RF are fully determined by the mean value and the covariance function. In particular, covariance functions (also known as kernels in the machine learning literature) are positive semidefinite functions that play a fundamental role in modeling relationships between data points in Gaussian RFs, kernel methods, and other machine learning algorithms.
For RFs defined in Euclidean spaces, covariance functions that depend solely on the distance between any pair of points within the domain are called isotropic or radially symmetric. This paper operates under this assumption, which greatly simplifies the discussion. However, isotropic covariances serve as the building blocks for more complex models, such as anisotropic, nonstationary (Paciorek and Schervish,, 2006), multivariate (Gneiting et al.,, 2010), or spatiotemporal extensions (Porcu et al.,, 2020), among others.
To date, there exists a rich catalog of parametric isotropic covariance functions (see, for instance, Rasmussen and Williams, (2006) or Chilès and Delfiner, (2012)), among which the Matérn ($\mathcal{MT}$ hereafter) model is by far the most widely used (see Porcu et al., (2024) for a comprehensive review). It has played a central role in spatial statistics for decades and, more recently, in other disciplines such as numerical analysis, approximation theory, and machine learning. A key advantage of the $\mathcal{MT}$ model is its continuous parameterization of the differentiability of the sample paths of the associated Gaussian RF through a positive smoothness parameter. The larger the smoothness parameter, the higher the level of differentiability of the sample paths. .
While the $\mathcal{MT}$ model remains widely used, particularly in the stochastic partial differential equations approach and its generalizations to non-Euclidean domains (Lindgren et al.,, 2011; Bolin et al.,, 2024), enhancements to the $\mathcal{MT}$ model have been proposed in recent years. These covariance models retain the benefits of the Matérn model while offering additional features (Porcu et al.,, 2024).
Among these, the generalized Wendland ($\mathcal{GW}$ hereafter) (Gneiting,, 2002; Bevilacqua et al.,, 2019) has gained popularity, as it allows for a continuous parameterization of the differentiability of the sample paths of the underlying Gaussian RF, similar to the $\mathcal{MT}$ model, while also being compactly supported. This is a desirable feature from a computational perspective, as it produces sparse covariance matrices (Duff et al.,, 1989; Davis,, 2006).
Several connections between the $\mathcal{MT}$ and $\mathcal{GW}$ models have been established. For instance, Bevilacqua et al., (2019) show that, under specific conditions, these models lead to equivalent Gaussian measures. Additionally, Bevilacqua et al., (2022) demonstrate that the $\mathcal{MT}$ model is a special limit case of a compactly supported reparameterization of the $\mathcal{GW}$ model. Estimation and prediction using $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied under both increasing and fixed-domain asymptotics (Mardia and Marshall,, 1984; Zhang,, 2004; Kaufman and Shaby,, 2013; Bevilacqua et al.,, 2019).
The focus of this paper is the study of the Gauss hypergeometric ($\mathcal{GH}$) covariance, a highly flexible kernel recently proposed by Emery and Alegría, (2022). These authors demonstrate that the $\mathcal{GH}$ model encompasses most well-known covariance models as special or asymptotic cases, including both the $\mathcal{MT}$ and $\mathcal{GW}$ models. In addition, they provide sufficient conditions for the validity of the model. In Porcu et al., (2024), the Gauss hypergeometric ($\mathcal{GH}$) kernel appears to be one of the most promising generalizations of the $\mathcal{MT}$ model. However, the estimation of the $\mathcal{GH}$ covariance model, particularly under fixed-domain asymptotics, has not yet been addressed.
This paper presents several contributions related to the $\mathcal{GH}$ covariance model. First, we provide necessary and sufficient conditions for the validity of the $\mathcal{GH}$ model, offering a complete characterization of the kernel. This result ensures that the model is well-defined and applicable to a wide range of scenarios.
Second, we show that this family of models suffers from identifiability issues under both increasing and fixed-domain asymptotics. These findings highlight a critical limitation of the $\mathcal{GH}$ model in practical applications, particularly when attempting to infer parameters from data. To address this problem, we propose a parsimonious version of the model that adheres to an optimality criterion. This criterion is based on the maximization of the isotropic integral range (Lantuéjoul,, 2002), which can be seen as a measure of the flexibility of the correlation model. For instance, if $d=1$, the maximum integral range corresponds to the maximum area under the unidimensional correlation function.
Using this approach, we derive a new class of compactly supported kernels, which we call the hypergeometric ($\mathcal{H}$ hereafter) model that can be seen as an improvement over the $\mathcal{GW}$ model. The model can also be seen as a generalization of the Euclid’s hat model proposed in Gneiting, (1999), which includes the celebrated spherical models widely used in geostatistics (Chilès and Delfiner,, 2012). Additionally, we show that two specific compact support reparameterizations allow us to recover the $\mathcal{MT}$ model as a special limit case, highlighting the advantages and disadvantages of each reparameterization.
Finally, for a fixed smoothness parameter, we establish strong consistency and the asymptotic distribution of the maximum likelihood (ML) estimator of the microergodic parameter associated with the $\mathcal{H}$ class.
[…]
The $\mathcal{H}$ model has been implemented in the GeoModels package (Bevilacqua et al.,, 2025) for the open-source R statistical environment.
The remainder of this paper is organized as follows. Section 2 provides background material on the $\mathcal{MT}$ and $\mathcal{GW}$ covariance models. Section 3 introduces the $\mathcal{GH}$ model and highlights the identifiability issue. Section 4 first provides necessary and sufficient conditions for the $\mathcal{GH}$ model. Then, following the criterion of maximal integral range, the $\mathcal{H}$ model is proposed. In Section 5, some closed-form expressions for the $\mathcal{H}$ model that do not depend on the computation of special functions are presented. In addition, two compactly supported reparameterizations that include the $\mathcal{MT}$ model as a special case are introduced. In Section 6, we provide the asymptotic properties of the ML estimator of the microergodic parameter. In Section 7, we report a simulation study that explores the finite-sample properties of the ML method when estimating the parameters of the proposed (reparameterized) $\mathcal{H}$ model. In Section 8, we present the analysis of a climate data set. Concluding remarks are provided in Section 9. All proofs have been deferred to Appendix A.
2 Background: the Matérn and Generalized Wendland correlation models
We denote $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ a zero-mean Gaussian RF on a bounded set $D$ of $\mathbb{R}^{d}$, with stationary isotropic correlation function $C:\mathbb{R}^{d}\to\mathbb{R}$. We consider the class $\Phi_{d}$ of continuous mappings $\phi:[0,\infty)\to\mathbb{R}$ with $\sigma^{2}:=\phi(0)=1$, such that
${\rm corr}\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),$
[…]
Schoenberg, (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and non-negative measure, $F$:
$\phi(x)=\int_{0}^{\infty}\Omega_{d}(xr)F(\text{d}r),\qquad x\geq 0,$
with $\Omega_{d}(x)=\Gamma(d/2)(2/x)^{d/2-1}J_{d/2-1}(x)$ and $J_{\nu}$ the Bessel function of the first kind of order $\nu$. The class $\Phi_{d}$ is nested, with the inclusion relation $\Phi_{1}\supset\Phi_{2}\supset\ldots\supset\Phi_{\infty}$ being strict, and where $\Phi_{\infty}:=\bigcap_{d\geq 1}\Phi_{d}$ is the class of continuous mappings $\phi$, the radial version of which is positive semidefinite on any $d$-dimensional Euclidean space.
Fourier transforms of radial versions of members of $\Phi_{d}$, for a given $d$, have a simple expression, as reported in Yaglom, (1987) or Stein, (1999). For a member $\phi$ of the class $\Phi_{d}$ such that $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$, we define its isotropic spectral density as
$\widehat{\phi}(z)=\frac{z^{1-d/2}}{(2\pi)^{\frac{d}{2}}}\int_{0}^{\infty}u^{d/2}J_{d/2-1}(uz)\phi(u){\rm d}u,\qquad z\geq 0.$ (1)
In the following we review two important examples:
  1. 1.
    The Matérn (${\cal MT}$) family
    The Matérn model is a two-parameter globally supported correlation function defined as (Matérn,, 1986):
    ${\cal MT}_{\nu,\alpha}(x)=\begin{cases}1,\quad x=0\\ \frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,\end{cases}$ (2)
    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun,, 1970, 9.6.22). The associated spectral density, $\widehat{{\cal MT}}_{\alpha,\nu}$, is given by
    $\widehat{{\cal MT}}_{\nu,\alpha}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{\alpha^{d}}{(1+\alpha^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
    Arguments in Stein, (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
    The importance of the Matérn model lies in the parameter $\nu$, which controls the differentiability (in the mean square sense) of the associated Gaussian RF. Specifically, for any integer $l=0,1,\ldots$, a Gaussian RF with a Matérn covariance function is $l$-times mean square differentiable, in any direction, if and only if $\nu>l$. When $\nu=l+1/2$, the Matérn covariance function simplifies into the product of an exponential covariance and a polynomial of order $l$, given by:
    ${\cal MT}_{l+1/2,\alpha}(x)=\exp\left(-\frac{x}{\alpha}\right)\sum_{i=0}^{l}\frac{(l+i)!}{2l!}\binom{l}{i}\left(\frac{2x}{\alpha}\right)^{l-i},\quad l=0,1,\ldots$
  2. 2.
    The Generalized Wendland (${\cal GW}$) family
    The ${\cal GW}$ model is a three-parameters compactly supported correlation function defined as (Gneiting,, 2002; Bevilacqua et al.,, 2019):
    $\begin{split}{\cal GW}_{\kappa,\mu,\beta}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{\beta^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{\beta^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and ${}_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun,, 1970, 15.1.1). The associated spectral density is given by:
    $\widehat{{\cal GW}}_{\kappa,\mu,\beta}(z)=L\beta^{d}\mathstrut_{1}F_{2}\Big{(}\lambda;\lambda+\frac{\mu}{2},\lambda+\frac{\mu}{2}+\frac{1}{2};-\frac{(z\beta)^{2}}{4}\Big{)},\quad z>0,$ (5)
    where ${}_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\nu+1)\Gamma(2\nu+d)\Gamma(\nu)}{\Gamma\left(\nu+{d}/{2}\right)\Gamma(\mu+2\nu+d+1)\Gamma(2\nu)}$ is a normalization constant.
    For a given $\kappa>-\frac{1}{2}$, a necessary and sufficient condition for ${\cal GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ is (Bevilacqua et al.,, 2024)
    $\mu\geq\mu_{\min}(\kappa,d):=\begin{cases}\frac{\sqrt{8\kappa+9}-1}{2}\text{ if $d=1$ and $-\frac{1}{2}<\kappa<0$}\\ \kappa+\frac{d+1}{2}\text{ otherwise.}\end{cases}$ (6)
    The ${\cal GW}_{\kappa,\mu,\beta}$ model, as in the ${\cal MT}_{\nu,\alpha}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$. Specifically, for any integer $l=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,\beta}$ model are $l$ times differentiable, in any direction, if and only if $\kappa>l-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
    When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$:
    ${\cal GW}_{k,\mu,\beta}=\left(1-\frac{x}{\beta}\right)^{\mu+k}_{+}P_{k}(x;\mu,\beta),\quad k=0,1,2,\ldots$ (7)
    where $P_{k}(h,\mu,\beta)$ is a polynomial of degree $k$ that can be expressed as:
    $P_{k}(x;\mu,\beta)=\frac{2^{k}k!}{(2k)!}\sum_{j=0}^{k}c_{j,k}(\mu)\left(\frac{x}{\beta}\right)^{k-j}\left(1-\frac{x}{\beta}\right)^{j},$ (8)
    and $c_{j,k}(\mu)$ are suitable constants (Bevilacqua et al.,, 2024).
    Recently, Bevilacqua et al., (2022) considered a compact support reparameterized version of the ${\cal GW}$ model, and proved that
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(\beta,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (9)
    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $[0,\infty)$.
    A simpler compact support reparameterization that leads to the ${\cal MT}$ model as a special limit case has been proposed in Emery et al., (2025):
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,\mu\beta}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (10)
    again with uniform convergence on $[0,\infty)$.
    As a consequence, the two reparameterized ${\cal GW}$ models can be viewed as a generalization of the ${\cal MT}$ model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
3 The Gauss hypergeometric model and the identifiability problem
The $\mathcal{GH}$ class in Emery and Alegría, (2022) has been defined starting from the following spectral density:
$\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}\mathstrut_{1}F_{2}\Big{(}\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\Big{)},\quad z\geq 0.$ (11)
Note that (11) has basically the same form of the spectral density of the ${\cal GW}$ model in (5). However, in (11) there is no dependence between the parameters of the $\mathstrut{}_{1}F_{2}$ function as in the ${\cal GW}$ case.
Using some results in Cho and Yun, (2018), Emery and Alegría, (2022) show that, for a given $a>0$, the following conditions:
  1. (a)
    $\delta>\frac{d}{2}$;
  2. (b)
    $2(\beta-\delta)(\gamma-\delta)\geq\delta$;
  3. (c)
    $\beta+\gamma\geq 3\delta+\frac{1}{2}$;
are sufficient for the positivity of the spectral density on $[0,\infty)$. The associated four-parameter isotropic correlation model is given by:
$\begin{split}\mathcal{GH}_{\delta,\beta,\gamma,a}(x)&=\frac{\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})}{\Gamma(\beta-\delta+\gamma-\frac{d}{2})\Gamma(\delta-\frac{d}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\beta-\delta+\gamma-\frac{d}{2}-1}\\ &\times{}_{2}F_{1}\left(\beta-\delta;\gamma-\delta;\beta-\delta+\gamma-\frac{d}{2};\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (12)
Here $a>0$ is a compact support parameter, $\delta>\frac{d}{2}$ is a smoothness parameter, and $\beta$ and $\gamma$ are two additional shape parameters.
As in the ${\cal MT}$ and ${\cal GW}$ cases, the ${\cal GH}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\delta$. Furthermore, it is not difficult to see that the ${\cal GH}$ model includes the ${\cal GW}$ and the ${\cal MT}$ models as special cases. For instance,
$\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+\frac{1}{2},a}(x)={\cal GW}_{\kappa,\mu,a}(x),$ (13)
where $\delta=\frac{d+1}{2}+\kappa$.
Although the ${\cal GH}$ model is highly flexible, it suffers from identifiability problems. This is apparent under an increasing domain scenario, where all the parameters can be estimated, since $\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a}$ due to the symmetry of the Gauss hypergeometric function: ${}_{2}F_{1}\left(a;b;\cdot;x\right)={}_{2}F_{1}\left(b;a;\cdot;x\right)$. This implies that there is not a one-to-one correspondence between the parametric space and the correlation function.
To study identifiability under fixed-domain asymptotics, the following theorem establishes the equivalence condition for two distinct Gaussian measures with $\mathcal{GH}$ correlation models that share a common smoothness parameter. Hereafter, we denote by $P(\rho_{i})$, $i=0,1$, a zero-mean Gaussian measure with an isotropic covariance function $\rho_{i}$.
Theorem 1.
Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures. If $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$, for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$, if and only if
$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}},$ (14)
where $L(\beta_{i},\gamma_{i})=\frac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$ with $i=0,1.$
An immediate consequence of Theorem 1 is that the microergodic parameter of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model is given by $\frac{\sigma^{2}L(\beta,\gamma)}{a^{2\delta-d}}$. Since $L(\beta,\gamma)=L(\gamma,\beta)$, it is clear that there is not a unique correspondence between the parametric space and the microergodic parameter, which implies that the microergodic parameter is not identifiable.
Summarizing our results, the $\mathcal{GH}$ correlation model is highly flexible, including the $\mathcal{GW}$ and the $\mathcal{MT}$ models as special cases. However, it is not identifiable under both increasing and fixed-domain scenarios. To fix this issue, in the next section, we will assume that the second shape parameter ($\gamma$) is a function of the first shape parameter ($\beta$).
4 Fixing the identifiability problem: the Hypergeometric correlation model
In this Section we first provide necessary and sufficient conditions for the validity of the ${\cal GH}$ model. Then we solve the identifiability problem by considering a parsimonious version that responds to an optimality criterion. With this goal in mind, we consider a ${\cal GW}$-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model, that is, we set
$\delta=\frac{d+1}{2}+\kappa,\quad\beta=\delta+\frac{\mu}{2},\quad\gamma=\beta+l,$ (15)
with $l\geq 0$. As outlined in (13), this parameterization includes the ${\cal GW}$ model as a special case when $l=\frac{1}{2}$.
The following theorem can be viewed as an improvement of Theorem 1 in Emery and Alegría, (2022) and provides necessary and sufficient conditions for the validity of the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model.
Theorem 2.
Let $d$ be a fixed dimension, $\delta=\frac{d+1}{2}+\kappa$ and $\kappa>-\frac{1}{2}$ a smoothness parameter. Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ if and only if
  1. 1.
    $\mu\geq\delta-l+\frac{1}{2}\quad$ if $\quad 0\leq l\leq\frac{d}{2}+\kappa$
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>\frac{d}{2}+\kappa$.
It should be stressed that necessary and sufficient conditions can also be found for $l<0$. They are symmetric with respect to the case $l\geq 0$ (see the proof of Theorem 2 in Appendix A). However, to solve the issue of identifiability, hereafter we consider the condition $l\geq 0$ as in Theorem 2.
If $l=\frac{1}{2}$ then the sufficient and necessary conditions of the ${\cal GW}_{\nu,\mu,a}$ model in (6) are obtained as special case. It is apparent that, to solve the identifiability problem, $l$ should be fixed (as in the ${\cal GW}$ model) or should depend on the other parameters and/or the dimension, so an interesting question is: which is the “best” $l\geq 0$ in the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ family?
This question can be responded by establishing an optimality criterion. Here we consider the criterion of maximum integral range that can be attained by a compactly supported correlation for a given smoothness and compact support. For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is defined as
$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\text{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{\mathbb{R}^{+}}x^{d-1}\phi(x)\text{d}x.$ (16)
$A$ is the isotropic version of the integral range proposed in Lantuéjoul, (1991) for stationary correlation functions. This measure has a close relation with the ergodicity of the RF, with the definition of an effective number of independent observations, and with the definition of long-range dependence (Yaglom,, 1987; Matheron,, 1989). In particular, following Lim and Teo, (2009), a RF with isotropic correlation $\phi$ is long-range dependent if $A=\infty$ and short-range dependent when $A<\infty$. If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$. Hristopoulos, (2020) interprets $A^{1/d}$ as an average measure of the distance over which two RF values are correlated.
When $d=1$, the integral range corresponds to the area under the unidimensional correlation function. A parametric correlation model that can cover a wide range of areas, from 00 to a specific upper bound, can be considered a flexible model (see Figure  1 and Theorem 6 hereunder). As a consequence, this can serve as a criterion for selecting between two different parametric correlation models, particularly when comparing two compactly supported correlation models that share the same smoothness and compact support parameters. In such cases, the correlation model that achieves a larger integral range is preferred.
For a parametric correlation model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we will denote its associated integral range as $A_{\boldsymbol{\theta},d}$. For instance, using formula 6.561.16 of Gradshteyn and Ryzhik, (2007), for the ${\cal MT}_{\nu,\alpha}$ correlation one has:
$A_{\nu,\alpha,d}=\frac{\alpha^{d}\pi^{d/2}\Gamma\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (17)
The following theorem provides an expression for the integral range associated with the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ correlation model.
Theorem 3.
The integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is given by:
$A_{\kappa,\mu,l,a,d}=a^{d}\left[\frac{\pi^{d/2}\Gamma(\frac{d+1}{2}+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+\frac{1}{2})\Gamma\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\left(\frac{\mu+1+d}{2}+\kappa+l\right)}\right].$ (18)
Note that, as expected, the compact support $a$ raised to power $d$ represents a scale factor in the integral range computation. In addition, using asymptotic properties of the Gamma function, it can be shown that $A_{\kappa,\mu,l,a,d}\sim\mu^{-d}$ as $\mu\to\infty$ and, as a consequence,
${\lim_{\mu\to\infty}}A_{\kappa,\mu,l,a,d}=0.$ (19)
This implies that, as the shape parameter $\mu$ increases, the integral range approaches zero, regardless of the compact support and smoothness parameters. The following theorem establishes that, for a given $l\geq 0$, the maximum integral range is obtained by fixing $\mu$ on the lower bound of its parametric space.
Theorem 4.
For an arbitrary $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ is attained when
  1. 1.
    $\mu=\mu_{1}(l):=\delta-l+\frac{1}{2}\quad$ if $\quad 0\leq l\leq\frac{d}{2}+\kappa$
  2. 2.
    $\mu=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>\frac{d}{2}+\kappa$.
For instance, let $\kappa\geq 0$. If $l=\frac{1}{2}$, then from (13) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of ${\cal GW}$ model (6).
As an example, Figure 1 (left part) depicts the ${\cal GW}_{0,\mu,1}$ model and the right part depicts the ${\cal GW}_{1,\mu,1}$ model for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$. In these examples $d=1$ and it is apparent that the model with maximum area corresponds to the case $\mu=\mu_{1}(\frac{1}{2})$ (red line) in both examples.
The next step is to find $l\geq 0$ that maximizes $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$.
Theorem 5.
The maximum of $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$ is attained when $l=d/2+\kappa$.
Refer to caption Refer to caption
Figure 1: Left part: ${\cal GW}_{0,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$ with $\delta=\frac{d+1}{2}+\kappa$ and $\kappa=0$. Right part: the same Figure but with $\kappa=1$. In both cases $d=1$.
As an example, Figure 2 compares $\mathcal{GH}_{\delta,\delta+\frac{\mu_{1}(l)}{2},\delta+\frac{\mu_{1}(l)}{2}+l,1}$ when $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines, respectively). In particular the left part depicts the case $\kappa=0$ and the right part depicts the case $\kappa=1$ ($d=2$ in both cases). It is apparent that the case $l=\frac{d}{2}+\kappa$ (blue line) corresponds to the model with higher integral range and improves the maximum integral range attained by the ${\cal GW}$ models (red line).
Refer to caption Refer to caption
Figure 2: Left part: $\mathcal{GH}_{\delta,\delta+\frac{\mu(l)}{2},\delta+\frac{\mu(l)}{2}+l,1}$ with $\delta=\frac{d+1}{2}+\kappa$, $\kappa=0$, $\mu(l)=\delta-l+\frac{1}{2}$ and $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines respectively). Right part: the same Figure but with $\kappa=1$. The case $l=\frac{1}{2}$ (red line) corresponds to the $\mathcal{GW}_{\kappa,\mu,1}$ model. The blue line corresponds to the correlation model that maximizes the integral range. In both cases $d=2$.
Summarizing our results, the model that can potentially achieve the maximum integral range is $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ and, in particular, the maximum is achieved when $\mu=\mu_{1}(d/2+\kappa)=1$. Taking into account (19), this implies that
$0\leq A_{\kappa,\mu,l,a,d}\leq a^{d}\left[\frac{\pi^{d/2}\,\Gamma\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\left(1+\kappa\right)\,\Gamma\left(1+\frac{d}{2}+2\kappa\right)}{2^{d}\,\Gamma\left(\kappa+\frac{1}{2}\right)\,\Gamma\left(1+\frac{d}{2}+\kappa\right)\,\Gamma\left(1+d+2\kappa\right)}\right].$
Accounting for (15), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting,, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría,, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron, (1965, I.4.18).
Hereafter, we call the correlation model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ as hypergeometric correlation and we denote as
$\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\quad x\geq 0,$ (20)
which is defined as
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma(\kappa+\frac{\mu+1}{2})\Gamma(2\kappa+\frac{d+\mu+1}{2})}{\Gamma(\mu+\frac{d+1}{2}+2\kappa)\Gamma(\kappa+\frac{1}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2};\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (21)
Necessary and sufficient validity conditions for $\mathcal{H}_{\kappa,\mu,a,d}$ can be easily obtained from Theorem 2 by setting $l=\frac{d}{2}+\kappa$. Specifically, given the smoothness parameter $\kappa>-\frac{1}{2}$, $\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}$ if and only if $\mu\geq 1$. Interestingly, the conditions are much simpler than the conditions (6) of the $\mathcal{GW}_{\kappa,\mu,a}$ model. In particular, the condition on $\mu$ does not depend on the dimension. The price to pay is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ depends on the dimension $d$.
The following theorem proves that the ${\cal H}_{\kappa,1,a,d}$ subclass of hypergeometric correlation functions with maximum integral range allows recovering any member of the ${\cal H}_{\kappa,\mu,a,d}$ class, as well as great part of the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ class, by scale mixtures. The chosen optimality criterion (maximizing the integral range) therefore coincides with the greatest flexibility from a modeling viewpoint.
Theorem 6.
Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$ and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ owing to Theorem 2. Then, $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is a beta-type scale mixture of hypergeometric correlation functions of the form $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
$\begin{split}&\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)=\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)}{\Gamma(\delta+\frac{1-d}{2})\Gamma(\delta+\frac{1}{2}+\kappa)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x){\rm d}u{\rm d}v,\quad x\geq 0.\end{split}$ (22)
In particular, for $l=\frac{d}{2}+\kappa$, one obtains
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu}{2}+\kappa)}{\Gamma(\delta+\frac{1-d}{2})\Gamma(\delta+\frac{1}{2}+\kappa)\Gamma^{2}(\frac{\mu-1}{2})}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x){\rm d}u{\rm d}v,\quad x\geq 0.\end{split}$
Clearly, the ${\cal H}_{\kappa,\mu,a,d}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$ in the same way as the ${\cal GW}_{\kappa,\mu,\beta}$ model. Figure 3 depicts two realizations of zero-mean and unit-variance Gaussian RFs with correlations ${\cal H}_{\kappa,4,0.2,2}$ for $\kappa=0,1,2$. The realizations were constructed via the Cholesky decomposition of the covariance matrix.
Refer to caption Refer to caption
Figure 3: Two Gaussian RF realizations with ${\cal H}_{\kappa,4,0.2,2}$ correlation model, for $\kappa=0,1$ (from left to right).
5 Some special cases and bridges to the Matérn model
The computation of the $\mathcal{H}_{\kappa,\mu,a,d}$ correlation function requires an efficient evaluation of the Gauss hypergeometric function, which is feasible (Johansson,, 2019) but could be cumbersome to (spatial) statisticians or machine learning practitioners used to handle closed-form parametric correlation models.
However, similarly to the $\mathcal{MT}$ and $\mathcal{GW}$ models, in some special cases, the computation can be considerably simplified. The following theorem considers the case when $\kappa$ is a non-negative integer, $d=1$ and $\mu\geq 1$.
Theorem 7.
If $\kappa=k$ is a non-negative integer, $d=1$ and $\mu\geq 1$, then:
$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\Gamma\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\left(k+0.5\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\times\sum\limits^{k}_{n=0}\frac{(n+k)!}{2^{2n}n!\Gamma(n+\frac{\mu}{2}+k+1)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
Unfortunately, general closed forms for any $d>1$, $\mu\geq 1$ and $\kappa$ a non-negative integer, are difficult to obtain. Some special cases can, however, be obtained using closed forms of the Gauss hypergeometric function. For instance, the following theorem provides a closed form expression when $d$ is an odd integer, $\mu\geq 1$ and $\kappa=0$.
Theorem 8.
Let $d$ an odd integer and $\mu\geq 1$. If $\kappa=0$ then
$\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum\limits^{\frac{d-1}{2}}_{n=0}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d-1}{2}+1\right)_{n}\left(1-\frac{x}{a}\right)^{n}_{+}}{2^{n}n!\Gamma\left(n+\frac{d-1}{2}+\mu+1\right)},$
with $C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\left(\frac{\mu+1}{2}\right)\Gamma\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}$.
Additional closed-form expressions, including the spherical models, for some fixed $\mu$ and $\kappa$ are given in Appendix B.
The ${\cal MT}$ model can be obtained as a special limit case of the proposed $\mathcal{H}_{\kappa,\mu,a,d}$ model after two suitable compact support reparameterizations. To obtain the first parameterization, the following theorem provides sufficient conditions of equivalence between two Gaussian measures with covariances $\sigma_{0}^{2}{\cal MT}_{\nu,\alpha}$ and $\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d}$, respectively. The proof can be found in Appendix A.
Theorem 9.
Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures. If $\nu=\kappa+\frac{1}{2}$ with $\nu>-\frac{1}{2}$ and
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}B(\kappa,\mu,d),$ (23)
then for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$ on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$, where
$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma(\frac{\mu+1}{2}+\kappa)\Gamma(\frac{\mu+d+1}{2}+2\kappa)}{\Gamma(\frac{\mu}{2})\Gamma(\frac{\mu+d}{2}+\kappa)}.$ (24)
Note that the equivalence condition (23) sharing a common variance parameter and taking into account the relation between the smoothness parameters $\nu=\kappa+\frac{1}{2}$, can be written as
$a=\alpha B(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
If we use the previous equation as a compact support reparameterization in the ${\cal H}$ model, then, using the same arguments as in Bevilacqua et al., (2022), it can be shown that
$\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa>-\frac{1}{2},$ (25)
with uniform convergence on $[0,\infty)$. In addition, following the same arguments in Emery et al., (2025), it can be shown that
$\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,\mu\alpha,d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa\geq-\frac{1}{2},$ (26)
with uniform convergence on $[0,\infty)$. As a consequence, the two reparameterized versions of the ${\cal H}$ model can be viewed as generalizations of the ${\cal MT}$ model with an additional parameter $\mu$ that, for given smoothness and spatial dependence parameters, allows to obtain compactly ($\mu<\infty$) or globally ($\mu=\infty$) supported covariance functions. Thus, under this parameterization, $\mu$ can be fixed by the user, particularly when dealing with large datasets, to achieve sparse covariance matrices and obtain computational benefits.
An illustration of (25) can be found in Figure 4 where the ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ model with increasing values of $\mu$ is compared with the ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$ model when $d=2$.
Refer to caption Refer to caption
Figure 4: The ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,2),2}$ for $\mu=1,2,10$ when $\kappa=0$ (left) and $\kappa=1$ (right). The red line is the ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$ correlation model ($\alpha=0.2$ in both cases).
It should be emphasized that, under this compact support reparameterization, the role of the shape parameter $\mu$ changes significantly. Let us assume, without loss of generality, that the scale dependence parameter $\alpha$ is equal to $1$. Under this assumption, the parameter $\mu$ (together with the smoothness parameter and the dimension $d$) determines the compact support. For large values of $\mu$, the compact support becomes larger, and both the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ and ${\cal H}_{\kappa,\mu,\mu,d}$ models approach the ${\cal MT}_{\kappa+\frac{1}{2},1}$ model. In this case, the maximum integral range is the one associated with the ${\cal MT}$ model in Equation (17).
Conversely, as $\mu$ approaches its lower bound, the compact support shrinks, and a potential smaller compact support results in a more flexible model. If we compare the compact support of $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ with that of $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ (see Equation 9) and given that $\mu$ is lower bounded by $1$ for the $\mathcal{H}$ model and by $\mu_{\min}(\kappa,d)$ for the $\mathcal{GW}$ model, it can be shown that
$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa)$
when $d\geq 2$ or $\kappa\geq 0$ and $b(1,1,\kappa,d)>g(1,\mu_{\min}(\kappa,d),\kappa)$ when $d=1$ and $-\frac{1}{2}<\kappa<0$. This implies that the $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model achieves a smaller compact support compared to the $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model for a significant portion of the parameter space. As a consequence, the $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model is more flexible from this perspective.
On the other hand, if we compare the smaller compact supports of $\mathcal{H}_{\kappa,\mu,\mu,d}$ and $\mathcal{GW}_{\kappa,\mu,\mu}$, then we reach the same conclusions since $1\leq\mu_{\min}(\kappa,d)$ when $d\geq 2$ or $\kappa\geq 0$ and $1>\mu_{\min}(\kappa,d)$ when $d=1$ and $-\frac{1}{2}<\kappa<0$.
We also compare the smaller compact supports that can be achieved by the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ and ${\cal H}_{\kappa,\mu,\mu,d}$ models, which are $b(1,1,\kappa,d)$ and $1$, respectively. Figure 5 illustrates the compact support $b(1,1,\kappa,d)$ for $-0.5<\kappa<1.5$ and $d=1,2,3,4,5$, compared to the reference value $1$. It can be observed that only for a small region of the smoothness parameter space and $d=1,2,3,4$, we obtain $b(1,1,\kappa,d)<1$.
Figure 5: The smaller compact support $b(1,1,\kappa,d)$ that can be achieved by the model ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ for $d=1,2,3,4,5$ and $-0.5<\kappa<1.5$. The red line is the smaller compact support that can be achieved by the ${\cal H}_{\kappa,\mu,\mu,d}$ model.
Summarizing our results, when comparing the two reparameterized versions of the $\mathcal{H}$ model with the analogous versions of the $\mathcal{GW}$ model that share the same smoothness parameter $\kappa$ and converge to the same $\mathcal{MT}$ model as the shape parameter $\mu$ becomes very large, it turns out that the reparameterized $\mathcal{H}$ models can achieve smaller compact supports. In addition the ${\cal H}_{\kappa,\mu,\mu,d}$ model can achieve smaller compact supports with respect to the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model.
6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
When considering the asymptotic properties of the ML estimator for the parameters of the $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ model, two different scenarios can be examined. Under increasing domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ can be estimated consistently. Furthermore, the asymptotic distribution of the ML estimator is Gaussian, with a covariance matrix equal to the inverse of the Fisher information matrix (Mardia and Marshall,, 1984).
Under fixed-domain, using the results of Theorem 1 and taking into account the parameterization (15) with $l=\frac{d}{2}+\kappa$, the equivalence condition for the two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing a common smoothness parameter, when $d=1,2,3$ and provided that $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, reads:
$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}},$ (27)
As a consequence only the microergodic parameter $\sigma^{2}B(\kappa,\mu,d)/a^{2\kappa+1}$ can be estimated consistently. On the other hand, assuming a fixed known shape parameter $\mu$, the microergodic parameter is given by $\sigma^{2}/a^{2\kappa+1}$.
In what follows, we provide the asymptotic properties of ML estimator of the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$. In particular, we will show that the microergodic parameter can be estimated consistently, and then we will assess the asymptotic distribution of the ML estimator.
Let $D$ be a bounded subset of $\mathbb{R}^{d}$ and let $\boldsymbol{Z}_{n}=(Z(\boldsymbol{s}_{1}),\ldots,Z(\boldsymbol{s}_{n}))^{\prime}$ be a finite realization of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, a zero-mean stationary Gaussian RF with a given parametric covariance function $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, with $\sigma^{2}>0$, $\boldsymbol{\tau}$ a parameter vector and $\phi$ a member of the class $\Phi_{d}$, with $\phi(0;\boldsymbol{\tau})=1$.
We then write $R_{n}(\boldsymbol{\tau})=[\phi(\|\boldsymbol{s}_{i}-\boldsymbol{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$ for the associated correlation matrix. The Gaussian log-likelihood function is defined as:
$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\left(n\log(2\pi\sigma^{2})+\log(|R_{n}(\boldsymbol{\tau})|)+\frac{1}{\sigma^{2}}\boldsymbol{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\boldsymbol{Z}_{n}\right).$ (28)
Under the $\cal H$ model, the Gaussian log-likelihood is obtained with $\phi(\cdot;\boldsymbol{\tau})\equiv\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$. Since in what follows $\mu$ and $\kappa$ are assumed known and fixed, for notation convenience, we write $\tau=a$.
Let $\hat{\sigma}^{2}_{n}$ and $\hat{a}_{n}$ be the ML estimator obtained maximizing $\mathcal{L}_{n}(\sigma^{2},a)$ for a fixed $\mu$ and $\kappa$. We consider two types of estimators. The first maximizes (28) with respect to $\sigma^{2}$ for a fixed arbitrary compact support $a>0$, obtaining the following estimator
$\hat{\sigma}_{n}^{2}(a)=\operatorname*{arg\,max}_{\sigma^{2}}\mathcal{L}_{n}(\sigma^{2},a)=\boldsymbol{Z}_{n}^{\prime}R_{n}(a)^{-1}\boldsymbol{Z}_{n}/n.$ (29)
Here $R_{n}(a)$ is the correlation matrix coming from the $\mathcal{H}_{\kappa,\mu,a,d}$ model. The following result establishes strong consistency and asymptotic distribution of the ML estimator $\hat{\sigma}_{n}^{2}(a)/a^{(1+2\kappa)}$ of the microergodic parameter. We omit the proof, as it follows the same steps as Theorem 8 in Bevilacqua et al., (2019).
Theorem 10.
Let $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ covariance model with $\mu\geq 1+d/2$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times(0,\infty)$. For a fixed $a>0$, let $\hat{\sigma}_{n}^{2}(a)$ as defined in Equation (29). Then, as $n\to\infty$,
  1. 1.
    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa}$ and
  2. 2.
    $\sqrt{n}(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,2(\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})^{2})$.
The second type of estimation technique considers the joint maximization of (28) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$ where $I=[a_{L},a_{U}]$ and $0<a_{L}<a_{U}<\infty$. The solution of this optimization problem is given by $(\hat{\sigma}_{n}^{2}(\hat{a}_{n}),\hat{a}_{n})$ where
$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\boldsymbol{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\boldsymbol{Z}_{n}/n$
and $\hat{a}_{n}=\operatorname*{arg\,max}_{a\in I}\mathcal{PL}_{n}(a)$. Here $\mathcal{PL}_{n}(a)$ is the profile log-likelihood:
$\mathcal{PL}_{n}(a)=-\frac{1}{2}\left(\log(2\pi)+n\log(\hat{\sigma}_{n}^{2}(a))+\log|R_{n}(a)|+n\right).$ (30)
In order to establish strong consistency and asymptotic distribution of the sequence of random variables $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$, we use the following lemma that establishes the monotone behavior of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ when viewed as a function of $a\in I$ under specific conditions on the $\mu$ parameter. The proof can be found in Appendix A.
Lemma 1.
Let $S_{n}=\{\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n}\in D\subset\mathbb{R}^{d}\}$ denote any set of distinct locations. For any $a_{1}<a_{2}\in I$ and for each $n$, $\hat{\sigma}_{n}^{2}(a_{1})/a_{1}^{1+2\kappa}\leq\hat{\sigma}_{n}^{2}(a_{2})/a_{2}^{1+2\kappa}$ if and only if $\mu\geq 4$.
Finally, the following theorem establishes the asymptotic properties of the ML estimator $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$. We omit the proof, as it follows the same steps as Theorem 9 in Bevilacqua et al., (2019) using the results in Lemma 1.
Theorem 11.
Let $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ covariance model with $\mu\geq 4$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$ where $I=[a_{L},a_{U}]$ with $0<a_{L}<a_{U}<\infty$. Let $(\hat{\sigma}_{n}^{2},\hat{a}_{n})^{\prime}$ maximize (28) over $(0,\infty)\times I$. Then as $n\to\infty$,
  1. 1.
    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa}$ and
  2. 2.
    $\sqrt{n}(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,2(\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})^{2})$.
We conclude this section with some comments on the reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$. In particular, assuming $\mu$ unknown and using (27), for the latter model the associated microergodic parameter is given by $\sigma^{2}B(\kappa,\mu,d)/(\alpha\mu)^{2\kappa+1},$ and for the former model, it is given by $\sigma^{2}/\alpha^{2\kappa+1}.$ Notably, the microergodic parameter in the second case does not depend on $\mu$. This implies that, for the $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ model, we have two microergodic parameters, $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$, that can be estimated consistently under fixed-domain asymptotics. As a consequence, from an inferential point of view, the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ should be preferred over the $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$ model. However, particularly for large datasets, the $\mu$ parameter should be fixed, as it allows the construction of sparse covariance matrices. In this case, the microergodic parameter is given by $\sigma^{2}/\alpha^{2\kappa+1}$ for both parameterizations.
7 A simulation study
In this section, we perform a simulation study on the estimation of the $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ family. We simulate 500 realizations, using Cholesky decomposition, of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $d=2$, setting $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$. We have considered a regular grid with increments $0.025$ over $[0,1]^{2}$, that is $1681$ grid points. Then the grid points have been perturbed, adding a uniform random value on $[-0.01,0.01]$ to each coordinate. We assume the parameters $\kappa$ and $\mu$ known and we focus on the ML estimation of $\sigma^{2}$ and $a$ and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing and fixed domain scenarios. Since the increasing as well as the fixed-domain frameworks can be mimicked by fixing the number of location sites over a given spatial domain and decreasing or increasing the spatial dependence (Zhang and Zimmerman,, 2005; Kaufman and Shaby,, 2013), we set the compact support $a=0.1,0.25,0.5$. We consider $a=0.1$ as an increasing domain scenario and $a=0.5$ as a fixed domain scenario.
Table 1 reports biases and standard deviations (in parentheses) of the 500 ML parameter estimates, namely $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$, $\kappa=0,1,2$, $\mu=4,6,8$ and $a=0.1,0.25,0.5$.
Table 1: Bias and standard deviation (in parenthesis) of ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ when $\kappa=0,1,2$, $\mu=4,6,8$, $a=0.1,0.25,0.5$ and $\sigma^{2}=1$.
$\kappa=0$ $\kappa=1$ $\kappa=2$
$\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$
$\widehat{\sigma}^{2}$ $a=0.1$ $0.0037$ $0.0025$ $0.0018$ $0.0060$ $0.0039$ $0.0028$ $0.0053$ $0.0038$ $0.0029$
$(0.0422)$ $(0.0382)$ $(0.0368)$ $(0.0485)$ $(0.0427)$ $(0.0398)$ $(0.0475)$ $(0.0432)$ $(0.0407)$
$a=0.25$ $0.0030$ $0.0047$ $0.0046$ $0.0085$ $0.0098$ $0.0095$ $0.0100$ $0.0116$ $0.0111$
$(0.0808)$ $(0.0622)$ $(0.0522)$ $(0.1082)$ $(0.0870)$ $(0.0737)$ $(0.1044)$ $(0.088)$ $(0.0773)$
$a=0.5$ $-0.0023$ $0.0024$ $0.0045$ $0.0029$ $0.0079$ $0.0101$ $0.0052$ $0.0091$ $0.0116$
$(0.1446)$ $(0.1103)$ $(0.0900)$ $(0.1921)$ $(0.1592)$ $(0.135)$ $(0.1892)$ $(0.1599)$ $(0.1397)$
$\widehat{a}$ $a=0.1$ $0.0006$ $0.0007$ $0.0007$ $0.0004$ $0.0004$ $0.0004$ $0.0002$ $0.0003$ $0.0003$
$(0.0049)$ $(0.0055)$ $(0.0068)$ $(0.0024)$ $(0.0028)$ $(0.0033)$ $(0.0017)$ $(0.0021)$ $(0.0026)$
$a=0.25$ $0.0007$ $0.0014$ $0.0016$ $0.0005$ $0.0008$ $0.0009$ $0.0004$ $0.0006$ $0.0007$
$(0.0214)$ $(0.0176)$ $(0.0157)$ $(0.0099)$ $(0.0086)$ $(0.0079)$ $(0.0059)$ $(0.0054)$ $(0.0052)$
$a=0.5$ $-0.0015$ $0.0010$ $0.0023$ $-0.0016$ $-0.0001$ $0.0008$ $-0.0009$ $-0.0001$ $0.0005$
$(0.0737)$ $(0.0578)$ $(0.0488)$ $(0.0329)$ $(0.0280)$ $(0.0245)$ $(0.0196)$ $(0.0171)$ $(0.0154)$
$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ $-0.0150$ $-0.0219$ $-0.0151$ $-3.9684$ $-4.4120$ $-3.7659$ $-496.06$ $-486.01$ $-346.94$
$(0.4370)$ $(0.5362)$ $(0.6792)$ $(51.3695)$ $(68.6746)$ $(88.9434)$ $(6397.8)$ $(8930.7)$ $(11628.7)$
$a=0.25$ $0.0045$ $0.0015$ $-0.0017$ $0.0228$ $-0.0470$ $-0.1218$ $-0.2553$ $-1.7130$ $-3.0083$
$(0.1499)$ $(0.1600)$ $(0.1708)$ $(2.5102)$ $(2.7402)$ $(2.9959)$ $(41.521)$ $(46.056)$ $(51.067)$
$a=0.5$ $0.0034$ $0.0033$ $0.0026$ $0.0113$ $0.0099$ $0.0065$ $0.0399$ $0.0319$ $0.0150$
$(0.0719)$ $(0.0740)$ $(0.0765)$ $(0.2923)$ $(0.3050)$ $(0.3187)$ $(1.1922)$ $(1.2484)$ $(1.3091)$
Overall, the ML estimation of $\sigma^{2}$ and $a$ is approximately unbiased, and the standard deviations increase when switching from the increasing domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$), as expected. Another general pattern that can be observed is that, for each $\kappa$, increasing the shape parameter $\mu$ results in a lower standard deviation.
For the microergodic parameter, it can be observed that under the increasing domain scenario ($a=0.1$), the ML estimation can be significantly unbiased but with a high standard deviation, particularly when $\kappa$ increases. When moving to the fixed-domain scenario ($a=0.5$), both the bias and the standard deviation are significantly reduced, as expected. In addition, the standard deviation of the microergodic parameter increases as the smoothness parameter $\kappa$ increases.
Figure 6 depicts the boxplots of the ML estimates of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (on the left) and $a=0.5$ (on the right) when $\kappa=0$. The Gaussian distribution of the estimates for $\sigma^{2}$ and $a$ is more evident under the increasing domain scenario, as expected. On the other hand, when switching to the fixed-domain scenario, the distribution of the ML estimates of the microergodic parameter becomes more centered around the true values, and the distribution appears Gaussian, as expected.
Refer to caption Refer to caption
($a=0.1$) ($a=0.5$)
Figure 6: Left part: Boxplots of ML estimates of $\sigma^{2}$, $a$ and the microergodic parameter $\frac{\sigma^{2}}{a^{2\kappa+1}}$ (from left to right) when estimating the covariance model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $\sigma^{2}=1$, for $\mu=4,6,8$ and $a=0.1$ (increasing domain scenario). Right part: the same Figure with $a=0.5$ (fixed domain scenario). In both cases $\kappa=0$.
8 A real data application
We consider data from WorldClim (www.worldclim.org), a global database of high spatial resolution weather and climate data for the years 1970–2000, which can be downloaded using the R package geodata (Hijmans et al.,, 2024). Specifically, we analyze mean temperature data for September observed at 725,330 georeferenced locations across Madagascar. To make ML estimation feasible, we select a random sample of 3,500 locations and perform the analysis using a UTM projection.
Following Li and Zhang, (2011), we first detrend the data using splines to remove the cyclic patterns of both variables along the two coordinates. We then treat the residuals $y(\boldsymbol{s}_{i})$, $i=1,\ldots,3500$, as realizations from a zero-mean RF $\{Y(\boldsymbol{s}):\boldsymbol{s}\in D\}$. Figure 7, from left to right, depicts the colored map, the boxplot, and the empirical semivariogram of the residuals. The graphics suggest that a weakly stationary RF with a marginal distribution capable of modeling heavy tails is potentially an appropriate model for our data.
Refer to caption Refer to caption Refer to caption
Figure 7: From left to right: spatial map of the residuals for the Madagascar mean temperature data, boxplot, and empirical semivariogram of the data.
Let $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ be a standard Gaussian RF with an isotropic parametric correlation function. In our analysis, we consider two specifications for $Y(\boldsymbol{s})$:
  • a zero-mean Gaussian RF, i.e., $Y(\boldsymbol{s})=\sigma Z(\boldsymbol{s})$, and
[…]
${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$ ${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$
$\sigma^{2}$ $1.2054$ $1.2066$ $1.2061$ $0.6998$ $0.6986$ $0.6984$
$(0.0553)$ $(0.0541)$ $(0.0540)$ $(0.0418)$ $(0.0403)$ $(0.0403)$
$\kappa$ $0.6079$ $0.0305$ $0.0313$ $0.5767$ $-0.0036$ $-0.0043$
$(0.0293)$ $(0.0307)$ $(0.0312)$ $(0.0265)$ $(0.0237)$ $(0.0237)$
$\beta$ $16.0229$ $85.8058$ $87.9751$ $18.5405$ $76.0127$ $77.6552$
$(1.3104)$ $(20.4440)$ $(20.9644)$ $(1.6183)$ $(12.0024)$ $(10.5911)$
$\mu$ $4.2150$ $3.8914$ $3.1346$ $2.8085$
$(1.3021)$ $(1.3026)$ $(0.6548)$ $(0.5764)$
$h$ $0.1885$ $0.1899$ $0.1898$
$(0.0172)$ $(0.0170)$ $0.0170$
loglik $-3900.01$ $-3896.88$ $-3896.94$ $-3689.88$ $-3683.81$ $-3683.94$
AIC $7806$ $7802$ $7802$ $7388$ $7378$ $7378$
RMSE $0.6593$ $0.6588$ $0.6588$ $0.6586$ $0.6581$ $0.6581$
MAE $0.4068$ $0.4065$ $0.4065$ $0.4073$ $0.4071$ $0.4070$
[…]
$Time$ $7.22$ $0.04$ $0.04$ $6.26$ $0.03$ $0.03$
[…]
Refer to caption Refer to caption
Refer to caption Refer to caption
[…]
9 Concluding Remarks
In this paper, we study the Gauss hypergeometric covariance model, a highly flexible kernel introduced by Emery and Alegría, (2022), which generalizes both the Matérn and Generalized Wendland models. We establish necessary and sufficient conditions for the validity of the model and highlight that the model suffers from identifiability issues under both increasing and fixed domain asymptotics. To address the identifiability problem, we introduce a new class of compactly supported kernels that satisfy a specific optimality criterion. In particular, the proposed model achieves the maximum integral range within the class of Gauss hypergeometric correlations. From this perspective, the proposed model can be regarded as an enhancement of the Generalized Wendland model.
The proposed model has the potential to outperform the Matérn family in terms of model fitting, prediction performance, and computational efficiency, as demonstrated in the real-data application.
Additionally, the proposed model can be easily implemented, as efficient numerical computation of the Gaussian hypergeometric function is available in various libraries, such as the GNU Scientific Library (Gough,, 2009), and in major statistical software packages, including R, MATLAB, and Python. In particular, the R package GeoModels (Bevilacqua et al.,, 2025), used in this paper for numerical experiments and applications, computes the proposed model using the Python implementation of the Gaussian hypergeometric function in the SciPy library (Virtanen et al.,, 2020).
From an inferential perspective, the results on the asymptotic properties of the maximum likelihood estimator of the microergodic parameter could be generalized to include a nugget effect, as in Tang et al., (2021). Additionally, the estimation of the smoothness parameter could, in principle, be studied following the ideas in Loh and Sun, (2023). In a Bayesian setup, the fixed-domain asymptotic posterior distribution of the microergodic parameter could be investigated following Li, (2022). Finally, the proposed model can be generalized in various directions, such as space-time, multivariate, or non-stationary contexts. These topics are left for future research.
Acknowledgement
We would like to thank Professor Igor Kondrashuk for his help in partially proving Lemma 1. This work was funded by the National Agency for Research and Development of Chile, under grants ANID FONDECYT 1240308 (M. Bevilacqua), ANID project Data Observatory Foundation DO210001 (M. Bevilacqua), MATH-AMSUD1167 22-MATH-06 (AMSUD220041) (M. Bevilacqua), ANID FONDECYT 1250862 (C. Caamaño-Carrillo) and ANID PIA AFB230001 (X. Emery).
Appendix A Proofs
A.1 Proof of Theorem 1
Proof.
Let $P(\rho_{i})$, $i=0,1$, be two zero-mean Gaussian measures with isotropic covariance function $\rho_{i}$ and associated spectral density $\widehat{\rho}_{i}$, $i=0,1$, as defined through (1). Using results in Skorokhod and Yadrenko, (1973) and Ibragimov and Rozanov, (1978), Stein, (2004) has shown that, if for some $a>0$, $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and for some finite and positive $c$,
$\int_{c}^{\infty}z^{d-1}\;\left\{\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right\}^{2}\;{\rm d}z<\infty,$ (31)
then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$. Then (31) is the key tool to prove Theorem 1.
We first need the results in Lemma 2 hereunder. They follow from Zastavnyi, (2006), Chernih and Hubbert, (2014), and from standard properties of Fourier transforms and give asymptotic expansion of the spectral density of the $\mathcal{GH}$ model. Their proofs are thus omitted. For two given functions $g_{1}(x)$ and $g_{2}(x)$, with $g_{1}(x)\asymp g_{2}(x)$ we mean that there exist two constants $c$ and $C$ such that $0<c<C<\infty$ and $c|g_{2}(x)|\leq|g_{1}(x)|\leq C|g_{2}(x)|$ for each $x$.
Lemma 2.
Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be the function defined at Equation (11). Then, for $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $12(\beta+\gamma)\geq 6\delta+$ 1
  1. 1.
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\sigma^{2}Ka^{d}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
        $+\,c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}$,  for $z\to\infty$;
  2. 2.
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta}$,  for $z\to\infty,$
where $c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}$, $c_{4}=\frac{\pi^{-\frac{1}{2}}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)}$ and $c_{5}=\frac{\pi}{2}(\frac{1}{2}+\delta-\beta-\gamma)$.
We are now ready to prove the Theorem. Let us start with the sufficient part of the assertion. From Lemma  2 (Point 2), we have:
$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}\leq c_{2},\quad i=0,1.$
We then proceed by direct construction, and, using Lemma 2 (Points 1 and 2) and (31) we find that, as $z\to\infty$,
$\Big{|}\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\Big{|}\leq\Big{|}\sigma_{1}^{2}K_{1}a_{1}^{d}\Big{[}c_{3}^{1}(za_{1})^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
$\displaystyle+\,c_{4}^{1}(za_{1})^{\lambda_{1}}\big{\{}\cos(za_{1}+c_{5}^{1})+\mathcal{O}(z^{-1})\big{\}}\Big{]}-\sigma_{0}^{2}K_{0}a_{0}^{d}\Big{[}c_{3}^{0}(za_{0})^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
$+\,c_{4}^{0}(za_{0})^{\lambda_{0}}\big{\{}\cos(za+c_{5}^{0})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\Big{|}\frac{z^{2\delta}}{c_{1}}$
$\leq\frac{1}{c_{1}}\Big{|}\big{[}c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big{]}\{1+\mathcal{O}(z^{-2})\}+\big{[}c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z-c_{5}^{1})$
$-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z-c_{5}^{0})\big{]}+\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}\Big{|},$
where $c_{3}^{i}=\frac{2^{2\delta}\Gamma(\beta_{i})\Gamma(\gamma_{i})}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$, $c_{4}^{i}=\frac{\pi^{-\frac{1}{2}}\Gamma(\beta_{i})\Gamma(\gamma_{i})}{2^{\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}}\Gamma(\delta)}$,$c_{5}^{i}=\frac{\pi}{2}(\frac{1}{2}+\delta-\beta_{i}-\gamma_{i})$ and $\lambda_{i}=\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}$, $i=0,1$.
Let us now write
$A(z)$ $=\big{[}c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta)}K_{0}\big{]}\{1+\mathcal{O}(z^{-2})\},$
$B(z)$ $=\big{[}c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z-c_{5}^{1})-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z-c_{5}^{0})\big{]}$
$D(z)$ $=\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}.$
Then, for some positive and finite $c$, Condition (31) can be written as
$\int_{c}^{\infty}z^{d-1}\bigg{(}\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\biggr{)}^{2}{\rm d}z<\infty,$ (32)
being equivalent to
$\int_{c}^{\infty}z^{d-1}\big{(}A(z)+B(z)+D(z)\big{)}^{2}{\rm d}z<\infty.$
Now, we have that $A(z)$ is identically equal to zero thanks to our condition. We claim that the partial integral that depends of $B(z)$ and $D(z)$ is finite if $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$. In fact, we have, for $z\to\infty$,
$|B(z)|$ $\leq c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\leq c_{6}\min\{z^{\lambda_{1}+2\delta},z^{\lambda_{0}+2\delta}\}$
and
$|D(z)|$ $\leq\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}$
$\leq c_{7}\min\{z^{\lambda_{1}+2\delta},z^{\lambda_{0}+2\delta}\},$
with $c_{i}$ being positive and finite constants, $i=6,7$. Thus, if $\frac{\sigma_{0}^{2}L_{0}}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L_{1}}{a_{1}^{2\delta-d}}$, we have
$\int_{c}^{\infty}$ $z^{d-1}\left(\frac{\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\right)^{2}{\rm d}z$ (33)
$\leq\int_{c}^{\infty}z^{d-1}(B(z)+D(z))^{2}{\rm d}z$
$\leq\int_{c}^{\infty}z^{d-1}\Big{(}B(z)^{2}+D(z)^{2}+2B(z)D(z)\Big{)}{\rm d}z.$
We conclude that (33) is finite if $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$. The sufficient part of our claim is thus proved.
The necessary part follows the proof of Zhang, (2004). We suppose that $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$, $\frac{\sigma_{0}^{2}L_{0}}{a_{0}^{2\delta-d}}\neq\frac{\sigma_{1}^{2}L_{1}}{a_{1}^{2\delta-d}}$ and let $\sigma_{2}^{2}=\sigma_{0}^{2}L_{0}\left(\frac{a_{1}}{a_{0}}\right)^{2\delta-d}$. Then $\mathcal{GH}_{\delta,\beta_{0},\gamma_{0},a_{0}}$ and $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{2}}$ induce two equivalent Gaussian measures. This, we need to show that $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{2}}$ and $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{1}}$ define two orthogonal Gaussian measures. The rest of the proof follows the same arguments in Zhang, (2004). ∎
A.2 Proof of Theorem 2
Proof.
Let us consider the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model and, for the moment, let us assume that $l$ is a real number. The sufficient conditions (a), (b), (c) given in Section 3, under the parameterization (15) are:
  1. (a)
    $\kappa>-\frac{1}{2}$;
  2. (b)
    $\mu(\mu+2l)\geq d+2\kappa+1$;
  3. (c)
    $\mu+l\geq\frac{d}{2}+\kappa+1$.
Solving the previous system of inequalities, we derive the following sufficient conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l\leq-\frac{(1+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(1+2\kappa)}{2}<l\leq\frac{(1+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l>\frac{(1+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l<-\frac{(d+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa)}{2}\leq l\leq\frac{(d+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
Necessary conditions can be obtained following Cho and Yun, (2018, Theorem 6.1), who show that if $\mathstrut{}_{1}F_{2}\left(a;b;c;-\frac{x^{2}}{4}\right)\geq 0$, then $b>a$, $c>a$, $b+c\geq 3a+\frac{1}{2}$. Using the spectral density (11), under the parameterization (15), we have
  1. (a)
    $\mu>0$;
  2. (b)
    $\mu>-2l$;
  3. (c)
    $\mu+l\geq\frac{d}{2}+\kappa+1$.
Solving the system of inequalities (under the restriction $\kappa>-\frac{1}{2}$) for $d\geq 1$, we derive the following necessary conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu>-2l$ and $l\leq-\frac{(3+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(3+2\kappa)}{2}<l\leq\frac{(3+2\kappa)}{2}$.
    • d)
      $\mu>0$ and $l>\frac{(3+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu>-2l$ and $l<-\frac{(d+2\kappa+2)}{2}$
    • c)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa+2)}{2}\leq l<\frac{(d+2\kappa+2)}{2}$.
    • d)
      $\mu>0$ and $l\geq\frac{(d+2\kappa+2)}{2}$.
Putting together the necessary and sufficient conditions for the cases $d=1$ and $d>1$, we obtain the following conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l\leq-\frac{(1+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(1+2\kappa)}{2}<l\leq\frac{(1+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l>\frac{(1+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l<-\frac{(d+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa)}{2}\leq l\leq\frac{(d+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
We observe that the necessary and sufficient conditions are symmetric around $l=0$, since $\mathcal{GH}_{\delta,\beta,\gamma,a}(x)=\mathcal{GH}_{\delta,\gamma,\beta,a}(x)$. Specifically, for $d=1$, conditions a) and c) are symmetric when $l<-\frac{(1+2\kappa)}{2}$ and $l>\frac{(1+2\kappa)}{2}$, respectively, while condition b) is symmetric when $-\frac{(d+2\kappa)}{2}<l\leq 0$ and $0\leq l\leq\frac{(d+2\kappa)}{2}$. These same symmetry patterns are also observed for the case $d>1$.
Using the symmetry of the conditions, we put together the necessary and sufficient conditions only for $l\geq 0$ obtaining the main result, that is for $\kappa>-\frac{1}{2}$ the model is valid if and only if
  1. 1.
    $\mu\geq\delta-(l-\frac{1}{2})$ and $0\leq l\leq\frac{(d+2\kappa)}{2}$
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
A.3 Proof of Theorem 3
Proof.
The result stems from formula 2.21.1.11 of Prudnikov et al., (1990) and the expressions (12) and (16) of the hypergeometric correlation and of the integral range.
A.4 Proof of Theorem 4
Hereafter, without loss of generality consider a unit compact support $a=1$. Let us define $LA_{\kappa,d}(\mu,l):=log(A_{\kappa,\mu,l,d,1})$ where
$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\Gamma\left(\frac{\mu+1}{2}+\kappa+d/2\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l+d/2\right)}.$
is the integral range.
[…]
  1. 1.
    $\mu\geq\delta-l+0.5\quad$ if $\quad 0\leq l\leq d/2+\kappa$
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>d/2+\kappa$.
This is a maximization problem over a bounded interval. Hereafter, with $\psi(x)=\frac{\text{d}log\Gamma(x)}{\text{d}x}$ we denote the digamma function which is strictly increasing for $x>0$. The fist derivative of $LA_{\kappa,d}(\mu,l)$ with respect to $\mu$ is given by
$\frac{\text{d}LA_{\kappa,d}(\mu,l)}{\text{d}\mu}=0.5[\psi(x)-\psi(x+d/2)+\psi(x+l)-\psi(x+l+d/2)],$
where $x=\frac{\mu+1}{2}+\kappa$. It is a negative function since $\psi(\cdot)$ is strictly increasing and this implies that the function $LA_{\kappa,d}(\mu,l)$ is monotone non-increasing as a function of $\mu$. As a consequence the maximum is achieved by $LA_{\kappa,d}(\mu,l)$ on the lower boundary of region 1) and region 2), that is
  1. 1.
[…]
  2. 2.
[…]
where $\mu_{1}(l)=\delta-l+0.5$ and $\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l$. This proves Theorem 4.
A.5 Proof of Theorem 5
We focus on the functions $LA_{\kappa,d}(\mu_{i}(l),l)$ for $i=1,2$ defined on the proof of Theorem 4 and the goal is to find the maximum with respect to $l$.
If $i=1$, the derivative of $LA_{\kappa,d}(\mu_{1}(l),l)$ for $0\leq l\leq d/2+\kappa$ is given by
$\frac{\text{d}LA_{\kappa,d}(\mu_{1}(l),l)}{\text{d}l}=0.5[\psi(x+l/2)-\psi(x-l/2)+\psi(x-l/2+d/2)-\psi(x+l/2+d/2)],$
where $x=\frac{d}{4}+\frac{3\kappa}{2}+1$. It can be shown that this function is non-negative using the fact that the derivative of the digamma function $\frac{\text{d}\psi(x)}{\text{d}x}$ is a decreasing function. This implies that the maximum of $LA_{\kappa,d}(\mu_{1}(l),l)$ is achieved on the upper bound of the parametric space, that is $l=d/2+\kappa$.
Using similar arguments, if $i=2$ it can be shown that the derivative of $LA_{\kappa,d}(\mu_{2}(l),l)$ for $l>d/2+\kappa$ is non-positive. As a consequence the maximum is achieved on the lower bound of the parametric space, that is when $l=d/2+\kappa+\epsilon$ where $\epsilon$ an arbitrary small positive constant.
Finally we can prove that $LA_{\kappa,d}(\mu_{1}(d/2+\kappa),d/2+\kappa))=LA_{\kappa,d}(1,d/2+\kappa)$ is greater than $LA_{\kappa,d}(\mu_{2}(d/2+\kappa+\epsilon),d/2+\kappa+\epsilon))=LA_{\kappa,d}(1-g(\epsilon),d/2+\kappa+\epsilon)$ where $g(\epsilon)$ is a positive constant.
A.6 Proof of Theorem 6
Proof.
Using (11) and Equation (25) in Emery and Alegría, (2022), the Fourier transform of the right-hand side member of (22) is found to be
$\begin{split}&\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)\Gamma(\delta)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{1}{2})\Gamma(2\delta)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{d+2\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathstrut_{1}F_{2}\Big{(}\delta;\delta+\frac{1}{2},2\delta;-\frac{uv(za)^{2}}{4}\Big{)}{\rm d}u{\rm d}v\\ &=\frac{\Gamma(\delta)\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{\mu}{2})\Gamma(\delta+\frac{\mu}{2}+l)}\mathstrut_{1}F_{2}\Big{(}\delta;\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l;-\frac{(za)^{2}}{4}\Big{)}\\ &=\widehat{\mathcal{GH}}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(z),\quad z\geq 0.\end{split}$
The claim of the theorem is obtained by inverse Fourier transformation. ∎
A.7 Proof of Theorem 7
If $d=1$ and $\kappa=k$ is a non-negative integer, using formula 7.3.1.70 of Prudnikov et al., (1990) and applying the identity
$P_{a}^{b}(x)=P^{b}_{-a-1}(x),$
where $P_{a}^{b}(x)$ denotes the associated Legendre function (of the first kind) with degree $a$ and order $b$ (see Abramowitz and Stegun, (1970, 8.3.1)), one can rewrite (21) as
$\mathcal{H}_{k,\mu,a,1}(x)$ $=$ $\frac{\Gamma(\frac{\mu}{2}+2k+1)}{\Gamma(k+\frac{1}{2})}\sqrt{\pi}\left(1-\frac{x^{2}}{a^{2}}\right)^{\frac{\mu}{2}+k}\left(\frac{x}{a}\right)^{k}P^{-\frac{\mu}{2}-k}_{k}\left(\frac{a^{2}+x^{2}}{2ax}\right).$
The claim of the Theorem then derives from the following result, which is valid for $a$ being a non-negative integer and $z>1$:
$P^{-b}_{a}(z)=\frac{(z-1)^{b/2}}{(z+1)^{b/2}}\sum_{j=0}^{a}\frac{(j+a)!(0.5(z-1))^{j}}{j!\Gamma(j+b+1)(a-j)!}.$ (35)
A.8 Proof of Theorem 8
Proof.
For $\mu\geq 1$ and $\kappa=0$, applying formula 7.3.1.41 from Prudnikov et al., (1990) to Equation (21), one gets
$\mathcal{H}_{0,\mu,a,d}(x)=\frac{\Gamma(\frac{1+\mu}{2})\Gamma(\frac{1+\mu+d}{2})}{\sqrt{\pi}}2^{\mu+\frac{d-1}{2}}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{\mu}{2}+\frac{d-1}{4}}\,P_{\frac{d-1}{2}}^{\frac{1-d}{2}-\mu}\left(\frac{x}{a}\right).$
The claim of the Theorem stems from the fact that, for $m$ being an integer, the associated Legendre function of degree $m$ is a terminating series:
$P_{m}^{\eta}(z)=\left(\frac{1+z}{1-z}\right)^{\frac{\eta}{2}}\sum\limits_{n=0}^{m}\frac{(-m)_{n}(m+1)_{n}}{n!\Gamma(n-\eta+1)}\left(\frac{1-z}{2}\right)^{n},\quad 0<z<1.$ (36)
A.9 Proof of Theorem 9
Proof.
For notation convenience we consider the notation ${\cal GH}_{\delta,\beta,\gamma,a}$ for the $\mathcal{H}_{\kappa,\mu,a,d}$ model, recalling that (see 20) if
$\delta=\frac{d+1}{2}+\kappa,\quad\beta=\delta+\frac{\mu}{2},\quad\gamma=\delta+\frac{\mu}{2}+d/2+\kappa,$ (37)
then $\mathcal{H}_{\kappa,\mu,a,d}=\mathcal{GH}_{\delta,\beta,\gamma,a}.$
In order to prove Theorem 9, as in the proof of Theorem 1, we use (31) to find conditions such that, for some positive and finite $c$,
$\int_{c}^{\infty}z^{d-1}\bigg{(}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\biggr{)}^{2}{\rm d}z<\infty.$ (38)
It is known that $\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha})(z)z^{\tau}$ is bounded away from 0 and $\infty$ as $z\to\infty$ for some $\tau>0$ (Zhang,, 2004). Let $\lambda$ be a parameter such as $\lambda=\frac{1}{2}+\delta-\beta-\gamma$. Using Points 1 of Lemma 2 , we have, as $z\to\infty$,
$\Bigg{|}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-\frac{d}{2}}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}(\alpha^{-2}+z^{2})^{\nu+\frac{d}{2}}-1\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{\frac{d}{2}}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\frac{(({\alpha z})^{-2}+1)^{\nu+\frac{d}{2}}}{z^{-2\nu-d}}-1\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{\frac{d}{2}}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\times$
$\quad\frac{\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}}{z^{-2\nu-d}}-1\Bigg{|}.$
$=$ $\Bigg{|}w_{1}z^{2\nu+d-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}$
$\quad+w_{2}z^{\lambda+2\nu+d}\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}-1\Bigg{|},$
where $w_{1}=\frac{\sigma_{1}^{2}Ka^{d-2\delta}\Gamma(\nu)c_{3}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-\frac{d}{2}}}$, $w_{2}=w_{1}c_{4}a^{\lambda}/c_{3}$.
Since $z^{2\nu+d}\big{[}(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\big{]}=\mathcal{O}(z^{2\nu+d-2})$, we have
$\int_{c}^{\infty}z^{d-1}\Bigg{|}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg{|}^{2}{\rm d}z$
$=\int_{c}^{\infty}z^{d-1}\Bigg{|}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})+\big{\{}w_{1}z^{2\nu+d-2\delta}-1\big{\}}+w_{1}z^{-2\delta}$
×$\quad\times\big{\{}\mathcal{O}(z^{2\nu+d-2})+\mathcal{O}(z^{2\nu+d-4})\big{\}}+w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5}^{\boldsymbol{\varsigma}})$
$\quad+\mathcal{O}(z^{-1})\big{\}}\Bigg{|}^{2}{\rm d}z.$
For assessing the last integral, the following is relevant:
  • (i)
    $w_{1}z^{2\nu+d-2\delta}-1=0$ if $\delta=\nu+\frac{d}{2}$ and $w_{1}=1$.
  • (ii)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}^{2}{\rm d}z<\infty$ if $d=1,2,3$ and $2\nu+d=2\delta$.
  • (iii)
    $\int_{c}^{\infty}z^{d-1}\Big{(}w_{1}z^{-2\delta}\big{\{}\mathcal{O}(z^{2\nu+d-2})+\mathcal{O}(z^{2\nu+d-4})\big{\}}\Big{)}^{2}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+\frac{1}{2}$.
  • (iv)
    $\int_{c}^{\infty}z^{d-1}\Big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5}^{\boldsymbol{\varsigma}})+\mathcal{O}(z^{-1})\big{\}}\Big{)}^{2}{\rm d}z<\infty$ if $\nu=\kappa+1/2$.
  • (v)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{-2})(\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d})\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
  • (vi)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}\big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
  • (vii)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{-2})\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{)}\big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}$$\times\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
This allows us to conclude that, for a given $\kappa>-\frac{1}{2}$, if $w_{1}=1$, $\nu=\kappa+1/2$ and $d=1,2,3$ then (38) holds and thus $P(\sigma^{2}_{0}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a})$. Condition $w_{1}=1$ is equivalent to
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}A}{a_{1}^{2\delta-d}},$
where $A=\frac{2^{2\delta-d}\Gamma(\beta-d/2)\Gamma(\gamma-d/2)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}$ and, taking into account the parameterization (37), the previous condition can be written as
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}B(\kappa,\mu,d),$
where $B(\kappa,\mu,d)$ is defined in (24).
A.10 Proof of Lemma 1
Proof.
The proof follows Kaufman and Shaby, (2013). Let $0<a_{1}<a_{2}$, with $a_{1},a_{2}\in I$. Then, for any $\boldsymbol{Z}_{n}$,
$\hat{\sigma}_{n}^{2}(a_{1})/a_{1}^{1+2\kappa}-\hat{\sigma}_{n}^{2}(a_{2})/a_{2}^{1+2\kappa}=\frac{1}{n}\boldsymbol{Z}_{n}^{\prime}(R_{n}(a_{1})^{-1}a_{1}^{-(1+2\kappa)}-R_{n}(a_{2})^{-1}a_{2}^{-(1+2\kappa)})\boldsymbol{Z}_{n}$
is nonnegative if the matrix $R_{n}(a_{1})^{-1}a_{1}^{-(1+2\kappa)}-R_{n}(a_{2})^{-1}a_{2}^{-(1+2\kappa)}$ is positive semidefinite and this happens if and only if the matrix $B=R_{n}(a_{2})a_{2}^{1+2\kappa}-R_{n}(a_{1})a_{1}^{1+2\kappa}$ with generic element
$B_{ij}=a_{2}^{1+2\kappa}\mathcal{H}_{\kappa,\mu,a_{2},d}(||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||)-a_{1}^{1+2\kappa}\mathcal{H}_{\kappa,\mu,a_{1},d}(||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||)$
is positive semidefinite. The last assertion holds if and only if the function
$f(z)=a_{2}^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0$ (39)
for all $z>0$. Here ^$\widehat{\mathcal{H}}_{\kappa,\mu,a,d}$ is the spectral density defined in (11) that is $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}$ under the parameterization (37). Condition (39) is true if and only if the function $f_{1}(t)=t^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)$ is an increasing function with respect to $t>0$ for all $z>0$. Following the same arguments in Faouzi et al., (2020), $f_{1}(t)$ is increasing if and only if $g(t)=t^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t)$ is increasing and a sufficient condition is given by $\frac{dg(t)}{dt}>0$. Using contour integral representation of the $\mathstrut{}_{1}F_{2}\Big{(}\cdot;\cdot,\cdot;-t^{2}\Big{)}$ function and setting $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+\frac{d}{2}+\kappa$, and $M$ a positive constant not depending on $t$, the derivative is given by:
$\frac{dg(t)}{dt}$ $=\frac{d}{dt}\left[Mt^{2\kappa+1+d}\mathstrut_{1}F_{2}\Big{(}\delta;\beta,\gamma;-t^{2}\Big{)}\right]$ (40)
[…]
$=M2\delta t^{2\kappa+d}\mathstrut_{1}F_{2}\left(\delta+1;\beta,\gamma;-t^{2}\right)$
$=M2\delta t^{2\kappa+d}\mathstrut_{1}F_{2}\left(\delta_{1};\beta_{1},\gamma_{1};-t^{2}\right),$
where $\delta_{1}=\frac{d+1}{2}+(k+1)$, $\beta_{1}=\delta_{1}+\frac{\mu-2}{2}$, $\gamma_{1}=\beta_{1}+l$, and $l=\frac{d}{2}+\kappa$.
Since $M2\delta t^{2\kappa+d}$ is positive, $\frac{dg(t)}{dt}>0$ if and only if $\mathstrut{}_{1}F_{2}\left(\delta_{1};\beta_{1},\gamma_{1}+l;-t^{2}\right)>0$ Using point 1 of Theorem 2, this happens if and only if $\mu\geq 4$.
Appendix B Closed-form expressions of the hypergeometric model
In Euclidean spaces of odd dimension. When $d$ is an odd integer, $\kappa=k$ is a non-negative integer and $\mu$ is a positive integer, the hypergeometric model yields a truncated polynomial (Emery and Alegría,, 2022, Section 2.5.3). Specifically, if $\mu$ is odd, then
$\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+1}{2}+k)}{\Gamma(\frac{1-\mu-d}{2}-2k)}\bigg{[}\sum_{n=0}^{k+\frac{\mu-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu-d}{2}-2k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+1}{2}+k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{k-1+\frac{\mu+d}{2}}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu}{2})}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu+d}{2}+k-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg{]}\text{ if $x<a$}\\ 0\text{ if $x\geq a$,}\end{cases}\end{split}$
while, if $\mu$ is even,
$\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+d+1}{2}+2k)}{\Gamma(\frac{1-\mu}{2}-k)}\bigg{[}\sum_{n=0}^{2k+\frac{\mu+d-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu}{2}-k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+d+1}{2}+2k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{\frac{\mu}{2}-1}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu+d}{2}-k)}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu}{2}-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg{]}\text{ if $x<a$}\\ 0\text{ if $x\geq a$.}\end{cases}\end{split}$
The cases $k=0$, $\mu=1$ and $d=1$, $3$ and $5$ correspond to the well-known triangular, spherical and pentaspherical models (Matérn,, 1986), which behave linearly near the origin:
$\mathcal{H}_{0,1,a,1}(x)=1-\frac{x}{a},\quad x\leq a,$
$\mathcal{H}_{0,1,a,3}(x)=1-\frac{3x}{2a}+\frac{x^{3}}{2a^{3}},\quad x\leq a,$
$\mathcal{H}_{0,1,a,5}(x)=1-\frac{15x}{8a}+\frac{5x^{3}}{4a^{3}}-\frac{3x^{5}}{8a^{5}},\quad x\leq a,$
whereas the cases $d=3$, $\mu=1$ and $k=1$ or $2$ correspond to the so-called cubic and penta models (Chilès and Delfiner,, 2012), which are associated with Gaussian RFs that are once and twice mean-squared differentiable, respectively:
$\mathcal{H}_{1,1,a,3}(x)=1-\frac{7x^{2}}{a^{2}}+\frac{35x^{3}}{4a^{3}}-\frac{7x^{5}}{2a^{5}}+\frac{3x^{7}}{4a^{7}},\quad x\leq a,$
$\mathcal{H}_{2,1,a,3}(x)=1-\frac{22x^{2}}{3a^{2}}+\frac{33x^{4}}{a^{4}}-\frac{77x^{5}}{2a^{5}}+\frac{33x^{7}}{2a^{7}}-\frac{11x^{9}}{2a^{9}}+\frac{5x^{11}}{6a^{11}},\quad x\leq a.$
In the 2D plane. If $d=2$ and $\mu=1$, the following closed-form expressions are found.
  1. 1.
    When $\kappa=0$, formula 7.3.2.92 of Prudnikov et al., (1990) gives:
    $\mathcal{H}_{0,1,a,2}(x)=\frac{2}{\pi}\left(\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\right),$ (41)
    which corresponds to the circular correlation model (Chilès and Delfiner,, 2012, 2.46), which behaves linearly near the origin.
  2. 2.
    When $\kappa=\frac{1}{2}$, formula 7.3.2.99 of Prudnikov et al., (1990) gives:
    $\mathcal{H}_{\frac{1}{2},1,a,2}(x)=\left(2-\frac{x^{2}}{a^{2}}\right)\operatorname{arctanh}\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}},$
    which is the montée (upgrading) (Matheron,, 1965, I.4.18) of order $1$ of the spherical correlation model in $\mathbb{R}^{3}$.
  3. 3.
    When $\kappa=1$, one obtains (Weisstein,, 2025):
    $\mathcal{H}_{1,1,a,2}(x)$ $=$ $\frac{2}{3\pi}\Bigg{\{}\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\left[15-4\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\left(3-\frac{x^{2}}{a^{2}}\right)\right]$ (42)
    $+3\left[6\left(1-\frac{x^{2}}{a^{2}}\right)_{+}-5\right]\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}\Bigg{\}},$
    which is the montée of order $2$ of the Euclid’s hat model in $\mathbb{R}^{4}$ and is associated with a Gaussian RF that is mean-squared differentiable.

### v2

@@
 Abstract
-We study the covariance model belonging to the Gauss hypergeometric ($\mathcal{GH}$) class, a highly flexible and compactly supported correlation model (kernel).
-This class includes the well-known Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) kernels as special cases.
-First, we provide necessary and sufficient conditions for the validity of the $\mathcal{GH}$ model.
-We then demonstrate that this family of models suffers from identifiability issues under both increasing and fixed-domain asymptotics.
-To address this problem, we propose a parsimonious version of the model that adheres to an optimality criterion.
-This approach results in a new class of compactly supported kernels, which can be seen as an improvement over the $\mathcal{GW}$ model.
-Additionally, we show that two specific compact support reparameterizations allow us to recover the $\mathcal{MT}$ model, highlighting the advantages and disadvantages of each reparameterization.
-[…]
+We study covariance functions in the Gauss hypergeometric ($\mathcal{GH}$) class, a flexible family that encompasses the Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) models.
+We derive sharp validity conditions, providing a complete characterization of the admissible parameter space, and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics.
+To resolve this issue, we introduce a parsimonious compactly supported subclass selected via a maximum integral range criterion.
+The resulting hypergeometric model can be viewed as a structural refinement of the $\mathcal{GW}$ family and admits compact-support reparameterizations that recover the $\mathcal{MT}$ model as a limit case.
+We further establish strong consistency and asymptotic normality of the maximum likelihood estimator of the associated microergodic parameter under fixed-domain asymptotics.
+Simulation experiments and a real-data application to climate data illustrate the finite-sample behavior and practical performance of the proposed model.
 Keywords: Fixed domain asymptotics, Gaussian random fields, Generalized Wendland model, Integral range, Sparse matrices.
 1 Introduction
-Gaussian processes, or Gaussian random fields (RFs), are widely used in various disciplines such as spatial statistics (Cressie,, 1993; Stein,, 1999; Banerjee et al.,, 2004), machine learning (Rasmussen and Williams,, 2006), numerical analysis (Wendland,, 2004), and computer experiments (Santner et al.,, 2003), among others.
-The Gaussian assumption implies that the finite-dimensional distributions of the RF are fully determined by the mean value and the covariance function.
-In particular, covariance functions (also known as kernels in the machine learning literature) are positive semidefinite functions that play a fundamental role in modeling relationships between data points in Gaussian RFs, kernel methods, and other machine learning algorithms.
-For RFs defined in Euclidean spaces, covariance functions that depend solely on the distance between any pair of points within the domain are called isotropic or radially symmetric.
-This paper operates under this assumption, which greatly simplifies the discussion.
-However, isotropic covariances serve as the building blocks for more complex models, such as anisotropic, nonstationary (Paciorek and Schervish,, 2006), multivariate (Gneiting et al.,, 2010), or spatiotemporal extensions (Porcu et al.,, 2020), among others.
-To date, there exists a rich catalog of parametric isotropic covariance functions (see, for instance, Rasmussen and Williams, (2006) or Chilès and Delfiner, (2012)), among which the Matérn ($\mathcal{MT}$ hereafter) model is by far the most widely used (see Porcu et al., (2024) for a comprehensive review).
-It has played a central role in spatial statistics for decades and, more recently, in other disciplines such as numerical analysis, approximation theory, and machine learning.
-A key advantage of the $\mathcal{MT}$ model is its continuous parameterization of the differentiability of the sample paths of the associated Gaussian RF through a positive smoothness parameter.
-The larger the smoothness parameter, the higher the level of differentiability of the sample paths.
-.
-While the $\mathcal{MT}$ model remains widely used, particularly in the stochastic partial differential equations approach and its generalizations to non-Euclidean domains (Lindgren et al.,, 2011; Bolin et al.,, 2024), enhancements to the $\mathcal{MT}$ model have been proposed in recent years.
-These covariance models retain the benefits of the Matérn model while offering additional features (Porcu et al.,, 2024).
-Among these, the generalized Wendland ($\mathcal{GW}$ hereafter) (Gneiting,, 2002; Bevilacqua et al.,, 2019) has gained popularity, as it allows for a continuous parameterization of the differentiability of the sample paths of the underlying Gaussian RF, similar to the $\mathcal{MT}$ model, while also being compactly supported.
-This is a desirable feature from a computational perspective, as it produces sparse covariance matrices (Duff et al.,, 1989; Davis,, 2006).
-Several connections between the $\mathcal{MT}$ and $\mathcal{GW}$ models have been established.
-For instance, Bevilacqua et al., (2019) show that, under specific conditions, these models lead to equivalent Gaussian measures.
-Additionally, Bevilacqua et al., (2022) demonstrate that the $\mathcal{MT}$ model is a special limit case of a compactly supported reparameterization of the $\mathcal{GW}$ model.
-Estimation and prediction using $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied under both increasing and fixed-domain asymptotics (Mardia and Marshall,, 1984; Zhang,, 2004; Kaufman and Shaby,, 2013; Bevilacqua et al.,, 2019).
-The focus of this paper is the study of the Gauss hypergeometric ($\mathcal{GH}$) covariance, a highly flexible kernel recently proposed by Emery and Alegría, (2022).
-These authors demonstrate that the $\mathcal{GH}$ model encompasses most well-known covariance models as special or asymptotic cases, including both the $\mathcal{MT}$ and $\mathcal{GW}$ models.
-In addition, they provide sufficient conditions for the validity of the model.
-In Porcu et al., (2024), the Gauss hypergeometric ($\mathcal{GH}$) kernel appears to be one of the most promising generalizations of the $\mathcal{MT}$ model.
-However, the estimation of the $\mathcal{GH}$ covariance model, particularly under fixed-domain asymptotics, has not yet been addressed.
-This paper presents several contributions related to the $\mathcal{GH}$ covariance model.
-First, we provide necessary and sufficient conditions for the validity of the $\mathcal{GH}$ model, offering a complete characterization of the kernel.
-This result ensures that the model is well-defined and applicable to a wide range of scenarios.
-Second, we show that this family of models suffers from identifiability issues under both increasing and fixed-domain asymptotics.
-These findings highlight a critical limitation of the $\mathcal{GH}$ model in practical applications, particularly when attempting to infer parameters from data.
-To address this problem, we propose a parsimonious version of the model that adheres to an optimality criterion.
-This criterion is based on the maximization of the isotropic integral range (Lantuéjoul,, 2002), which can be seen as a measure of the flexibility of the correlation model.
-For instance, if $d=1$, the maximum integral range corresponds to the maximum area under the unidimensional correlation function.
-Using this approach, we derive a new class of compactly supported kernels, which we call the hypergeometric ($\mathcal{H}$ hereafter) model that can be seen as an improvement over the $\mathcal{GW}$ model.
-The model can also be seen as a generalization of the Euclid’s hat model proposed in Gneiting, (1999), which includes the celebrated spherical models widely used in geostatistics (Chilès and Delfiner,, 2012).
-Additionally, we show that two specific compact support reparameterizations allow us to recover the $\mathcal{MT}$ model as a special limit case, highlighting the advantages and disadvantages of each reparameterization.
-Finally, for a fixed smoothness parameter, we establish strong consistency and the asymptotic distribution of the maximum likelihood (ML) estimator of the microergodic parameter associated with the $\mathcal{H}$ class.
-[…]
-The $\mathcal{H}$ model has been implemented in the GeoModels package (Bevilacqua et al.,, 2025) for the open-source R statistical environment.
-The remainder of this paper is organized as follows.
-Section 2 provides background material on the $\mathcal{MT}$ and $\mathcal{GW}$ covariance models.
-Section 3 introduces the $\mathcal{GH}$ model and highlights the identifiability issue.
-Section 4 first provides necessary and sufficient conditions for the $\mathcal{GH}$ model.
-Then, following the criterion of maximal integral range, the $\mathcal{H}$ model is proposed.
-In Section 5, some closed-form expressions for the $\mathcal{H}$ model that do not depend on the computation of special functions are presented.
-In addition, two compactly supported reparameterizations that include the $\mathcal{MT}$ model as a special case are introduced.
-In Section 6, we provide the asymptotic properties of the ML estimator of the microergodic parameter.
-In Section 7, we report a simulation study that explores the finite-sample properties of the ML method when estimating the parameters of the proposed (reparameterized) $\mathcal{H}$ model.
-In Section 8, we present the analysis of a climate data set.
-Concluding remarks are provided in Section 9.
-All proofs have been deferred to Appendix A.
+Gaussian random fields (RFs) are widely used in spatial statistics, machine learning, numerical analysis, and computer experiments (Cressie, 1993; Stein, 1999; Banerjee et al., 2004; Rasmussen and Williams, 2006; Wendland, 2004; Santner et al., 2003).
+Under Gaussianity, the dependence structure is fully characterized by the covariance function, whose specification is central for modeling, prediction, and uncertainty quantification.
+For RFs on Euclidean spaces, isotropic covariance (kernel) functions— depending only on Euclidean distance—provide a tractable and widely used modeling framework.
+Among parametric isotropic models, the Matérn ($\mathcal{MT}$) class is arguably the most prominent (Porcu et al., 2024), owing to its smoothness parameter, which continuously controls the mean-square differentiability of the sample paths.
+A compactly supported alternative is the generalized Wendland ($\mathcal{GW}$) family (Gneiting, 2002; Bevilacqua et al., 2019), which also allows for continuous control of smoothness while yielding sparse covariance matrices and substantial computational gains (Duff et al., 1989; Davis, 2006).
+Connections between $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied (Bevilacqua et al., 2019, 2022).
+We focus on the Gauss hypergeometric ($\mathcal{GH}$) covariance model introduced by Emery and Alegría (2022).
+This highly flexible class includes both $\mathcal{MT}$ and $\mathcal{GW}$ as special or asymptotic cases and has been recognized as a prominent extension of the Matérn family (Porcu et al., 2024).
+While sufficient validity conditions were derived in Emery and Alegría (2022), a complete characterization of the parameter space and a detailed inferential analysis remain open.
+This paper provides several theoretical and methodological contributions.
+First, we refine the validity conditions of the $\mathcal{GH}$ model.
+We show that the sufficient conditions of Emery and Alegría (2022) are also necessary in the parameter range where the two sets of conditions agree.
+Second, we clarify that the $\mathcal{GH}$ family is not identifiable in the classical parametric sense.
+The parameter-to-correlation mapping is not injective, and this structural non-identifiability must be distinguished from the well-known inferential limitations under fixed-domain asymptotics, where Gaussian measures may remain equivalent on bounded domains.
+In such settings, only microergodic parameter combinations are consistently estimable.
+To resolve the structural non-identifiability, we introduce an optimality-driven parsimonious subclass of $\mathcal{GH}$.
+For fixed smoothness and compact support, we select the member that maximizes the isotropic integral range, that is, the total mass of the correlation over $\mathbb{R}^{d}$.
+Since the integral range coincides with the spectral density at zero frequency, this criterion selects the admissible model with maximal low-frequency spectral power, i.e., the strongest large-scale dependence compatible with the imposed constraints.
+The integral range provides a global and threshold-free measure of dependence (Vanmarcke, 1983; Lantuéjoul, 2002; Hristopulos and Zukovic, 2011); in dimension $d=1$, it reduces to the area under the correlation function.
+The proposed selection principle is intrinsic to the model class and does not depend on sampling design or loss functions.
+The resulting hypergeometric ($\mathcal{H}$) model constitutes the main contribution of this paper.
+It provides a canonical and parsimonious representative of the $\mathcal{GH}$ class, preserving compact support, continuous smoothness parameterization, and modeling flexibility.
+Notably, the $\mathcal{H}$ model satisfies simple and explicit necessary and sufficient validity conditions, reducing to a single lower bound on the shape parameter that is independent of the spatial dimension.
+In addition, it admits a beta-type scale-mixture representation and may be interpreted as a generalization of the Euclid’s hat model (Gneiting, 1999), encompassing the classical spherical covariance functions widely used in geostatistics (Chilès and Delfiner, 2012).
+We further show that a compact-support reparameterization of the $\mathcal{H}$ model yields the Matérn model as a special limit case, thereby extending to the hypergeometric framework the Wendland-to-Matérn connection proved in Bevilacqua et al.
+(2022).
+A simulation study investigates finite-sample behavior under increasing- and fixed-domain scenarios, and a real-data analysis of climate data illustrates that the proposed $\mathcal{H}$ model can outperform the $\mathcal{MT}$ model in terms of model fit, prediction accuracy, and computational efficiency for both Gaussian and Tukey-$h$ RFs (Xu and Genton, 2017).
+The $\mathcal{H}$ model is implemented in the GeoModels package (Bevilacqua et al., 2026).
+The remainder of the paper is organized as follows.
+Section 2 reviews the $\mathcal{MT}$ and $\mathcal{GW}$ models.
+Section 3 introduces the $\mathcal{GH}$ model and discusses identifiability.
+Section 4 provides validity conditions and derives the $\mathcal{H}$ subclass.
+Section 5 presents closed-form expressions and compact-support reparameterizations.
+Section 6 establishes asymptotic properties of the microergodic estimator.
+Sections 7 and 8 report simulation and real-data results.
+All proofs are collected in Appendix A.
 2 Background: the Matérn and Generalized Wendland correlation models
@@
 ${\rm corr}\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),$
-[…]
-Schoenberg, (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and non-negative measure, $F$:
+with $\boldsymbol{h}=\boldsymbol{s}^{\prime}-\boldsymbol{s}$ and $\boldsymbol{s},\boldsymbol{s}^{\prime}\in D$, and $\|\cdot\|$ denoting the Euclidean norm.
+Gaussian RFs with such covariance functions are called weakly stationary and isotropic.
+Schoenberg (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and nonnegative measure, $F$:
 $\phi(x)=\int_{0}^{\infty}\Omega_{d}(xr)F(\text{d}r),\qquad x\geq 0,$
@@
 The class $\Phi_{d}$ is nested, with the inclusion relation $\Phi_{1}\supset\Phi_{2}\supset\ldots\supset\Phi_{\infty}$ being strict, and where $\Phi_{\infty}:=\bigcap_{d\geq 1}\Phi_{d}$ is the class of continuous mappings $\phi$, the radial version of which is positive semidefinite on any $d$-dimensional Euclidean space.
-Fourier transforms of radial versions of members of $\Phi_{d}$, for a given $d$, have a simple expression, as reported in Yaglom, (1987) or Stein, (1999).
+Fourier transforms of radial versions of members of $\Phi_{d}$, for a given $d$, have a simple expression, as reported in Yaglom (1987) or Stein (1999).
 For a member $\phi$ of the class $\Phi_{d}$ such that $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$, we define its isotropic spectral density as
@@
     The Matérn ($\cal MT$}) family
-    The Matérn model is a two-parameter globally supported correlation function defined as (Matérn,, 1986):
-    ${\cal MT}_{\nu,\alpha}(x)=\begin{cases}1,\quad x=0\\ \frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,\end{cases}$ (2)
-    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun,, 1970, 9.6.22).
+    The Matérn model is a two-parameter globally supported correlation function defined as (Matérn, 1986):
+    ${\cal MT}_{\nu,\alpha}(x)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,$ (2)
+    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun, 1970, 9.6.22).
 The associated spectral density, $\widehat{{\cal MT}}_{\alpha,\nu}$, is given by
     $\widehat{{\cal MT}}_{\nu,\alpha}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{\alpha^{d}}{(1+\alpha^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
-    Arguments in Stein, (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
-    The importance of the Matérn model lies in the parameter $\nu$, which controls the differentiability (in the mean square sense) of the associated Gaussian RF.
-Specifically, for any integer $l=0,1,\ldots$, a Gaussian RF with a Matérn covariance function is $l$-times mean square differentiable, in any direction, if and only if $\nu>l$.
-When $\nu=l+1/2$, the Matérn covariance function simplifies into the product of an exponential covariance and a polynomial of order $l$, given by:
-    ${\cal MT}_{l+1/2,\alpha}(x)=\exp\left(-\frac{x}{\alpha}\right)\sum_{i=0}^{l}\frac{(l+i)!}{2l!}\binom{l}{i}\left(\frac{2x}{\alpha}\right)^{l-i},\quad l=0,1,\ldots$
+    Arguments in Stein (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
+    A key feature of the Matérn model is the smoothness parameter $\nu$, which determines the mean-square differentiability of the sample paths of the associated Gaussian RF.
+In particular, for any integer $l=0,1,\ldots$, the sample paths are $l$-times mean-square differentiable (in any direction) if and only if $\nu>l$.
+In the special case $\nu=l+\tfrac{1}{2}$, the correlation function reduces to an exponential term multiplied by a polynomial of degree $l$.
   2. 2.
     The Generalized Wendland ($\cal GW$}) family
-    The ${\cal GW}$ model is a three-parameters compactly supported correlation function defined as (Gneiting,, 2002; Bevilacqua et al.,, 2019):
+    The $\cal GW$} model is a three-parameters compactly supported correlation function defined as (Gneiting, 2002; Bevilacqua et al., 2019):
     $\begin{split}{\cal GW}_{\kappa,\mu,\beta}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{\beta^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{\beta^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
-    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and ${}_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun,, 1970, 15.1.1).
+    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and $_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun, 1970, 15.1.1).
 The associated spectral density is given by:
@@
     where $_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\nu+1)\Gamma(2\nu+d)\Gamma(\nu)}{\Gamma\left(\nu+{d}/{2}\right)\Gamma(\mu+2\nu+d+1)\Gamma(2\nu)}$ is a normalization constant.
-    For a given $\kappa>-\frac{1}{2}$, a necessary and sufficient condition for ${\cal GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ is (Bevilacqua et al.,, 2024)
-    $\mu\geq\mu_{\min}(\kappa,d):=\begin{cases}\frac{\sqrt{8\kappa+9}-1}{2}\text{ if $d=1$ and $-\frac{1}{2}<\kappa<0$}\\ \kappa+\frac{d+1}{2}\text{ otherwise.}\end{cases}$ (6)
+    For a given $\kappa>-\frac{1}{2}$, the validity conditions for $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ follow as a special case of Theorem 2 by setting $l=\frac{1}{2}$:
+    • (A)
+      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ if and only if $\mu\geq\kappa+\frac{d+1}{2}$, whenever $\kappa\geq 0$ or $d\geq 2$;
+    • (B)
+      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{1}$ if $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$, and only if $\mu\geq\kappa+1$, for $\frac{1}{2}<\kappa<0$.
+    Since $\kappa+1<\frac{\sqrt{8\kappa+9}-1}{2}$ for all $\frac{1}{2}<\kappa<0$, a gap exists between the necessary and sufficient conditions.
+This partially corrects Bevilacqua et al.
+(2024), which claimed in case (B) that $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$ is both necessary and sufficient for $d=1$.
+Whether this lower bound is also necessary remains an open problem.
     The ${\cal GW}_{\kappa,\mu,\beta}$ model, as in the ${\cal MT}_{\nu,\alpha}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$.
 Specifically, for any integer $l=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,\beta}$ model are $l$ times differentiable, in any direction, if and only if $\kappa>l-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
-    When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$:
-    ${\cal GW}_{k,\mu,\beta}=\left(1-\frac{x}{\beta}\right)^{\mu+k}_{+}P_{k}(x;\mu,\beta),\quad k=0,1,2,\ldots$ (7)
-    where $P_{k}(h,\mu,\beta)$ is a polynomial of degree $k$ that can be expressed as:
-    $P_{k}(x;\mu,\beta)=\frac{2^{k}k!}{(2k)!}\sum_{j=0}^{k}c_{j,k}(\mu)\left(\frac{x}{\beta}\right)^{k-j}\left(1-\frac{x}{\beta}\right)^{j},$ (8)
-    and $c_{j,k}(\mu)$ are suitable constants (Bevilacqua et al.,, 2024).
-    Recently, Bevilacqua et al., (2022) considered a compact support reparameterized version of the ${\cal GW}$ model, and proved that
+    When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$.
+    Recently, Bevilacqua et al. (2022) considered a compact support reparameterized version of the $\cal GW$} model, and proved that
     $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(\beta,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (6)
-    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $[0,\infty)$.
-    A simpler compact support reparameterization that leads to the ${\cal MT}$ model as a special limit case has been proposed in Emery et al., (2025):
-    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,\mu\beta}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (10)
-    again with uniform convergence on $[0,\infty)$.
-    As a consequence, the two reparameterized ${\cal GW}$ models can be viewed as a generalization of the ${\cal MT}$ model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
+    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $0,\infty)$.
+    As a consequence, the reparameterized $\cal GW$} model can be viewed as a generalization of the $\cal MT$} model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
 3 The Gauss hypergeometric model and the identifiability problem
-The $\mathcal{GH}$ class in Emery and Alegría, (2022) has been defined starting from the following spectral density:
+The $\mathcal{GH}$ class in Emery and Alegría (2022) has been defined starting from the following spectral density:
 $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}\mathstrut_{1}F_{2}\Big(\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\Big),\quad z\geq 0.$ (7)
@@
 However, in (7) there is no dependence between the parameters of the $\mathstrut{}_{1}F_{2}$ function as in the $\cal GW$} case.
-Using some results in Cho and Yun, (2018), Emery and Alegría, (2022) show that, for a given $a>0$, the following conditions:
+Using some results in Cho and Yun (2018), Emery and Alegría (2022) show that, for a given $a>0$, the following conditions:
   1. (a)
@@
     $\beta+\gamma\geq 3\delta+\frac{1}{2}$;
-are sufficient for the positivity of the spectral density on $[0,\infty)$. The associated four-parameter isotropic correlation model is given by:
+are sufficient for the positivity of the spectral density on $0,\infty)$. The associated four-parameter isotropic correlation model is given by:
 $\begin{split}\mathcal{GH}_{\delta,\beta,\gamma,a}(x)&=\frac{\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})}{\Gamma(\beta-\delta+\gamma-\frac{d}{2})\Gamma(\delta-\frac{d}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\beta-\delta+\gamma-\frac{d}{2}-1}\\ &\times{}_{2}F_{1}\left(\beta-\delta;\gamma-\delta;\beta-\delta+\gamma-\frac{d}{2};\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (8)
@@
 As in the $\cal MT$} and $\cal GW$} cases, the $\cal G$H} model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\delta$.
-Furthermore, it is not difficult to see that the ${\cal GH}$ model includes the ${\cal GW}$ and the ${\cal MT}$ models as special cases.
-For instance,
+Furthermore, it is not difficult to see that the $\cal G$H} model includes the $\cal GW$} and the $\cal MT$} models as special cases. For instance,
 $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+\frac{1}{2},a}(x)={\cal GW}_{\kappa,\mu,a}(x),$ (9)
 where $\delta=\frac{d+1}{2}+\kappa$.
-Although the ${\cal GH}$ model is highly flexible, it suffers from identifiability problems.
-This is apparent under an increasing domain scenario, where all the parameters can be estimated, since $\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a}$ due to the symmetry of the Gauss hypergeometric function: ${}_{2}F_{1}\left(a;b;\cdot;x\right)={}_{2}F_{1}\left(b;a;\cdot;x\right)$.
-This implies that there is not a one-to-one correspondence between the parametric space and the correlation function.
-To study identifiability under fixed-domain asymptotics, the following theorem establishes the equivalence condition for two distinct Gaussian measures with $\mathcal{GH}$ correlation models that share a common smoothness parameter.
+Although the $\cal G$H} family is highly flexible, it is not identifiable in the usual parametric sense.
+Indeed, even under increasing-domain asymptotics—where one would in principle expect all parameters to be estimable—the parameter-to-correlation map is not injective.
+In particular,
+$\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a},$
+because the Gauss hypergeometric function is symmetric in its first two arguments:
+$_{2}F_{1}(a,b;\cdot\,;x)={}_{2}F_{1}(b,a;\cdot\,;x).$
+Hence distinct parameter values (obtained by swapping $\beta$ and $\gamma$) induce exactly the same correlation function. This implies that additional constraints are required.
+A related inferential limitation arises under fixed-domain (infill) asymptotics, where distinct parameter values may induce equivalent Gaussian measures.
+The following theorem establishes the equivalence condition for two distinct Gaussian measures with $\mathcal{GH}$ correlation models that share a common smoothness parameter.
 Hereafter, we denote by $P(\rho_{i})$, $i=0,1$, a zero-mean Gaussian measure with an isotropic covariance function $\rho_{i}$.
 Theorem 1.
-Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures.
-If $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$, for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$, if and only if
+Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures and let $\delta>\frac{d}{2}$.
+If $\frac{d+1}{2}+3\delta<\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0}),$ for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, if and only if
 $\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}},$ (10)
-where $L(\beta_{i},\gamma_{i})=\frac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$ with $i=0,1.$
-An immediate consequence of Theorem 1 is that the microergodic parameter of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model is given by $\frac{\sigma^{2}L(\beta,\gamma)}{a^{2\delta-d}}$.
-Since $L(\beta,\gamma)=L(\gamma,\beta)$, it is clear that there is not a unique correspondence between the parametric space and the microergodic parameter, which implies that the microergodic parameter is not identifiable.
-Summarizing our results, the $\mathcal{GH}$ correlation model is highly flexible, including the $\mathcal{GW}$ and the $\mathcal{MT}$ models as special cases.
-However, it is not identifiable under both increasing and fixed-domain scenarios.
-To fix this issue, in the next section, we will assume that the second shape parameter ($\gamma$) is a function of the first shape parameter ($\beta$).
+where $L(\beta_{i},\gamma_{i})=\dfrac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$, $i=0,1$.
+The restriction to $d=1,2,3$ is typical when studying equivalence of Gaussian measures for isotropic models on $\mathbb{R}^{d}$, and is not specific to the $\cal G$H} class.
+Indeed, the same dimensional restriction appears in the equivalence results for the $\cal MT$} model (Zhang, 2004), and stems from the integrability conditions required by the criteria of Stein (2004).
+An immediate consequence of Theorem 1 is that, for fixed $\delta$, the microergodic parameter is given by
+$\frac{\sigma^{2}L(\beta,\gamma)}{a^{2\delta-d}}.$
+Because $L(\beta,\gamma)=L(\gamma,\beta)$, this quantity does not distinguish $(\beta,\gamma)$ from $(\gamma,\beta)$.
+Thus, while the microergodic parameter is the relevant identifiable feature under fixed-domain asymptotics, the individual shape parameters $\beta$ and $\gamma$ are not identifiable without an additional constraint.
+Summarizing, the $\mathcal{GH}$ correlation model is highly flexible and includes the $\mathcal{GW}$ and $\mathcal{MT}$ models as special cases.
+However, it is not identifiable in the usual parametric sense because swapping $(\beta,\gamma)$ leaves the correlation unchanged.
+This issue can be observed under both increasing and fixed domain scenario.
+To this end, in the next Section we assume that the second shape parameter ($\gamma$) is a function of the first ($\beta$).
 4 Fixing the identifiability problem: the Hypergeometric correlation model
-In this Section we first provide necessary and sufficient conditions for the validity of the ${\cal GH}$ model.
-Then we solve the identifiability problem by considering a parsimonious version that responds to an optimality criterion.
-With this goal in mind, we consider a ${\cal GW}$-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model, that is, we set
-$\delta=\frac{d+1}{2}+\kappa,\quad\beta=\delta+\frac{\mu}{2},\quad\gamma=\beta+l,$ (15)
-with $l\geq 0$. As outlined in (13), this parameterization includes the ${\cal GW}$ model as a special case when $l=\frac{1}{2}$.
-The following theorem can be viewed as an improvement of Theorem 1 in Emery and Alegría, (2022) and provides necessary and sufficient conditions for the validity of the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model.
+In this section we first refine the validity conditions of the $\mathcal{GH}$ model, clarifying the relationship between necessary and sufficient constraints.
+We then address the identifiability issue discussed in the previous section by introducing a parsimonious subclass selected according to an explicit optimality criterion.
+To this end, we adopt a $\cal GW$}-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model and set
+$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\beta+l.$ (11)
+As noted in (9), this parametrization includes the $\cal GW$} model as the special case $l=\frac{1}{2}$.
+The following theorem improves Theorem 1 in Emery and Alegría (2022) by providing sharp validity conditions for $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
 Theorem 2.
-Let $d$ be a fixed dimension, $\delta=\frac{d+1}{2}+\kappa$ and $\kappa>-\frac{1}{2}$ a smoothness parameter.
-Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ if and only if
+Let $d$ be fixed, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $a>0$, and $l\geq 0$. Then:
   1. 1.
-    $\mu\geq\delta-l+\frac{1}{2}\quad$ if $\quad 0\leq l\leq\frac{d}{2}+\kappa$
+    If $0\leq l\leq\frac{d}{2}+\kappa$, then
+    $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq\delta-l+\frac{1}{2}.$
   2. 2.
-    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>\frac{d}{2}+\kappa$.
-It should be stressed that necessary and sufficient conditions can also be found for $l<0$.
-They are symmetric with respect to the case $l\geq 0$ (see the proof of Theorem 2 in Appendix A).
-However, to solve the issue of identifiability, hereafter we consider the condition $l\geq 0$ as in Theorem 2.
-If $l=\frac{1}{2}$ then the sufficient and necessary conditions of the ${\cal GW}_{\nu,\mu,a}$ model in (6) are obtained as special case.
-It is apparent that, to solve the identifiability problem, $l$ should be fixed (as in the ${\cal GW}$ model) or should depend on the other parameters and/or the dimension, so an interesting question is: which is the “best” $l\geq 0$ in the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ family?
-This question can be responded by establishing an optimality criterion.
-Here we consider the criterion of maximum integral range that can be attained by a compactly supported correlation for a given smoothness and compact support.
-For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is defined as
-$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\text{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{\mathbb{R}^{+}}x^{d-1}\phi(x)\text{d}x.$ (16)
-$A$ is the isotropic version of the integral range proposed in Lantuéjoul, (1991) for stationary correlation functions.
-This measure has a close relation with the ergodicity of the RF, with the definition of an effective number of independent observations, and with the definition of long-range dependence (Yaglom,, 1987; Matheron,, 1989).
-In particular, following Lim and Teo, (2009), a RF with isotropic correlation $\phi$ is long-range dependent if $A=\infty$ and short-range dependent when $A<\infty$.
-If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$.
-Hristopoulos, (2020) interprets $A^{1/d}$ as an average measure of the distance over which two RF values are correlated.
-When $d=1$, the integral range corresponds to the area under the unidimensional correlation function.
-A parametric correlation model that can cover a wide range of areas, from 00 to a specific upper bound, can be considered a flexible model (see Figure  1 and Theorem 6 hereunder).
-As a consequence, this can serve as a criterion for selecting between two different parametric correlation models, particularly when comparing two compactly supported correlation models that share the same smoothness and compact support parameters.
-In such cases, the correlation model that achieves a larger integral range is preferred.
-For a parametric correlation model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we will denote its associated integral range as $A_{\boldsymbol{\theta},d}$.
-For instance, using formula 6.561.16 of Gradshteyn and Ryzhik, (2007), for the ${\cal MT}_{\nu,\alpha}$ correlation one has:
+    If $l>\frac{d}{2}+\kappa$, then
+    $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if}\quad\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l,$
+    and only if
+    $\mu\geq\delta-l+\frac{1}{2}.$
+For $l=\frac{1}{2}$, the necessary and sufficient validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ discussed in Section 2 are recovered as a special case.
+Specifically, since $l=\frac{1}{2}\leq\frac{d}{2}+\kappa$ if and only if $\kappa\geq\frac{1}{2}-\frac{d}{2}$, which holds for all $d\geq 2$ and $\kappa>-\frac{1}{2}$, or for $d=1$ and $\kappa\geq 0$, Regime 1 of Theorem 2 gives the necessary and sufficient condition
+$\mu\geq\frac{d+1}{2}+\kappa,$
+For $d=1$ and $\frac{1}{2}<\kappa<0$, instead $l=\frac{1}{2}>\frac{d}{2}+\kappa$, so Regime 2 applies, yielding the sufficient condition $\mu\geq 0.5(\sqrt{8\kappa+9}-1)$ and the necessary condition $\mu\geq\kappa+1$, consistently with the discussion in Section 2.
+As discussed in Section 3, the model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under the interchange $\beta\leftrightarrow\gamma$, corresponding to $l\mapsto-l$ in the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
+The admissible region is therefore symmetric with respect to $l=0$, and restricting attention to $l\geq 0$ entails no loss of generality.
+Translating Theorem 2 back to the original parameterization $\mathcal{GH}_{\delta,\beta,\gamma,a}$, the following equivalent conditions hold for $\delta>d/2$ and $a>0$:
+  1. 1.
+    If $0\leq\gamma-\beta\leq\delta-\frac{1}{2}$, then
+    $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if and only if}\quad\beta+\gamma\geq 3\delta+\frac{1}{2}.$
+  2. 2.
+    If $\gamma-\beta>\delta-\frac{1}{2}$, then
+    $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if}\quad\beta+\gamma\geq 2\delta+\sqrt{(\gamma-\beta)^{2}+2\delta},$
+    and only if
+    $\beta+\gamma\geq 3\delta+\frac{1}{2}.$
+In particular, in Regime 1 ($0\leq\gamma-\beta\leq\delta-\frac{1}{2}$), condition (c) of Section 3 implies condition (b), so that (a) and (c) alone characterize validity.
+Hence, within this regime, the sufficient conditions of Emery and Alegría (2022) reduce to (a) and (c) and are therefore also necessary.
+In Regime 2 ($\gamma-\beta>\delta-\frac{1}{2}$), whether the gap between sufficiency and necessity can be closed remains an open question.
+We now turn to the structural identifiability issue for the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model.
+To resolve it, the parameter $l$ must either be fixed (as in the $\cal GW$} model) or linked to the remaining parameters and/or to the spatial dimension.
+This naturally raises the question: which choice of $l\geq 0$ is optimal ?
+We address this question by introducing an explicit optimality criterion that selects a single parsimonious representative once smoothness and compact support are fixed.
+We adopt the criterion of maximum integral range attainable by a compactly supported correlation with given smoothness and support.
+For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is
+$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\,\mathrm{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{0}^{\infty}x^{d-1}\phi(x)\,\mathrm{d}x.$ (12)
+This quantity, originally introduced by Lantuéjoul (1991), provides a global and threshold-free measure of dependence (Yaglom, 1987; Matheron, 1989; Hristopulos and Zukovic, 2011).
+If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$, so that the integral range coincides with the spectral density at zero frequency.
+When $d=1$, AA reduces to the area under the correlation function.
+For compactly supported correlations, maximizing $A$ therefore selects, among all admissible models with fixed smoothness and support, the one with maximal low-frequency spectral power, that is, the strongest large-scale dependence compatible with the imposed constraints.
+While alternative optimality principles could be considered, they would typically depend on the sampling design or on a specific loss function.
+In contrast, the integral range criterion is intrinsic to the model class and yields a reproducible, one-to-one specification within the family.
+For a parametric model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we denote the associated integral range by $A_{\boldsymbol{\theta},d}$. For example, for the Matérn correlation one has
 $A_{\nu,\alpha,d}=\frac{\alpha^{d}\pi^{d/2}\Gamma\!\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (13)
-The following theorem provides an expression for the integral range associated with the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ correlation model.
+The following theorem provides the integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
 Theorem 3.
@@
 Theorem 4.
-For an arbitrary $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ is attained when
+Let $\kappa>0$, $a>0$, and $d\in\mathbb{N}$ be fixed.
+For each $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ with respect to $\mu$ (over its admissible parameter space) is attained at $\mu^{\star}(l)$, where:
   1. 1.
-    $\mu=\mu_{1}(l):=\delta-l+\frac{1}{2}\quad$ if $\quad 0\leq l\leq\frac{d}{2}+\kappa$
+    if $0\leq l\leq\dfrac{d}{2}+\kappa$, then
+    $\mu^{\star}(l)=\mu_{1}(l):=\delta-l+\frac{1}{2};$
   2. 2.
-    $\mu=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>\frac{d}{2}+\kappa$.
+    if $l>\dfrac{d}{2}+\kappa$, then
+    $\mu^{\star}(l)=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l.$
 For instance, let $\kappa\geq 0$.
-If $l=\frac{1}{2}$, then from (13) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of ${\cal GW}$ model (6).
+If $l=\frac{1}{2}$, then from (9) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of $\cal GW$} model in this case.
 As an example, Figure 1 (left part) depicts the ${\cal GW}_{0,\mu,1}$ model and the right part depicts the ${\cal GW}_{1,\mu,1}$ model for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$.
@@
 Theorem 5.
-The maximum of $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$ is attained when $l=d/2+\kappa$.
+Let $\kappa>-1/2$, $a>0$, and $d\in\mathbb{N}$ be fixed, and set $\delta=\frac{d}{2}+\kappa$.
+  1. 1.
+    The function $l\mapsto A_{\kappa,\mu_{1}(l),l,a,d}$ is non-decreasing on $0,\delta]$, hence
+    $\max_{0\leq l\leq\delta}A_{\kappa,\mu_{1}(l),l,a,d}=A_{\kappa,\mu_{1}(\delta),\delta,a,d}=A_{\kappa,1,\delta,a,d}.$
+  2. 2.
+    The function $l\mapsto A_{\kappa,\mu_{2}(l),l,a,d}$ is non-increasing on $(\delta,\infty)$ and
+    $\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}=A_{\kappa,1,\delta,a,d}.$
+In particular, the global maximum over all admissible $(\mu,l)$ is attained at $(\mu,l)=(1,\delta)$ and is equal to $A_{\kappa,1,\delta,a,d}$.
 Refer to caption Refer to caption
@@
 $0\leq A_{\kappa,\mu,l,a,d}\leq a^{d}\left[\frac{\pi^{d/2}\,\Gamma\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\left(1+\kappa\right)\,\Gamma\left(1+\frac{d}{2}+2\kappa\right)}{2^{d}\,\Gamma\left(\kappa+\frac{1}{2}\right)\,\Gamma\left(1+\frac{d}{2}+\kappa\right)\,\Gamma\left(1+d+2\kappa\right)}\right].$
-Accounting for (15), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting,, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría,, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron, (1965, I.4.18).
-Hereafter, we call the correlation model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ as hypergeometric correlation and we denote as
-$\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\quad x\geq 0,$ (20)
-which is defined as
-$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma(\kappa+\frac{\mu+1}{2})\Gamma(2\kappa+\frac{d+\mu+1}{2})}{\Gamma(\mu+\frac{d+1}{2}+2\kappa)\Gamma(\kappa+\frac{1}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2};\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (21)
-Necessary and sufficient validity conditions for $\mathcal{H}_{\kappa,\mu,a,d}$ can be easily obtained from Theorem 2 by setting $l=\frac{d}{2}+\kappa$.
-Specifically, given the smoothness parameter $\kappa>-\frac{1}{2}$, $\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}$ if and only if $\mu\geq 1$.
-Interestingly, the conditions are much simpler than the conditions (6) of the $\mathcal{GW}_{\kappa,\mu,a}$ model.
-In particular, the condition on $\mu$ does not depend on the dimension.
-The price to pay is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ depends on the dimension $d$.
-The following theorem proves that the ${\cal H}_{\kappa,1,a,d}$ subclass of hypergeometric correlation functions with maximum integral range allows recovering any member of the ${\cal H}_{\kappa,\mu,a,d}$ class, as well as great part of the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ class, by scale mixtures.
-The chosen optimality criterion (maximizing the integral range) therefore coincides with the greatest flexibility from a modeling viewpoint.
+Accounting for (11), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron (1965, I.4.18).
+Fixing $l=\frac{d}{2}+\kappa$, we define the hypergeometric correlation by
+$\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\qquad x\geq 0.$ (16)
+An explicit expression is given by
+$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)}{\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\frac{1}{2}\right)}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\[4.0pt] &\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\qquad x\geq 0.\end{split}$ (17)
+Notably, the model $\mathcal{H}_{\kappa,\mu,a,d}$ always lies in Regime 1 of Theorem 2, where validity is completely characterized.
+Indeed, by setting $l=\frac{d}{2}+\kappa$ in Theorem 2, we obtain the condition
+$\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq 1.$
+These validity conditions are substantially simpler than those for the $\mathcal{GW}_{\kappa,\mu,a}$ model discussed in Section 2.
+In particular, the lower bound on $\mu$ is independent of the spatial dimension.
+The trade-off is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ explicitly depends on $d$.
+The next theorem shows that the subclass ${\cal H}_{\kappa,1,a,d}$, obtained by maximizing the integral range, retains substantial modeling flexibility.
+Through beta-type scale mixtures, it can generate any member of ${\cal H}_{\kappa,\mu,a,d}$ and a broad subclass of the Gauss–hypergeometric family $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
+Thus, the maximal integral range criterion does not restrict modeling capacity, but instead selects a parsimonious representative that remains rich within the admissible class.
 Theorem 6.
-Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$ and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ owing to Theorem 2.
-Then, $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is a beta-type scale mixture of hypergeometric correlation functions of the form $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
-$\begin{split}&\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)=\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)}{\Gamma(\delta+\frac{1-d}{2})\Gamma(\delta+\frac{1}{2}+\kappa)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x){\rm d}u{\rm d}v,\quad x\geq 0.\end{split}$ (22)
-In particular, for $l=\frac{d}{2}+\kappa$, one obtains
-$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu}{2}+\kappa)}{\Gamma(\delta+\frac{1-d}{2})\Gamma(\delta+\frac{1}{2}+\kappa)\Gamma^{2}(\frac{\mu-1}{2})}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x){\rm d}u{\rm d}v,\quad x\geq 0.\end{split}$
+Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$, and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ by Theorem 2.
+Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ admits a beta-type scale mixture representation in terms of hypergeometric correlations $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
+$\begin{split}\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu-d}{2}+l\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu-1}{2}\right)\Gamma\!\left(\frac{\mu-1-d}{2}-\kappa+l\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$ (18)
+In particular, when $l=\frac{d}{2}+\kappa$, one obtains
+$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu}{2}+\kappa\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma^{2}\!\left(\frac{\mu-1}{2}\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$
 Clearly, the ${\cal H}_{\kappa,\mu,a,d}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$ in the same way as the ${\cal GW}_{\kappa,\mu,\beta}$ model.
@@
 5 Some special cases and bridges to the Matérn model
-The computation of the $\mathcal{H}_{\kappa,\mu,a,d}$ correlation function requires an efficient evaluation of the Gauss hypergeometric function, which is feasible (Johansson,, 2019) but could be cumbersome to (spatial) statisticians or machine learning practitioners used to handle closed-form parametric correlation models.
-However, similarly to the $\mathcal{MT}$ and $\mathcal{GW}$ models, in some special cases, the computation can be considerably simplified.
-The following theorem considers the case when $\kappa$ is a non-negative integer, $d=1$ and $\mu\geq 1$.
+The computation of the correlation function $\mathcal{H}_{\kappa,\mu,a,d}$ requires the evaluation of the Gauss hypergeometric function.
+Although efficient numerical implementations are available (Johansson, 2019), this may be less familiar to spatial statisticians and machine learning practitioners accustomed to closed-form parametric correlation models.
+As in the $\mathcal{MT}$ and $\mathcal{GW}$ families, substantial simplifications arise in certain special cases.
+The following theorem considers the case $\kappa=k\in\mathbb{N}_{0}$, $d=1$, and $\mu\geq 1$.
 Theorem 7.
-If $\kappa=k$ is a non-negative integer, $d=1$ and $\mu\geq 1$, then:
-$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\Gamma\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\left(k+0.5\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\times\sum\limits^{k}_{n=0}\frac{(n+k)!}{2^{2n}n!\Gamma(n+\frac{\mu}{2}+k+1)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
-Unfortunately, general closed forms for any $d>1$, $\mu\geq 1$ and $\kappa$ a non-negative integer, are difficult to obtain.
-Some special cases can, however, be obtained using closed forms of the Gauss hypergeometric function.
-For instance, the following theorem provides a closed form expression when $d$ is an odd integer, $\mu\geq 1$ and $\kappa=0$.
+If $\kappa=k$ is a nonnegative integer, $d=1$, and $\mu\geq 1$, then
+$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\!\left(k+\tfrac{1}{2}\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\quad\times\sum_{n=0}^{k}\frac{(n+k)!}{2^{2n}n!\,\Gamma\!\left(n+\frac{\mu}{2}+k+1\right)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
+General closed forms for arbitrary $d>1$, $\mu\geq 1$, and integer $\kappa$ are more difficult to obtain.
+Nevertheless, additional simplifications can be derived using known identities of the Gauss hypergeometric function.
+For instance, when $d$ is odd, $\mu\geq 1$, and $\kappa=0$, the following closed-form expression holds.
 Theorem 8.
-Let $d$ an odd integer and $\mu\geq 1$. If $\kappa=0$ then
-$\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum\limits^{\frac{d-1}{2}}_{n=0}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d-1}{2}+1\right)_{n}\left(1-\frac{x}{a}\right)^{n}_{+}}{2^{n}n!\Gamma\left(n+\frac{d-1}{2}+\mu+1\right)},$
-with $C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\left(\frac{\mu+1}{2}\right)\Gamma\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}$.
-Additional closed-form expressions, including the spherical models, for some fixed $\mu$ and $\kappa$ are given in Appendix B.
-The ${\cal MT}$ model can be obtained as a special limit case of the proposed $\mathcal{H}_{\kappa,\mu,a,d}$ model after two suitable compact support reparameterizations.
-To obtain the first parameterization, the following theorem provides sufficient conditions of equivalence between two Gaussian measures with covariances $\sigma_{0}^{2}{\cal MT}_{\nu,\alpha}$ and $\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d}$, respectively.
-The proof can be found in Appendix A.
+Let $d$ be an odd integer and $\mu\geq 1$. If $\kappa=0$, then
+$\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum_{n=0}^{\frac{d-1}{2}}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d+1}{2}\right)_{n}\left(1-\frac{x}{a}\right)_{+}^{n}}{2^{n}n!\,\Gamma\!\left(n+\frac{d-1}{2}+\mu+1\right)},$
+where
+$C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\!\left(\frac{\mu+1}{2}\right)\Gamma\!\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}.$
+Additional closed-form expressions, including spherical models, are reported in Appendix B.
+The Matérn model can be recovered as a limiting case of $\mathcal{H}_{\kappa,\mu,a,d}$ under an appropriate compact-support reparameterization.
+To establish this result, we first characterize the equivalence between the Gaussian measures $P(\sigma_{0}^{2}\mathcal{MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$.
+The proof is provided in Appendix A.
 Theorem 9.
-Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures. If $\nu=\kappa+\frac{1}{2}$ with $\nu>-\frac{1}{2}$ and
+Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures, and let $\kappa>-\frac{1}{2}$ and $\nu=\kappa+\frac{1}{2}$.
+For any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ are equivalent on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$, if and only if
 $\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}\,B(\kappa,\mu,d),$ (19)
-then for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$ on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$, where
-$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma(\frac{\mu+1}{2}+\kappa)\Gamma(\frac{\mu+d+1}{2}+2\kappa)}{\Gamma(\frac{\mu}{2})\Gamma(\frac{\mu+d}{2}+\kappa)}.$ (24)
-Note that the equivalence condition (23) sharing a common variance parameter and taking into account the relation between the smoothness parameters $\nu=\kappa+\frac{1}{2}$, can be written as
+where
+$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+d+1}{2}+2\kappa\right)}{\Gamma\!\left(\frac{\mu}{2}\right)\Gamma\!\left(\frac{\mu+d}{2}+\kappa\right)}.$ (20)
+Assuming a common variance and imposing the relation $\nu=\kappa+\frac{1}{2}$, the equivalence condition (19) is equivalently expressed as
 $a=\alpha\,B(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
-If we use the previous equation as a compact support reparameterization in the ${\cal H}$ model, then, using the same arguments as in Bevilacqua et al., (2022), it can be shown that
+Using this compact-support reparameterization and following the arguments in Bevilacqua et al. (2022), one obtains
 $\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa>-\frac{1}{2},$ (21)
-with uniform convergence on $[0,\infty)$. In addition, following the same arguments in Emery et al., (2025), it can be shown that
-$\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,\mu\alpha,d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa\geq-\frac{1}{2},$ (26)
-with uniform convergence on $[0,\infty)$.
-As a consequence, the two reparameterized versions of the ${\cal H}$ model can be viewed as generalizations of the ${\cal MT}$ model with an additional parameter $\mu$ that, for given smoothness and spatial dependence parameters, allows to obtain compactly ($\mu<\infty$) or globally ($\mu=\infty$) supported covariance functions.
-Thus, under this parameterization, $\mu$ can be fixed by the user, particularly when dealing with large datasets, to achieve sparse covariance matrices and obtain computational benefits.
-An illustration of (25) can be found in Figure 4 where the ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ model with increasing values of $\mu$ is compared with the ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$ model when $d=2$.
-Refer to caption Refer to caption
-Figure 4: The ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,2),2}$ for $\mu=1,2,10$ when $\kappa=0$ (left) and $\kappa=1$ (right).
-The red line is the ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$ correlation model ($\alpha=0.2$ in both cases).
-It should be emphasized that, under this compact support reparameterization, the role of the shape parameter $\mu$ changes significantly.
-Let us assume, without loss of generality, that the scale dependence parameter $\alpha$ is equal to $1$.
-Under this assumption, the parameter $\mu$ (together with the smoothness parameter and the dimension $d$) determines the compact support.
-For large values of $\mu$, the compact support becomes larger, and both the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ and ${\cal H}_{\kappa,\mu,\mu,d}$ models approach the ${\cal MT}_{\kappa+\frac{1}{2},1}$ model.
-In this case, the maximum integral range is the one associated with the ${\cal MT}$ model in Equation (17).
-Conversely, as $\mu$ approaches its lower bound, the compact support shrinks, and a potential smaller compact support results in a more flexible model.
-If we compare the compact support of $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ with that of $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ (see Equation 9) and given that $\mu$ is lower bounded by $1$ for the $\mathcal{H}$ model and by $\mu_{\min}(\kappa,d)$ for the $\mathcal{GW}$ model, it can be shown that
-$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa)$
-when $d\geq 2$ or $\kappa\geq 0$ and $b(1,1,\kappa,d)>g(1,\mu_{\min}(\kappa,d),\kappa)$ when $d=1$ and $-\frac{1}{2}<\kappa<0$.
-This implies that the $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model achieves a smaller compact support compared to the $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model for a significant portion of the parameter space.
-As a consequence, the $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model is more flexible from this perspective.
-On the other hand, if we compare the smaller compact supports of $\mathcal{H}_{\kappa,\mu,\mu,d}$ and $\mathcal{GW}_{\kappa,\mu,\mu}$, then we reach the same conclusions since $1\leq\mu_{\min}(\kappa,d)$ when $d\geq 2$ or $\kappa\geq 0$ and $1>\mu_{\min}(\kappa,d)$ when $d=1$ and $-\frac{1}{2}<\kappa<0$.
-We also compare the smaller compact supports that can be achieved by the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ and ${\cal H}_{\kappa,\mu,\mu,d}$ models, which are $b(1,1,\kappa,d)$ and $1$, respectively.
-Figure 5 illustrates the compact support $b(1,1,\kappa,d)$ for $-0.5<\kappa<1.5$ and $d=1,2,3,4,5$, compared to the reference value $1$.
-It can be observed that only for a small region of the smoothness parameter space and $d=1,2,3,4$, we obtain $b(1,1,\kappa,d)<1$.
-Figure 5: The smaller compact support $b(1,1,\kappa,d)$ that can be achieved by the model ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ for $d=1,2,3,4,5$ and $-0.5<\kappa<1.5$.
-The red line is the smaller compact support that can be achieved by the ${\cal H}_{\kappa,\mu,\mu,d}$ model.
-Summarizing our results, when comparing the two reparameterized versions of the $\mathcal{H}$ model with the analogous versions of the $\mathcal{GW}$ model that share the same smoothness parameter $\kappa$ and converge to the same $\mathcal{MT}$ model as the shape parameter $\mu$ becomes very large, it turns out that the reparameterized $\mathcal{H}$ models can achieve smaller compact supports.
-In addition the ${\cal H}_{\kappa,\mu,\mu,d}$ model can achieve smaller compact supports with respect to the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model.
+with uniform convergence on $0,\infty)$.
+[…]
+It is worth emphasizing that, under this compact-support reparameterization, the role of the shape parameter $\mu$ changes substantially.
+For fixed smoothness $\kappa$ and scale parameter $\alpha=1$, $\mu$ determines the compact support.
+As $\mu\to\infty$, ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ converges to the Matérn model ${\cal MT}_{\kappa+\frac{1}{2},1}$, with maximal integral range given by (13).
+Conversely, as $\mu$ approaches its lower bound, the compact support decreases, potentially yielding more flexible local dependence structures.
+A comparison with the corresponding $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model (see (6)) shows that, over a substantial portion of the parameter space, the reparameterized $\mathcal{H}$ models can achieve smaller compact supports.
+In particular, when $d\geq 2$ or $\kappa\geq 0$,
+$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa),$
+where $\mu_{\min}(\kappa,d)=\kappa+\frac{d+1}{2}$.
+The case $d=1$ and $\frac{1}{2}<\kappa<0$ is excluded here, as the exact lower bound $\mu_{\min}$ for the $\mathcal{GW}$ model remains an open problem in that regime.
 6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
-When considering the asymptotic properties of the ML estimator for the parameters of the $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ model, two different scenarios can be examined.
-Under increasing domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ can be estimated consistently.
-Furthermore, the asymptotic distribution of the ML estimator is Gaussian, with a covariance matrix equal to the inverse of the Fisher information matrix (Mardia and Marshall,, 1984).
-Under fixed-domain, using the results of Theorem 1 and taking into account the parameterization (15) with $l=\frac{d}{2}+\kappa$, the equivalence condition for the two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing a common smoothness parameter, when $d=1,2,3$ and provided that $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, reads:
-$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}},$ (27)
-As a consequence only the microergodic parameter $\sigma^{2}B(\kappa,\mu,d)/a^{2\kappa+1}$ can be estimated consistently.
-On the other hand, assuming a fixed known shape parameter $\mu$, the microergodic parameter is given by $\sigma^{2}/a^{2\kappa+1}$.
-In what follows, we provide the asymptotic properties of ML estimator of the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$.
-In particular, we will show that the microergodic parameter can be estimated consistently, and then we will assess the asymptotic distribution of the ML estimator.
-Let $D$ be a bounded subset of $\mathbb{R}^{d}$ and let $\boldsymbol{Z}_{n}=(Z(\boldsymbol{s}_{1}),\ldots,Z(\boldsymbol{s}_{n}))^{\prime}$ be a finite realization of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, a zero-mean stationary Gaussian RF with a given parametric covariance function $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, with $\sigma^{2}>0$, $\boldsymbol{\tau}$ a parameter vector and $\phi$ a member of the class $\Phi_{d}$, with $\phi(0;\boldsymbol{\tau})=1$.
-We then write $R_{n}(\boldsymbol{\tau})=[\phi(\|\boldsymbol{s}_{i}-\boldsymbol{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$ for the associated correlation matrix.
-The Gaussian log-likelihood function is defined as:
-$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\left(n\log(2\pi\sigma^{2})+\log(|R_{n}(\boldsymbol{\tau})|)+\frac{1}{\sigma^{2}}\boldsymbol{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\boldsymbol{Z}_{n}\right).$ (28)
-Under the $\cal H$ model, the Gaussian log-likelihood is obtained with $\phi(\cdot;\boldsymbol{\tau})\equiv\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$.
-Since in what follows $\mu$ and $\kappa$ are assumed known and fixed, for notation convenience, we write $\tau=a$.
-Let $\hat{\sigma}^{2}_{n}$ and $\hat{a}_{n}$ be the ML estimator obtained maximizing $\mathcal{L}_{n}(\sigma^{2},a)$ for a fixed $\mu$ and $\kappa$.
-We consider two types of estimators.
-The first maximizes (28) with respect to $\sigma^{2}$ for a fixed arbitrary compact support $a>0$, obtaining the following estimator
-$\hat{\sigma}_{n}^{2}(a)=\operatorname*{arg\,max}_{\sigma^{2}}\mathcal{L}_{n}(\sigma^{2},a)=\boldsymbol{Z}_{n}^{\prime}R_{n}(a)^{-1}\boldsymbol{Z}_{n}/n.$ (29)
-Here $R_{n}(a)$ is the correlation matrix coming from the $\mathcal{H}_{\kappa,\mu,a,d}$ model.
-The following result establishes strong consistency and asymptotic distribution of the ML estimator $\hat{\sigma}_{n}^{2}(a)/a^{(1+2\kappa)}$ of the microergodic parameter.
-We omit the proof, as it follows the same steps as Theorem 8 in Bevilacqua et al., (2019).
+We study the asymptotic properties of the maximum likelihood (ML) estimator for the parameters of the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ under both increasing-domain and fixed-domain asymptotics.
+Under increasing-domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ are consistently estimable, and the ML estimator is asymptotically Gaussian with covariance matrix equal to the inverse Fisher information (Mardia and Marshall, 1984).
+Under fixed-domain asymptotics, the situation differs.
+Using Theorem 1 and the parameterization (11) with $l=\frac{d}{2}+\kappa$, the equivalence condition for two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing the same smoothness parameter, for $d=1,2,3$ and $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, is
+$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}}.$ (22)
+Hence only the microergodic parameter
+$\frac{\sigma^{2}B(\kappa,\mu,d)}{a^{2\kappa+1}}$
+is consistently estimable. If $\mu$ is fixed and known, this reduces to $\sigma^{2}/a^{2\kappa+1}$.
+Let $D\subset\mathbb{R}^{d}$ be bounded and $\mathbf{Z}_{n}=(Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n}))^{\prime}$ a realization of a zero-mean stationary Gaussian RF with covariance $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, $\phi\in\Phi_{d}$, $\phi(0;\boldsymbol{\tau})=1$.
+Denote by
+$R_{n}(\boldsymbol{\tau})=[\phi(\|\mathbf{s}_{i}-\mathbf{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$
+the correlation matrix. The Gaussian log-likelihood is
+$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\Big(n\log(2\pi\sigma^{2})+\log|R_{n}(\boldsymbol{\tau})|+\frac{1}{\sigma^{2}}\mathbf{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\mathbf{Z}_{n}\Big).$ (23)
+Under the $\mathcal{H}$ model, $\phi(\cdot;\boldsymbol{\tau})=\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$.
+In what follows, $\kappa$ and $\mu$ are assumed fixed and known, so we write $\tau=a$.
+Let $\hat{\sigma}_{n}^{2}(a)$ denote the ML estimator of $\sigma^{2}$ for fixed $a>0$:
+$\hat{\sigma}_{n}^{2}(a)=\mathbf{Z}_{n}^{\prime}R_{n}(a)^{-1}\mathbf{Z}_{n}/n.$ (24)
+The following result establishes strong consistency and asymptotic normality of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$. The proof follows the same arguments as Theorem 8 in Bevilacqua et al. (2019).
 Theorem 10.
-Let $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ covariance model with $\mu\geq 1+d/2$.
-Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times(0,\infty)$.
-For a fixed $a>0$, let $\hat{\sigma}_{n}^{2}(a)$ as defined in Equation (29).
+Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 1+\frac{d}{2}$.
+Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)^{2}$.
+Then, for fixed $a>0$, as $n\to\infty$,
+  1. 1.
+    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
+  2. 2.
+    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
+We next consider joint maximization of (23) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$, $I=[a_{L},a_{U}]$, $0<a_{L}<a_{U}<\infty$. Let
+$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\mathbf{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\mathbf{Z}_{n}/n,\quad\hat{a}_{n}=\arg\max_{a\in I}\mathcal{PL}_{n}(a),$
+where $\mathcal{PL}_{n}(a)$ is the profile log-likelihood
+$\mathcal{PL}_{n}(a)=-\frac{1}{2}\big(\log(2\pi)+n\log\hat{\sigma}_{n}^{2}(a)+\log|R_{n}(a)|+n\big).$ (25)
+The following lemma establishes a monotonicity property of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ as a function of $a$. Its proof is given in Appendix A.
+Lemma 1.
+Let $S_{n}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\subset D$. For any $a_{1}<a_{2}$ in $I$,
+$\frac{\hat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}\leq\frac{\hat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}\quad\text{if and only if}\quad\mu\geq 4.$
+The next theorem establishes the asymptotic properties of $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$. The proof follows Theorem 9 in Bevilacqua et al. (2019) combined with Lemma 1.
+Theorem 11.
+Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 4$.
+Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$.
 Then, as $n\to\infty$,
   1. 1.
-    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa}$ and
+    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
   2. 2.
-    $\sqrt{n}(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,2(\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})^{2})$.
-The second type of estimation technique considers the joint maximization of (28) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$ where $I=[a_{L},a_{U}]$ and $0<a_{L}<a_{U}<\infty$.
-The solution of this optimization problem is given by $(\hat{\sigma}_{n}^{2}(\hat{a}_{n}),\hat{a}_{n})$ where
-$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\boldsymbol{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\boldsymbol{Z}_{n}/n$
-and $\hat{a}_{n}=\operatorname*{arg\,max}_{a\in I}\mathcal{PL}_{n}(a)$. Here $\mathcal{PL}_{n}(a)$ is the profile log-likelihood:
-$\mathcal{PL}_{n}(a)=-\frac{1}{2}\left(\log(2\pi)+n\log(\hat{\sigma}_{n}^{2}(a))+\log|R_{n}(a)|+n\right).$ (30)
-In order to establish strong consistency and asymptotic distribution of the sequence of random variables $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$, we use the following lemma that establishes the monotone behavior of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ when viewed as a function of $a\in I$ under specific conditions on the $\mu$ parameter.
-The proof can be found in Appendix A.
-Lemma 1.
-Let $S_{n}=\{\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n}\in D\subset\mathbb{R}^{d}\}$ denote any set of distinct locations.
-For any $a_{1}<a_{2}\in I$ and for each $n$, $\hat{\sigma}_{n}^{2}(a_{1})/a_{1}^{1+2\kappa}\leq\hat{\sigma}_{n}^{2}(a_{2})/a_{2}^{1+2\kappa}$ if and only if $\mu\geq 4$.
-Finally, the following theorem establishes the asymptotic properties of the ML estimator $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$.
-We omit the proof, as it follows the same steps as Theorem 9 in Bevilacqua et al., (2019) using the results in Lemma 1.
-Theorem 11.
-Let $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ covariance model with $\mu\geq 4$.
-Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$ where $I=[a_{L},a_{U}]$ with $0<a_{L}<a_{U}<\infty$.
-Let $(\hat{\sigma}_{n}^{2},\hat{a}_{n})^{\prime}$ maximize (28) over $(0,\infty)\times I$.
-Then as $n\to\infty$,
-  1. 1.
-    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa}$ and
-  2. 2.
-    $\sqrt{n}(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,2(\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})^{2})$.
-We conclude this section with some comments on the reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$.
-In particular, assuming $\mu$ unknown and using (27), for the latter model the associated microergodic parameter is given by $\sigma^{2}B(\kappa,\mu,d)/(\alpha\mu)^{2\kappa+1},$ and for the former model, it is given by $\sigma^{2}/\alpha^{2\kappa+1}.$ Notably, the microergodic parameter in the second case does not depend on $\mu$.
-This implies that, for the $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ model, we have two microergodic parameters, $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$, that can be estimated consistently under fixed-domain asymptotics.
-As a consequence, from an inferential point of view, the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ should be preferred over the $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$ model.
-However, particularly for large datasets, the $\mu$ parameter should be fixed, as it allows the construction of sparse covariance matrices.
-In this case, the microergodic parameter is given by $\sigma^{2}/\alpha^{2\kappa+1}$ for both parameterizations.
+    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
+The condition $\mu\geq 4$ ensures the monotonicity property in Lemma 1 and is technical in nature.
+Finally, we compare the two reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$.
+Under fixed-domain asymptotics and using (22), the associated microergodic parameters are
+$\frac{\sigma^{2}}{\alpha^{2\kappa+1}}\quad\text{and}\quad\frac{\sigma^{2}B(\kappa,\mu,d)}{(\alpha\mu)^{2\kappa+1}},$
+respectively.
+In the first case, the microergodic parameter does not depend on $\mu$, so both $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$ are consistently estimable.
+From an inferential perspective, $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ is therefore preferable.
+For large datasets, however, $\mu$ may be fixed to induce sparsity, in which case the microergodic parameter reduces to $\sigma^{2}/\alpha^{2\kappa+1}$ under both parameterizations.
 7 A simulation study
-In this section, we perform a simulation study on the estimation of the $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ family.
-We simulate 500 realizations, using Cholesky decomposition, of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $d=2$, setting $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$.
-We have considered a regular grid with increments $0.025$ over $[0,1]^{2}$, that is $1681$ grid points.
-Then the grid points have been perturbed, adding a uniform random value on $[-0.01,0.01]$ to each coordinate.
-We assume the parameters $\kappa$ and $\mu$ known and we focus on the ML estimation of $\sigma^{2}$ and $a$ and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing and fixed domain scenarios.
-Since the increasing as well as the fixed-domain frameworks can be mimicked by fixing the number of location sites over a given spatial domain and decreasing or increasing the spatial dependence (Zhang and Zimmerman,, 2005; Kaufman and Shaby,, 2013), we set the compact support $a=0.1,0.25,0.5$.
-We consider $a=0.1$ as an increasing domain scenario and $a=0.5$ as a fixed domain scenario.
-Table 1 reports biases and standard deviations (in parentheses) of the 500 ML parameter estimates, namely $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$, $\kappa=0,1,2$, $\mu=4,6,8$ and $a=0.1,0.25,0.5$.
+We investigate the finite-sample performance of ML estimation for the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$.
+We simulate 500 realizations of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ for $d=2$, $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$.
+The process is generated via Cholesky decomposition on a regular grid with spacing $0.025$ over $0,1]^{2}$ (1681 locations), followed by independent perturbations of each coordinate by a uniform random variable on [0.01,0.01][-0.01,0.01].
+Throughout the study, $\kappa$ and $\mu$ are assumed known.
+We focus on ML estimation of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing- and fixed-domain scenarios.
+Following Zhang and Zimmerman (2005); Kaufman and Shaby (2013), these regimes can be mimicked by fixing the sampling grid and varying the spatial dependence.
+Accordingly, we consider $a=0.1,0.25,0.5$, interpreting $a=0.1$ as increasing-domain and $a=0.5$ as fixed-domain behavior.
+Table 1 reports the bias and standard deviation (in parentheses) of the 500 ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$, for $i=1,\ldots,500$.
 Table 1: Bias and standard deviation (in parenthesis) of ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ when $\kappa=0,1,2$, $\mu=4,6,8$, $a=0.1,0.25,0.5$ and $\sigma^{2}=1$.
@@
 $(0.0808)$ $(0.0622)$ $(0.0522)$ $(0.1082)$ $(0.0870)$ $(0.0737)$ $(0.1044)$ $(0.088)$ $(0.0773)$
-$a=0.5$ $-0.0023$ $0.0024$ $0.0045$ $0.0029$ $0.0079$ $0.0101$ $0.0052$ $0.0091$ $0.0116$
+$a=0.5$ 0.0023-0.0023 $0.0024$ $0.0045$ $0.0029$ $0.0079$ $0.0101$ $0.0052$ $0.0091$ $0.0116$
 $(0.1446)$ $(0.1103)$ $(0.0900)$ $(0.1921)$ $(0.1592)$ $(0.135)$ $(0.1892)$ $(0.1599)$ $(0.1397)$
@@
 $(0.0214)$ $(0.0176)$ $(0.0157)$ $(0.0099)$ $(0.0086)$ $(0.0079)$ $(0.0059)$ $(0.0054)$ $(0.0052)$
-$a=0.5$ $-0.0015$ $0.0010$ $0.0023$ $-0.0016$ $-0.0001$ $0.0008$ $-0.0009$ $-0.0001$ $0.0005$
+$a=0.5$ 0.0015-0.0015 $0.0010$ $0.0023$ 0.0016-0.0016 0.0001-0.0001 $0.0008$ 0.0009-0.0009 0.0001-0.0001 $0.0005$
 $(0.0737)$ $(0.0578)$ $(0.0488)$ $(0.0329)$ $(0.0280)$ $(0.0245)$ $(0.0196)$ $(0.0171)$ $(0.0154)$
-$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ $-0.0150$ $-0.0219$ $-0.0151$ $-3.9684$ $-4.4120$ $-3.7659$ $-496.06$ $-486.01$ $-346.94$
+$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ 0.0150-0.0150 0.0219-0.0219 0.0151-0.0151 3.9684-3.9684 4.4120-4.4120 3.7659-3.7659 496.06-496.06 486.01-486.01 346.94-346.94
 $(0.4370)$ $(0.5362)$ $(0.6792)$ $(51.3695)$ $(68.6746)$ $(88.9434)$ $(6397.8)$ $(8930.7)$ $(11628.7)$
-$a=0.25$ $0.0045$ $0.0015$ $-0.0017$ $0.0228$ $-0.0470$ $-0.1218$ $-0.2553$ $-1.7130$ $-3.0083$
+$a=0.25$ $0.0045$ $0.0015$ 0.0017-0.0017 $0.0228$ 0.0470-0.0470 0.1218-0.1218 0.2553-0.2553 1.7130-1.7130 3.0083-3.0083
 $(0.1499)$ $(0.1600)$ $(0.1708)$ $(2.5102)$ $(2.7402)$ $(2.9959)$ $(41.521)$ $(46.056)$ $(51.067)$
@@
 $(0.0719)$ $(0.0740)$ $(0.0765)$ $(0.2923)$ $(0.3050)$ $(0.3187)$ $(1.1922)$ $(1.2484)$ $(1.3091)$
-Overall, the ML estimation of $\sigma^{2}$ and $a$ is approximately unbiased, and the standard deviations increase when switching from the increasing domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$), as expected.
-Another general pattern that can be observed is that, for each $\kappa$, increasing the shape parameter $\mu$ results in a lower standard deviation.
-For the microergodic parameter, it can be observed that under the increasing domain scenario ($a=0.1$), the ML estimation can be significantly unbiased but with a high standard deviation, particularly when $\kappa$ increases.
-When moving to the fixed-domain scenario ($a=0.5$), both the bias and the standard deviation are significantly reduced, as expected.
-In addition, the standard deviation of the microergodic parameter increases as the smoothness parameter $\kappa$ increases.
-Figure 6 depicts the boxplots of the ML estimates of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (on the left) and $a=0.5$ (on the right) when $\kappa=0$.
-The Gaussian distribution of the estimates for $\sigma^{2}$ and $a$ is more evident under the increasing domain scenario, as expected.
-On the other hand, when switching to the fixed-domain scenario, the distribution of the ML estimates of the microergodic parameter becomes more centered around the true values, and the distribution appears Gaussian, as expected.
+Overall, the ML estimators of $\sigma^{2}$ and $a$ are approximately unbiased across all configurations.
+As expected, their standard deviations increase when moving from the increasing-domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$).
+For fixed $\kappa$, larger values of $\mu$ generally reduce variability.
+Regarding the microergodic parameter, under the increasing-domain scenario ($a=0.1$) the estimator exhibits moderate bias and relatively high variance, particularly for larger $\kappa$.
+Under the fixed-domain scenario ($a=0.5$), both bias and variance decrease substantially, and the sampling distribution appears closer to Gaussian, in agreement with the asymptotic theory.
+Moreover, the variability of the microergodic estimator increases with the smoothness parameter $\kappa$.
+Figure 4 displays boxplots of the ML estimates of $\sigma^{2}$, $a$, and $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (left) and $a=0.5$ (right), when $\kappa=0$.
+Under the increasing-domain scenario, the distributions of $\widehat{\sigma}^{2}$ and $\widehat{a}$ are approximately Gaussian.
+In the fixed-domain setting, the microergodic estimator becomes more concentrated around the true value and exhibits an approximately Gaussian distribution, as predicted by the theory.
 Refer to caption Refer to caption
@@
 8 A real data application
-We consider data from WorldClim (www.worldclim.org), a global database of high spatial resolution weather and climate data for the years 1970–2000, which can be downloaded using the R package geodata (Hijmans et al.,, 2024).
-Specifically, we analyze mean temperature data for September observed at 725,330 georeferenced locations across Madagascar.
-To make ML estimation feasible, we select a random sample of 3,500 locations and perform the analysis using a UTM projection.
-Following Li and Zhang, (2011), we first detrend the data using splines to remove the cyclic patterns of both variables along the two coordinates.
-We then treat the residuals $y(\boldsymbol{s}_{i})$, $i=1,\ldots,3500$, as realizations from a zero-mean RF $\{Y(\boldsymbol{s}):\boldsymbol{s}\in D\}$.
-Figure 7, from left to right, depicts the colored map, the boxplot, and the empirical semivariogram of the residuals.
-The graphics suggest that a weakly stationary RF with a marginal distribution capable of modeling heavy tails is potentially an appropriate model for our data.
-Refer to caption Refer to caption Refer to caption
-Figure 7: From left to right: spatial map of the residuals for the Madagascar mean temperature data, boxplot, and empirical semivariogram of the data.
-Let $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ be a standard Gaussian RF with an isotropic parametric correlation function. In our analysis, we consider two specifications for $Y(\boldsymbol{s})$:
-  • a zero-mean Gaussian RF, i.e., $Y(\boldsymbol{s})=\sigma Z(\boldsymbol{s})$, and
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+We consider data from WorldClim (www.worldclim.org), which provides high-resolution gridded climate surfaces for 1970–2000 (Fick and Hijmans, 2017) and can be accessed in R via geodata (Hijmans et al., 2024).
+We analyze September mean temperature over Madagascar.
+The full dataset consists of 725,330 grid-cell predictions obtained by interpolating weather-station data using thin-plate splines with geographic and satellite-derived covariates (Fick and Hijmans, 2017).
+To make ML estimation computationally feasible, we randomly sample 3,500 grid cells and use UTM-projected coordinates.
+Following Li and Zhang (2011), the data are first detrended using spline regression to remove large-scale spatial trends.
+The residuals $y(\mathbf{s}_{i}),i=1,\ldots,3500$, are then treated as observations from a zero-mean RF $\{Y(\mathbf{s}):\mathbf{s}\in D\}$.
+Figure 5 displays the spatial map of the residuals and the empirical semivariogram.
+The plots suggest that a weakly stationary RF with a potentially heavy-tailed marginal distribution provides a reasonable modeling framework.
+Refer to caption Refer to caption
+Figure 5: Left: spatial map of residuals. Right: empirical semivariogram.
+Let $Z(\mathbf{s})$ be a standard Gaussian RF with isotropic correlation function. We consider two specifications for $Y(\mathbf{s})$:
+  • Gaussian RF: $Y(\mathbf{s})=\sigma Z(\mathbf{s})$;
+  • Tukey-$h$ RF (Xu and Genton, 2017; Caamaño-Carrillo et al., 2024): $Y(\mathbf{s})=\sigma Z(\mathbf{s})\exp\!\left(\tfrac{hZ^{2}(\mathbf{s})}{2}\right)$, where $h\in[0,1/2)$ controls tail heaviness.
+Since the transformation is monotone, the Tukey-$h$ RF preserves the geometric properties of the underlying Gaussian field.
+For the correlation of $Z(\mathbf{s})$, we compare three models: ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$, and the proposed ${\cal H}_{\kappa,\mu,\beta,2}$.
+Here $\kappa$ denotes smoothness, $\beta$ the scale parameter, and $\mu$ the shape parameter.
+We adopt an increasing-domain framework and estimate all covariance parameters by ML.
+Standard errors are computed from the inverse Fisher information matrix.
+For the Tukey-$h$ RF, ML estimation is implemented via the Lambert-$W$ function (see equation 10 in Caamaño-Carrillo et al.
+(2024)).
+Table 2 reports ML estimates, associated standard errors, maximized log-likelihood values, and AIC values.
+Overall, the AIC criterion favors the Tukey-$h$ specification, irrespective of the correlation model.
+Among the correlation structures, the compactly supported ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$ achieve the best AIC values.
+Figure 6 compares the normalized histogram of the residuals with the fitted density (left panels) and the empirical semivariogram with the fitted semivariogram (right panels), for the Gaussian (first row) and Tukey-$h$ (second row) models using the estimated ${\cal H}_{\kappa,\mu,\beta,2}$ correlation.
+The Tukey-$h$ RF provides a noticeably better fit to both marginal tail behavior and spatial dependence.
+Table 2: ML estimates with associated standard error (in parentheses), RMSE, MAE and percentage of zero values in the covariance matrix for the Gaussian and Tukey-$h$ RFs with underlying correlation functions ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$.
+Last row reports the time (in seconds) needed to compute the Cholesky factor of the covariance matrix.
+Gaussian RF Tukey-$h$ RF
 ${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$ ${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$
@@
 $(0.0553)$ $(0.0541)$ $(0.0540)$ $(0.0418)$ $(0.0403)$ $(0.0403)$
-$\kappa$ $0.6079$ $0.0305$ $0.0313$ $0.5767$ $-0.0036$ $-0.0043$
+$\kappa$ $0.6079$ $0.0305$ $0.0313$ $0.5767$ 0.0036-0.0036 0.0043-0.0043
 $(0.0293)$ $(0.0307)$ $(0.0312)$ $(0.0265)$ $(0.0237)$ $(0.0237)$
@@
 $(0.0172)$ $(0.0170)$ $0.0170$
-loglik $-3900.01$ $-3896.88$ $-3896.94$ $-3689.88$ $-3683.81$ $-3683.94$
+loglik 3900.01-3900.01 3896.88-3896.88 3896.94-3896.94 3689.88-3689.88 3683.81-3683.81 3683.94-3683.94
 AIC $7806$ $7802$ $7802$ $7388$ $7378$ $7378$
@@
 MAE $0.4068$ $0.4065$ $0.4065$ $0.4073$ $0.4071$ $0.4070$
-[…]
-$Time$ $7.22$ $0.04$ $0.04$ $6.26$ $0.03$ $0.03$
-[…]
+$\%$ 0 $96.45$ $96.23$ 0 $97.18$ $97.06$
+TimeTime $7.22$ $0.04$ $0.04$ $6.26$ $0.03$ $0.03$
+We assess predictive performance via repeated cross-validation.
+For each of 100 repetitions, 85% of locations are used for parameter estimation and the remaining 15% for prediction.
+Table 2 reports the average RMSE and MAE.
+The best predictive performance is achieved by the Tukey-$h$ RF combined with compactly supported correlations ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$.
 […]
@@
 Refer to caption Refer to caption
-[…]
+Figure 6: First row: normalized histogram versus estimated density and empirical semivariogram versus estimated semivariogram when using an estimated Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ correlation model (from left to right).
+Second row: the same graphics using a Tukey-$h$ RF.
 9 Concluding Remarks
-In this paper, we study the Gauss hypergeometric covariance model, a highly flexible kernel introduced by Emery and Alegría, (2022), which generalizes both the Matérn and Generalized Wendland models.
-We establish necessary and sufficient conditions for the validity of the model and highlight that the model suffers from identifiability issues under both increasing and fixed domain asymptotics.
-To address the identifiability problem, we introduce a new class of compactly supported kernels that satisfy a specific optimality criterion.
-In particular, the proposed model achieves the maximum integral range within the class of Gauss hypergeometric correlations.
-From this perspective, the proposed model can be regarded as an enhancement of the Generalized Wendland model.
-The proposed model has the potential to outperform the Matérn family in terms of model fitting, prediction performance, and computational efficiency, as demonstrated in the real-data application.
-Additionally, the proposed model can be easily implemented, as efficient numerical computation of the Gaussian hypergeometric function is available in various libraries, such as the GNU Scientific Library (Gough,, 2009), and in major statistical software packages, including R, MATLAB, and Python.
-In particular, the R package GeoModels (Bevilacqua et al.,, 2025), used in this paper for numerical experiments and applications, computes the proposed model using the Python implementation of the Gaussian hypergeometric function in the SciPy library (Virtanen et al.,, 2020).
-From an inferential perspective, the results on the asymptotic properties of the maximum likelihood estimator of the microergodic parameter could be generalized to include a nugget effect, as in Tang et al., (2021).
-Additionally, the estimation of the smoothness parameter could, in principle, be studied following the ideas in Loh and Sun, (2023).
-In a Bayesian setup, the fixed-domain asymptotic posterior distribution of the microergodic parameter could be investigated following Li, (2022).
-Finally, the proposed model can be generalized in various directions, such as space-time, multivariate, or non-stationary contexts.
-These topics are left for future research.
+In this paper we study the Gauss hypergeometric covariance model, a flexible class introduced by Emery and Alegría (2022) that encompasses both the Matérn and Generalized Wendland families.
+We derive necessary and sufficient validity conditions and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics.
+To address this issue, we introduce a parsimonious compactly supported subclass selected through an optimality criterion based on maximum integral range.
+In this sense, the proposed hypergeometric model can be viewed as a structural refinement of the Generalized Wendland construction.
+Although simple closed-form expressions are not generally available for nonnegative integer smoothness parameters, efficient numerical evaluation of the Gauss hypergeometric function is widely supported in modern scientific computing libraries.
+In particular, reliable implementations are available in the GNU Scientific Library (Gough, 2009) and in major environments such as R, MATLAB, and Python.
+The GeoModels package (Bevilacqua et al., 2026), used in this paper for numerical experiments and applications, computes the proposed model via the SciPy implementation of the hypergeometric function (Virtanen et al., 2020).
 Acknowledgement
-We would like to thank Professor Igor Kondrashuk for his help in partially proving Lemma 1.
 This work was funded by the National Agency for Research and Development of Chile, under grants ANID FONDECYT 1240308 (M.
@@
 Let $P(\rho_{i})$, $i=0,1$, be two zero-mean Gaussian measures with isotropic covariance function $\rho_{i}$ and associated spectral density $\widehat{\rho}_{i}$, $i=0,1$, as defined through (1).
-Using results in Skorokhod and Yadrenko, (1973) and Ibragimov and Rozanov, (1978), Stein, (2004) has shown that, if for some $a>0$, $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and for some finite and positive $c$,
-$\int_{c}^{\infty}z^{d-1}\;\left\{\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right\}^{2}\;{\rm d}z<\infty,$ (31)
-then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$.
-Then (31) is the key tool to prove Theorem 1.
-We first need the results in Lemma 2 hereunder.
-They follow from Zastavnyi, (2006), Chernih and Hubbert, (2014), and from standard properties of Fourier transforms and give asymptotic expansion of the spectral density of the $\mathcal{GH}$ model.
-Their proofs are thus omitted.
-For two given functions $g_{1}(x)$ and $g_{2}(x)$, with $g_{1}(x)\asymp g_{2}(x)$ we mean that there exist two constants $c$ and $C$ such that $0<c<C<\infty$ and $c|g_{2}(x)|\leq|g_{1}(x)|\leq C|g_{2}(x)|$ for each $x$.
+Using Skorokhod and Yadrenko (1973); Ibragimov and Rozanov (1978); Stein (2004), if for some $a>0$ the function $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and if for some finite and positive $c$,
+$\int_{c}^{\infty}z^{d-1}\left(\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right)^{2}\,{\rm d}z<\infty,$ (26)
+then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$.
+Condition (26) is the key tool.
+We first need the following result.
 Lemma 2.
-Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be the function defined at Equation (11). Then, for $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $12(\beta+\gamma)\geq 6\delta+$ 1
+Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be defined at (7). Assume $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $2(\beta+\gamma)\geq 6\delta+1$. Then:
   1. 1.
-    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\sigma^{2}Ka^{d}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
-        $+\,c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}$,  for $z\to\infty$;
+    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ $=\sigma^{2}Ka^{d}\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
+    $\qquad+c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big],$
+    as $z\to\infty$;
   2. 2.
-    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta}$,  for $z\to\infty,$
-where $c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}$, $c_{4}=\frac{\pi^{-\frac{1}{2}}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)}$ and $c_{5}=\frac{\pi}{2}(\frac{1}{2}+\delta-\beta-\gamma)$.
-We are now ready to prove the Theorem. Let us start with the sufficient part of the assertion. From Lemma  2 (Point 2), we have:
-$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}\leq c_{2},\quad i=0,1.$
-We then proceed by direct construction, and, using Lemma 2 (Points 1 and 2) and (31) we find that, as $z\to\infty$,
-$\Big{|}\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\Big{|}\leq\Big{|}\sigma_{1}^{2}K_{1}a_{1}^{d}\Big{[}c_{3}^{1}(za_{1})^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
-$\displaystyle+\,c_{4}^{1}(za_{1})^{\lambda_{1}}\big{\{}\cos(za_{1}+c_{5}^{1})+\mathcal{O}(z^{-1})\big{\}}\Big{]}-\sigma_{0}^{2}K_{0}a_{0}^{d}\Big{[}c_{3}^{0}(za_{0})^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
-$+\,c_{4}^{0}(za_{0})^{\lambda_{0}}\big{\{}\cos(za+c_{5}^{0})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\Big{|}\frac{z^{2\delta}}{c_{1}}$
-$\leq\frac{1}{c_{1}}\Big{|}\big{[}c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big{]}\{1+\mathcal{O}(z^{-2})\}+\big{[}c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z-c_{5}^{1})$
-$-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z-c_{5}^{0})\big{]}+\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}\Big{|},$
-where $c_{3}^{i}=\frac{2^{2\delta}\Gamma(\beta_{i})\Gamma(\gamma_{i})}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$, $c_{4}^{i}=\frac{\pi^{-\frac{1}{2}}\Gamma(\beta_{i})\Gamma(\gamma_{i})}{2^{\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}}\Gamma(\delta)}$,$c_{5}^{i}=\frac{\pi}{2}(\frac{1}{2}+\delta-\beta_{i}-\gamma_{i})$ and $\lambda_{i}=\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}$, $i=0,1$.
-Let us now write
-$A(z)$ $=\big{[}c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta)}K_{0}\big{]}\{1+\mathcal{O}(z^{-2})\},$
-$B(z)$ $=\big{[}c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z-c_{5}^{1})-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z-c_{5}^{0})\big{]}$
-$D(z)$ $=\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}.$
-Then, for some positive and finite $c$, Condition (31) can be written as
-$\int_{c}^{\infty}z^{d-1}\bigg{(}\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\biggr{)}^{2}{\rm d}z<\infty,$ (32)
-being equivalent to
+    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta},\qquad z\to\infty.$
+Here
+$c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)},\qquad c_{4}=\frac{\pi^{-1/2}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)},$
+$c_{5}=\frac{\pi}{2}\left(\frac{1}{2}+\delta-\beta-\gamma\right).$
+We are now ready to prove the theorem.
+From Lemma 2 (Point 2), there exist constants $0<c_{1}<c_{2}<\infty$ such that, for $z$ large,
+$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}(z)\leq c_{2},\qquad i=0,1.$ (27)
+In particular, $\widehat{\rho}_{0}(z)z^{2\delta}$ is bounded away from 0 and $\infty$ as $z\to\infty$.
+Using Lemma 2 (Points 1 and 2), and collecting terms, one obtains for large $z$
+$\left|\frac{\sigma_{1}^{2}\widehat{\varphi}_{1}(z)-\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}{\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}\right|\leq\frac{1}{c_{1}}\left|A(z)+B(z)+D(z)\right|,$
+where
+$A(z)$ $=\big[c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big]\{1+\mathcal{O}(z^{-2})\},$
+$B(z)$ $=\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z+c_{5}^{1})$
+$\quad-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z+c_{5}^{0})\big],$
+$D(z)$ $=\mathcal{O}(z^{-1})\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\big].$
+Hence condition (26) becomes
 $\int_{c}^{\infty}z^{d-1}\big(A(z)+B(z)+D(z)\big)^{2}\,{\rm d}z<\infty.$
-Now, we have that $A(z)$ is identically equal to zero thanks to our condition.
-We claim that the partial integral that depends of $B(z)$ and $D(z)$ is finite if $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$.
-In fact, we have, for $z\to\infty$,
-$|B(z)|$ $\leq c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\leq c_{6}\min\{z^{\lambda_{1}+2\delta},z^{\lambda_{0}+2\delta}\}$
-and
-$|D(z)|$ $\leq\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}$
-$\leq c_{7}\min\{z^{\lambda_{1}+2\delta},z^{\lambda_{0}+2\delta}\},$
-with $c_{i}$ being positive and finite constants, $i=6,7$. Thus, if $\frac{\sigma_{0}^{2}L_{0}}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L_{1}}{a_{1}^{2\delta-d}}$, we have
-$\int_{c}^{\infty}$ $z^{d-1}\left(\frac{\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\right)^{2}{\rm d}z$ (33)
-$\leq\int_{c}^{\infty}z^{d-1}(B(z)+D(z))^{2}{\rm d}z$
-$\leq\int_{c}^{\infty}z^{d-1}\Big{(}B(z)^{2}+D(z)^{2}+2B(z)D(z)\Big{)}{\rm d}z.$
-We conclude that (33) is finite if $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$. The sufficient part of our claim is thus proved.
-The necessary part follows the proof of Zhang, (2004).
-We suppose that $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$, $\frac{\sigma_{0}^{2}L_{0}}{a_{0}^{2\delta-d}}\neq\frac{\sigma_{1}^{2}L_{1}}{a_{1}^{2\delta-d}}$ and let $\sigma_{2}^{2}=\sigma_{0}^{2}L_{0}\left(\frac{a_{1}}{a_{0}}\right)^{2\delta-d}$.
-Then $\mathcal{GH}_{\delta,\beta_{0},\gamma_{0},a_{0}}$ and $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{2}}$ induce two equivalent Gaussian measures.
-This, we need to show that $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{2}}$ and $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{1}}$ define two orthogonal Gaussian measures.
-The rest of the proof follows the same arguments in Zhang, (2004).
-∎
+Assume now (10), i.e.
+$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}}.$
+This is equivalent to
+$c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}=c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0},$
+so that $A(z)\equiv 0$.
+Using $(x+y)^{2}\leq 2x^{2}+2y^{2}$, it suffices to check
+$\int_{c}^{\infty}z^{d-1+2(\lambda_{i}+2\delta)}\,dz<\infty,\qquad i=0,1,$
+which holds iff
+$\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta.$
+Therefore, under
+$\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})>\frac{d+1}{2}+3\delta,$
+condition (26) holds and the sufficient part is proved.
+The necessary part follows from Zhang (2004).
 A.2 Proof of Theorem 2
 Proof.
-Let us consider the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model and, for the moment, let us assume that $l$ is a real number.
-The sufficient conditions (a), (b), (c) given in Section 3, under the parameterization (15) are:
+Consider the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ in dimension $d$, where $\delta=\frac{d+1}{2}+\kappa$, $\kappa>-\frac{1}{2}$, $l\geq 0$.
+Sufficiency. Under $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, the sufficient conditions of Section 3 become:
   1. (a)
-    $\kappa>-\frac{1}{2}$;
+    $\kappa>-\frac{1}{2}$,
   2. (b)
-    $\mu(\mu+2l)\geq d+2\kappa+1$;
+    $\mu(\mu+2l)\geq d+2\kappa+1$,
   3. (c)
     $\mu+l\geq\frac{d}{2}+\kappa+1$.
-Solving the previous system of inequalities, we derive the following sufficient conditions:
-  1. 1.
-    For $d=1$
-    • a)
-      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l\leq-\frac{(1+2\kappa)}{2}$
-    • b)
-      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(1+2\kappa)}{2}<l\leq\frac{(1+2\kappa)}{2}$.
-    • c)
-      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l>\frac{(1+2\kappa)}{2}$.
-  2. 2.
-    For $d>1$
-    • a)
-      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l<-\frac{(d+2\kappa)}{2}$
-    • b)
-      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa)}{2}\leq l\leq\frac{(d+2\kappa)}{2}$.
-    • c)
-      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
-Necessary conditions can be obtained following Cho and Yun, (2018, Theorem 6.1), who show that if $\mathstrut{}_{1}F_{2}\left(a;b;c;-\frac{x^{2}}{4}\right)\geq 0$, then $b>a$, $c>a$, $b+c\geq 3a+\frac{1}{2}$.
-Using the spectral density (11), under the parameterization (15), we have
-  1. (a)
-    $\mu>0$;
-  2. (b)
-    $\mu>-2l$;
-  3. (c)
-    $\mu+l\geq\frac{d}{2}+\kappa+1$.
-Solving the system of inequalities (under the restriction $\kappa>-\frac{1}{2}$) for $d\geq 1$, we derive the following necessary conditions:
-  1. 1.
-    For $d=1$
-    • a)
-      $\mu>-2l$ and $l\leq-\frac{(3+2\kappa)}{2}$
-    • b)
-      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(3+2\kappa)}{2}<l\leq\frac{(3+2\kappa)}{2}$.
-    • d)
-      $\mu>0$ and $l>\frac{(3+2\kappa)}{2}$.
-  2. 2.
-    For $d>1$
-    • a)
-      $\mu>-2l$ and $l<-\frac{(d+2\kappa+2)}{2}$
-    • c)
-      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa+2)}{2}\leq l<\frac{(d+2\kappa+2)}{2}$.
-    • d)
-      $\mu>0$ and $l\geq\frac{(d+2\kappa+2)}{2}$.
-Putting together the necessary and sufficient conditions for the cases $d=1$ and $d>1$, we obtain the following conditions:
-  1. 1.
-    For $d=1$
-    • a)
-      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l\leq-\frac{(1+2\kappa)}{2}$
-    • b)
-      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(1+2\kappa)}{2}<l\leq\frac{(1+2\kappa)}{2}$.
-    • c)
-      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l>\frac{(1+2\kappa)}{2}$.
-  2. 2.
-    For $d>1$
-    • a)
-      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l<-\frac{(d+2\kappa)}{2}$
-    • b)
-      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa)}{2}\leq l\leq\frac{(d+2\kappa)}{2}$.
-    • c)
-      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
-We observe that the necessary and sufficient conditions are symmetric around $l=0$, since $\mathcal{GH}_{\delta,\beta,\gamma,a}(x)=\mathcal{GH}_{\delta,\gamma,\beta,a}(x)$.
-Specifically, for $d=1$, conditions a) and c) are symmetric when $l<-\frac{(1+2\kappa)}{2}$ and $l>\frac{(1+2\kappa)}{2}$, respectively, while condition b) is symmetric when $-\frac{(d+2\kappa)}{2}<l\leq 0$ and $0\leq l\leq\frac{(d+2\kappa)}{2}$.
-These same symmetry patterns are also observed for the case $d>1$.
-Using the symmetry of the conditions, we put together the necessary and sufficient conditions only for $l\geq 0$ obtaining the main result, that is for $\kappa>-\frac{1}{2}$ the model is valid if and only if
-  1. 1.
-    $\mu\geq\delta-(l-\frac{1}{2})$ and $0\leq l\leq\frac{(d+2\kappa)}{2}$
-  2. 2.
-    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
+Condition (b) is equivalent to the quadratic inequality
+$\mu^{2}+2l\mu-(d+2\kappa+1)\geq 0,$
+whose positive branch gives
+$\mu\geq-l+\sqrt{l^{2}+d+2\kappa+1}.$
+Only the positive branch is admissible since condition (c) implies $\mu>-l$ when $l\geq 0$. For each fixed $l$, the sufficient conditions therefore reduce to
+$\mu\geq\max\!\Big\{\frac{d}{2}+\kappa+1-l,\;-l+\sqrt{l^{2}+d+2\kappa+1}\Big\}.$
+A direct algebraic comparison shows that
+$\frac{d}{2}+\kappa+1-l\geq-l+\sqrt{l^{2}+d+2\kappa+1}\quad\Longleftrightarrow\quad 0\leq l\leq\frac{d}{2}+\kappa.$
+Hence condition (c) is binding when $0\leq l\leq\frac{d}{2}+\kappa$, giving $\mu\geq\delta-l+\frac{1}{2}$, and condition (b) is binding when $l>\frac{d}{2}+\kappa$, giving $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$.
+This yields exactly the two regimes in points 1 and 2.
+Necessity.
+Assume $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}$.
+Since the Gamma prefactors in the spectral density are positive whenever $\beta>\delta$ and $\gamma>\delta$, nonnegativity of the spectral density requires
+$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\right)\geq 0\quad\text{for all }z\geq 0.$
+By Cho and Yun (2018, Theorem 3.1), this implies $\beta>\delta$, $\gamma>\delta$, and $\beta+\gamma\geq 3\delta+\frac{1}{2}$.
+The third condition, under the parameterization $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, gives
+$\mu+l\geq\frac{d}{2}+\kappa+1,\quad\text{i.e.,}\quad\mu\geq\delta-l+\frac{1}{2},$
+which is the necessary lower bound for $\mu$ in both regimes.
+In Regime 1 ($0\leq l\leq\frac{d}{2}+\kappa$), this coincides with the sufficient condition in point 1, so the result is an if and only if.
+In Regime 2 ($l>\frac{d}{2}+\kappa$), this provides a necessary lower bound that is strictly weaker than the sufficient condition in point 2; whether the gap can be closed remains an open question.
+Reduction to $l\geq 0$.
+The model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under interchanging $\beta$ and $\gamma$, which corresponds to $l\mapsto-l$.
+The admissible region is therefore symmetric with respect to $l=0$, and restricting to $l\geq 0$ entails no loss of generality.
+This completes the proof. ∎
 A.3 Proof of Theorem 3
 Proof.
-The result stems from formula 2.21.1.11 of Prudnikov et al., (1990) and the expressions (12) and (16) of the hypergeometric correlation and of the integral range.
+Without loss of generality set $a=1$; the general case follows by the substitution $x\mapsto ax$, which contributes the factor $a^{d}$.
+Using the change of variable $t=1-x^{2}$ (so $x^{d-1}\,dx=\tfrac{1}{2}(1-t)^{(d-2)/2}\,dt$) in the definition of the integral range (12), and substituting (8), one obtains
+$A_{\kappa,\mu,l,1,d}$ $=\frac{\pi^{d/2}C_{0}}{2^{d}\Gamma(d/2)}\int_{0}^{1}(1-t)^{\frac{d-2}{2}}\,t^{\mu+\frac{d-1}{2}+2\kappa}$ (28)
+$\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;t\right)\,dt.$
+where $C_{0}=\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)\big/\!\left[\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\tfrac{1}{2}\right)\right]$.
+We apply formula 2.21.1.11 of Prudnikov et al. (1990),
+$\int_{0}^{1}t^{\sigma-1}(1-t)^{\rho-1}\,{}_{2}F_{1}(a,b;c;t)\,dt=\frac{\Gamma(\rho)\Gamma(\sigma)}{\Gamma(\rho+\sigma)}\,{}_{3}F_{2}(a,b,\sigma;\,c,\,\rho+\sigma;\,1),$ (29)
+with the identifications
+$\sigma=\mu+\tfrac{d+1}{2}+2\kappa,\quad\rho=\tfrac{d}{2},\quad a=\tfrac{\mu}{2},\quad b=\tfrac{\mu+d}{2}+\kappa,\quad c=\mu+\tfrac{d+1}{2}+2\kappa.$
+Since $c=\sigma$, the third upper parameter and the first lower parameter of $_{3}F_{2}$ coincide, so the series telescopes term by term:
+$_{3}F_{2}(a,b,\sigma;\,\sigma,\,\rho+\sigma;\,1)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}(\sigma)_{k}}{k!\,(\sigma)_{k}(\rho+\sigma)_{k}}=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{k!\,(\rho+\sigma)_{k}}={}_{2}F_{1}(a,b;\,\rho+\sigma;\,1).$
+Gauss’s summation theorem then gives, provided $\rho+\sigma-a-b>0$,
+$_{2}F_{1}(a,b;\rho+\sigma;1)=\frac{\Gamma(\rho+\sigma)\,\Gamma(\rho+\sigma-a-b)}{\Gamma(\rho+\sigma-a)\,\Gamma(\rho+\sigma-b)}.$
+One verifies directly that
+$\rho+\sigma-a-b=\kappa+\tfrac{1}{2}+l>0\quad(\kappa>-\tfrac{1}{2},\;l\geq 0),$
+$\rho+\sigma-a=\tfrac{\mu+1}{2}+\kappa+d,\qquad\rho+\sigma-b=\tfrac{\mu+1}{2}+\kappa+\tfrac{d}{2}.$
+Substituting back into (28), the factors $\Gamma(\sigma)$ and $\Gamma(\rho+\sigma)$ cancel with those arising from (29) and from $C_{0}$.
+Using the duplication formula $\Gamma(z)\Gamma(z+\tfrac{1}{2})=\sqrt{\pi}\,2^{1-2z}\Gamma(2z)$ to simplify the remaining Gamma factors, one obtains
+$A_{\kappa,\mu,l,1,d}=\frac{\pi^{d/2}\,\Gamma\!\left(\frac{d+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\,\Gamma\!\left(\kappa+\frac{1}{2}\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa+l\right)},$
+which, after restoring the factor $a^{d}$, gives (14). ∎
 A.4 Proof of Theorem 4
-Hereafter, without loss of generality consider a unit compact support $a=1$. Let us define $LA_{\kappa,d}(\mu,l):=log(A_{\kappa,\mu,l,d,1})$ where
-$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\Gamma\left(\frac{\mu+1}{2}+\kappa+d/2\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l+d/2\right)}.$
-is the integral range.
-[…]
+Proof.
+Without loss of generality, set the compact support to $a=1$. Define
+$LA_{\kappa,d}(\mu,l):=\log(A_{\kappa,\mu,l,d,1}),$
+where
+$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+\frac{d}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l+\frac{d}{2}\right)}.$
+We maximize $LA_{\kappa,d}(\mu,l)$ over the admissible set in $\mu$ for a fixed $l\geq 0$, as given by Theorem 2. Let $\delta=\frac{d}{2}+\kappa$. Then
   1. 1.
-    $\mu\geq\delta-l+0.5\quad$ if $\quad 0\leq l\leq d/2+\kappa$
+    $\mu\geq\delta-l+\tfrac{1}{2}$ if $0\leq l\leq\delta$;
   2. 2.
-    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>d/2+\kappa$.
-This is a maximization problem over a bounded interval.
-Hereafter, with $\psi(x)=\frac{\text{d}log\Gamma(x)}{\text{d}x}$ we denote the digamma function which is strictly increasing for $x>0$.
-The fist derivative of $LA_{\kappa,d}(\mu,l)$ with respect to $\mu$ is given by
-$\frac{\text{d}LA_{\kappa,d}(\mu,l)}{\text{d}\mu}=0.5[\psi(x)-\psi(x+d/2)+\psi(x+l)-\psi(x+l+d/2)],$
-where $x=\frac{\mu+1}{2}+\kappa$.
-It is a negative function since $\psi(\cdot)$ is strictly increasing and this implies that the function $LA_{\kappa,d}(\mu,l)$ is monotone non-increasing as a function of $\mu$.
-As a consequence the maximum is achieved by $LA_{\kappa,d}(\mu,l)$ on the lower boundary of region 1) and region 2), that is
-  1. 1.
-[…]
-  2. 2.
-[…]
-where $\mu_{1}(l)=\delta-l+0.5$ and $\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l$. This proves Theorem 4.
+    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ if $l>\delta$.
+This is a maximization problem over the half-line $\mu_{\min}(l),\infty)$.
+Let $\psi(x)=\frac{d}{dx}\log\Gamma(x)$ be the digamma function, which is strictly increasing for $x>0$.
+Set $x=\frac{\mu+1}{2}+\kappa$.
+Under the validity conditions we have $x>0$ and $x+l>0$.
+Differentiating with respect to $\mu$ yields
+$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)=\frac{1}{2}\Big[\psi(x)-\psi(x+\tfrac{d}{2})+\psi(x+l)-\psi(x+l+\tfrac{d}{2})\Big].$
+Since $\psi$ is increasing and $\tfrac{d}{2}>0$, we have $\psi(x)-\psi(x+\tfrac{d}{2})\leq 0$ and $\psi(x+l)-\psi(x+l+\tfrac{d}{2})\leq 0$, hence $\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)\leq 0$ for all admissible $\mu$.
+Therefore $LA_{\kappa,d}(\mu,l)$ is non-increasing in $\mu$, and its maximum is achieved at the smallest admissible value of $\mu$, i.e.
+$\mu_{1}(l)=\delta-l+\tfrac{1}{2}\quad(0\leq l\leq\delta),\qquad\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l\quad(l>\delta).$
+This proves Theorem 4. ∎
 A.5 Proof of Theorem 5
-We focus on the functions $LA_{\kappa,d}(\mu_{i}(l),l)$ for $i=1,2$ defined on the proof of Theorem 4 and the goal is to find the maximum with respect to $l$.
-If $i=1$, the derivative of $LA_{\kappa,d}(\mu_{1}(l),l)$ for $0\leq l\leq d/2+\kappa$ is given by
-$\frac{\text{d}LA_{\kappa,d}(\mu_{1}(l),l)}{\text{d}l}=0.5[\psi(x+l/2)-\psi(x-l/2)+\psi(x-l/2+d/2)-\psi(x+l/2+d/2)],$
-where $x=\frac{d}{4}+\frac{3\kappa}{2}+1$.
-It can be shown that this function is non-negative using the fact that the derivative of the digamma function $\frac{\text{d}\psi(x)}{\text{d}x}$ is a decreasing function.
-This implies that the maximum of $LA_{\kappa,d}(\mu_{1}(l),l)$ is achieved on the upper bound of the parametric space, that is $l=d/2+\kappa$.
-Using similar arguments, if $i=2$ it can be shown that the derivative of $LA_{\kappa,d}(\mu_{2}(l),l)$ for $l>d/2+\kappa$ is non-positive.
-As a consequence the maximum is achieved on the lower bound of the parametric space, that is when $l=d/2+\kappa+\epsilon$ where $\epsilon$ an arbitrary small positive constant.
-Finally we can prove that $LA_{\kappa,d}(\mu_{1}(d/2+\kappa),d/2+\kappa))=LA_{\kappa,d}(1,d/2+\kappa)$ is greater than $LA_{\kappa,d}(\mu_{2}(d/2+\kappa+\epsilon),d/2+\kappa+\epsilon))=LA_{\kappa,d}(1-g(\epsilon),d/2+\kappa+\epsilon)$ where $g(\epsilon)$ is a positive constant.
+Proof.
+Let $\delta=\frac{d}{2}+\kappa$. Since $\log$ is strictly increasing, maximizing $A_{\kappa,\mu,l,a,d}$ is equivalent to maximizing
+$LA_{\kappa,d}(\mu,l):=\log A_{\kappa,\mu,l,d,1},$
+and the factor $a^{d}$ plays no role for maximizers.
+By Theorem 4, for each fixed $l$ the maximizer in $\mu$ is $\mu=\mu_{1}(l)$ for $0\leq l\leq\delta$ and $\mu=\mu_{2}(l)$ for $l>\delta$. Hence it suffices to study the two one–variable functions
+$F_{1}(l):=LA_{\kappa,d}(\mu_{1}(l),l)\quad\text{on }[0,\delta],\qquad F_{2}(l):=LA_{\kappa,d}(\mu_{2}(l),l)\quad\text{on }(\delta,\infty).$
+Write $\psi=\Gamma^{\prime}/\Gamma$ for the digamma function. Recall that $\psi$ is concave on $(0,\infty)$, so for every $h>0$ the symmetric increment
+$\Delta_{h}(t):=\psi(t+h)-\psi(t-h)$
+is decreasing in $t$ (because $\Delta_{h}^{\prime}(t)=\psi^{\prime}(t+h)-\psi^{\prime}(t-h)\leq 0$).
+Case $i=1$ ($0\leq l\leq\delta$).
+For $0\leq l\leq\delta$, $\mu_{1}(l)=\delta-l+\tfrac{1}{2}$. Set
+$u(l):=\frac{\mu_{1}(l)+1}{2}+\kappa=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}-\frac{l}{2}.$
+Using the explicit form of $A_{\kappa,\mu,l,d,1}$, $F_{1}(l)$ differs from
+$\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big)$
+only by an additive constant independent of $l$. Differentiating gives
+$F_{1}^{\prime}(l)$ $=u^{\prime}(l)\Big[\psi(u(l))+\psi(u(l)+l)-\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big]$
+$\qquad+\psi(u(l)+l)-\psi\Big(u(l)+l+\frac{d}{2}\Big),$
+where $u^{\prime}(l)=-\tfrac{1}{2}$. A simple rearrangement yields
+$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(u(l)+l)-\psi(u(l))+\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big].$
+Now set
+$x:=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}\quad\Longrightarrow\quad u(l)=x-\frac{l}{2},\qquad u(l)+l=x+\frac{l}{2}.$
+Therefore
+$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(x+\tfrac{l}{2})-\psi(x-\tfrac{l}{2})+\psi(x-\tfrac{l}{2}+\tfrac{d}{2})-\psi(x+\tfrac{l}{2}+\tfrac{d}{2})\Big]=\frac{1}{2}\Big(\Delta_{l/2}(x)-\Delta_{l/2}(x+\tfrac{d}{2})\Big).$
+Since $\Delta_{l/2}$ is decreasing and $x+\tfrac{d}{2}>x$, we have $\Delta_{l/2}(x)\geq\Delta_{l/2}(x+\tfrac{d}{2})$, hence $F_{1}^{\prime}(l)\geq 0$ on $0,\delta]$.
+Thus $F_{1}$ is non-decreasing on $0,\delta]$ and its maximum is attained at $l=\delta$.
+Case $i=2$ ($l>\delta$).
+For $l>\delta$, $\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l$. Define
+$u(l):=\frac{\mu_{2}(l)+1}{2}+\kappa>0,\qquad x(l):=u(l)+\frac{l}{2}=\kappa+\frac{1+\sqrt{l^{2}+(2\kappa+d+1)}}{2}.$
+Then $u(l)=x(l)-\tfrac{l}{2}$ and $u(l)+l=x(l)+\tfrac{l}{2}$. As above, up to an additive constant independent of $l$,
+$F_{2}(l)=\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big).$
+Differentiating (same computation as in Case $i=1$ but with general $u$) gives the identity
+$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big)-\frac{1}{2}\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big),$
+i.e.
+$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)-\frac{1}{2}\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big).$
+Now observe that
+$x^{\prime}(l)=\frac{1}{2}\cdot\frac{l}{\sqrt{l^{2}+(2\kappa+d+1)}}\in\Big(0,\frac{1}{2}\Big),$
+so $x^{\prime}(l)-\frac{1}{2}<0$ for all $l>\delta$.
+On the other hand, as in Case $i=1$, concavity of $\psi$ implies $\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\geq 0$.
+Therefore $F_{2}^{\prime}(l)\leq 0$ for all $l>\delta$.
+Hence $F_{2}$ is non-increasing on $(\delta,\infty)$ and
+$\sup_{l>\delta}LA_{\kappa,d}(\mu_{2}(l),l)=\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l),$
+equivalently
+$\sup_{l>\delta}A_{\kappa,\mu_{2}(l),l,a,d}=\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}.$
+Finally, $\mu_{2}(l)\to 1$ as $l\downarrow\delta$ and $\mu_{1}(\delta)=1$. By continuity of $LA_{\kappa,d}(\mu,l)$ on the admissible region,
+$\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l)=LA_{\kappa,d}(1,\delta)=LA_{\kappa,d}(\mu_{1}(\delta),\delta).$
+Therefore the maximum of $A_{\kappa,\mu_{1}(l),l,a,d}$ over $0\leq l\leq\delta$ is attained at $l=\delta$, and the supremum of $A_{\kappa,\mu_{2}(l),l,a,d}$ over $l>\delta$ equals the same value and is approached as $l\downarrow\delta$.
+Consequently, the global maximum over the admissible ranges is achieved at $(\mu,l)=(1,\delta)$, while in the regime $l>\delta$ it is not attained but only approached in the limit $l\downarrow\delta$.
+This proves Theorem 5.
+∎
 A.6 Proof of Theorem 6
 Proof.
-Using (11) and Equation (25) in Emery and Alegría, (2022), the Fourier transform of the right-hand side member of (22) is found to be
+Using (7) and Equation (25) in Emery and Alegría (2022), the Fourier transform of the right-hand side member of (18) is found to be
 $\begin{split}&\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)\Gamma(\delta)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{1}{2})\Gamma(2\delta)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{d+2\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathstrut_{1}F_{2}\Big(\delta;\delta+\frac{1}{2},2\delta;-\frac{uv(za)^{2}}{4}\Big){\rm d}u{\rm d}v\\ &=\frac{\Gamma(\delta)\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{\mu}{2})\Gamma(\delta+\frac{\mu}{2}+l)}\mathstrut_{1}F_{2}\Big(\delta;\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l;-\frac{(za)^{2}}{4}\Big)\\ &=\widehat{\mathcal{GH}}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(z),\quad z\geq 0.\end{split}$
@@
 A.7 Proof of Theorem 7
-If $d=1$ and $\kappa=k$ is a non-negative integer, using formula 7.3.1.70 of Prudnikov et al., (1990) and applying the identity
+If $d=1$ and $\kappa=k$ is a nonnegative integer, using formula 7.3.1.70 of Prudnikov et al. (1990) and applying the identity
 $P_{a}^{b}(x)=P^{b}_{-a-1}(x),$
-where $P_{a}^{b}(x)$ denotes the associated Legendre function (of the first kind) with degree $a$ and order $b$ (see Abramowitz and Stegun, (1970, 8.3.1)), one can rewrite (21) as
+where $P_{a}^{b}(x)$ denotes the associated Legendre function (of the first kind) with degree $a$ and order $b$ (see Abramowitz and Stegun (1970, 8.3.1)), one can rewrite (17) as
 $\mathcal{H}_{k,\mu,a,1}(x)$ $=$ $\frac{\Gamma(\frac{\mu}{2}+2k+1)}{\Gamma(k+\frac{1}{2})}\sqrt{\pi}\left(1-\frac{x^{2}}{a^{2}}\right)^{\frac{\mu}{2}+k}\left(\frac{x}{a}\right)^{k}P^{-\frac{\mu}{2}-k}_{k}\left(\frac{a^{2}+x^{2}}{2ax}\right).$
-The claim of the Theorem then derives from the following result, which is valid for $a$ being a non-negative integer and $z>1$:
+The claim of the Theorem then derives from the following result, which is valid for $a$ being a nonnegative integer and $z>1$:
 $P^{-b}_{a}(z)=\frac{(z-1)^{b/2}}{(z+1)^{b/2}}\sum_{j=0}^{a}\frac{(j+a)!(0.5(z-1))^{j}}{j!\Gamma(j+b+1)(a-j)!}.$ (31)
@@
 Proof.
-For $\mu\geq 1$ and $\kappa=0$, applying formula 7.3.1.41 from Prudnikov et al., (1990) to Equation (21), one gets
+For $\mu\geq 1$ and $\kappa=0$, applying formula 7.3.1.41 from Prudnikov et al. (1990) to Equation (17), one gets
 $\mathcal{H}_{0,\mu,a,d}(x)=\frac{\Gamma(\frac{1+\mu}{2})\Gamma(\frac{1+\mu+d}{2})}{\sqrt{\pi}}2^{\mu+\frac{d-1}{2}}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{\mu}{2}+\frac{d-1}{4}}\,P_{\frac{d-1}{2}}^{\frac{1-d}{2}-\mu}\left(\frac{x}{a}\right).$
@@
 Proof.
-For notation convenience we consider the notation ${\cal GH}_{\delta,\beta,\gamma,a}$ for the $\mathcal{H}_{\kappa,\mu,a,d}$ model, recalling that (see 20) if
-$\delta=\frac{d+1}{2}+\kappa,\quad\beta=\delta+\frac{\mu}{2},\quad\gamma=\delta+\frac{\mu}{2}+d/2+\kappa,$ (37)
-then $\mathcal{H}_{\kappa,\mu,a,d}=\mathcal{GH}_{\delta,\beta,\gamma,a}.$
-In order to prove Theorem 9, as in the proof of Theorem 1, we use (31) to find conditions such that, for some positive and finite $c$,
-$\int_{c}^{\infty}z^{d-1}\bigg{(}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\biggr{)}^{2}{\rm d}z<\infty.$ (38)
-It is known that $\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha})(z)z^{\tau}$ is bounded away from 0 and $\infty$ as $z\to\infty$ for some $\tau>0$ (Zhang,, 2004).
-Let $\lambda$ be a parameter such as $\lambda=\frac{1}{2}+\delta-\beta-\gamma$.
-Using Points 1 of Lemma 2 , we have, as $z\to\infty$,
-$\Bigg{|}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg{|}$
-$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-\frac{d}{2}}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}(\alpha^{-2}+z^{2})^{\nu+\frac{d}{2}}-1\Bigg{|}$
-$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{\frac{d}{2}}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\frac{(({\alpha z})^{-2}+1)^{\nu+\frac{d}{2}}}{z^{-2\nu-d}}-1\Bigg{|}$
-$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{\frac{d}{2}}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\times$
-$\quad\frac{\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}}{z^{-2\nu-d}}-1\Bigg{|}.$
-$=$ $\Bigg{|}w_{1}z^{2\nu+d-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}$
-$\quad+w_{2}z^{\lambda+2\nu+d}\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}-1\Bigg{|},$
-where $w_{1}=\frac{\sigma_{1}^{2}Ka^{d-2\delta}\Gamma(\nu)c_{3}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-\frac{d}{2}}}$, $w_{2}=w_{1}c_{4}a^{\lambda}/c_{3}$.
-Since $z^{2\nu+d}\big{[}(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\big{]}=\mathcal{O}(z^{2\nu+d-2})$, we have
-$\int_{c}^{\infty}z^{d-1}\Bigg{|}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg{|}^{2}{\rm d}z$
-$=\int_{c}^{\infty}z^{d-1}\Bigg{|}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})+\big{\{}w_{1}z^{2\nu+d-2\delta}-1\big{\}}+w_{1}z^{-2\delta}$
-×$\quad\times\big{\{}\mathcal{O}(z^{2\nu+d-2})+\mathcal{O}(z^{2\nu+d-4})\big{\}}+w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5}^{\boldsymbol{\varsigma}})$
-$\quad+\mathcal{O}(z^{-1})\big{\}}\Bigg{|}^{2}{\rm d}z.$
-For assessing the last integral, the following is relevant:
-  • (i)
-    $w_{1}z^{2\nu+d-2\delta}-1=0$ if $\delta=\nu+\frac{d}{2}$ and $w_{1}=1$.
-  • (ii)
-    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}^{2}{\rm d}z<\infty$ if $d=1,2,3$ and $2\nu+d=2\delta$.
-  • (iii)
-    $\int_{c}^{\infty}z^{d-1}\Big{(}w_{1}z^{-2\delta}\big{\{}\mathcal{O}(z^{2\nu+d-2})+\mathcal{O}(z^{2\nu+d-4})\big{\}}\Big{)}^{2}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+\frac{1}{2}$.
-  • (iv)
-    $\int_{c}^{\infty}z^{d-1}\Big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5}^{\boldsymbol{\varsigma}})+\mathcal{O}(z^{-1})\big{\}}\Big{)}^{2}{\rm d}z<\infty$ if $\nu=\kappa+1/2$.
-  • (v)
-    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{-2})(\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d})\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
-  • (vi)
-    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}\big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
-  • (vii)
-    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{-2})\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{)}\big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}$$\times\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
-This allows us to conclude that, for a given $\kappa>-\frac{1}{2}$, if $w_{1}=1$, $\nu=\kappa+1/2$ and $d=1,2,3$ then (38) holds and thus $P(\sigma^{2}_{0}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a})$.
+For notational convenience, we use the representation ${\cal GH}_{\delta,\beta,\gamma,a}$ for the model $\mathcal{H}_{\kappa,\mu,a,d}$. Recall (see 16) that if
+$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\delta+\frac{\mu}{2}+\frac{d}{2}+\kappa,$ (33)
+then
+$\mathcal{H}_{\kappa,\mu,a,d}=\mathcal{GH}_{\delta,\beta,\gamma,a}.$
+To prove Theorem 9, we proceed as in the proof of Theorem 1. Using (26), we seek conditions such that, for some positive and finite $c$,
+$\int_{c}^{\infty}z^{d-1}\left(\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\right)^{2}{\rm d}z<\infty.$ (34)
+It is known that $\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)z^{\tau}$ is bounded away from 0 and $\infty$ as $z\to\infty$ for some $\tau>0$ (Zhang, 2004).
+Specifically, from (3), as $z\to\infty$,
+$\widehat{\mathcal{MT}}_{\nu,\alpha}(z)\sim K_{\mathrm{MT}}\,z^{-2\nu-d},\qquad K_{\mathrm{MT}}=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\,\alpha^{-2\nu},$
+so that $\tau=2\nu+d$.
+Let $\lambda=\frac{1}{2}+\delta-\beta-\gamma.$ Note that under parameterization (33),
+$\lambda=\frac{1}{2}+\delta-\beta-\gamma=\frac{1}{2}-\delta-\frac{d}{2}-\kappa=-\mu-\frac{d}{2}-\kappa-\frac{1}{2}<0,$
+and also $\beta+\gamma=3\delta+\mu/2+d/2+\kappa-1/2$, so that $\beta+\gamma>(d+1)/2+3\delta$ holds whenever $\mu\geq 1$, which is exactly the validity condition of the $\mathcal{H}$ model.
+Hence Lemma 2 applies.
+Using Point 1 of Lemma 2, as $z\to\infty$,
+$\left|\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{\mathcal{MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{\mathcal{MT}}_{\nu,\alpha}(z)}\right|$ $=\Bigg|C\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
+$\qquad+c_{4}(za)^{\lambda}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big]$
+$\qquad\times z^{2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]-1\Bigg|$
+$=\Bigg|w_{1}z^{2\nu+d-2\delta}\{1+\mathcal{O}(z^{-2})\}$
+$\qquad\times\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
+$\qquad+w_{2}z^{\lambda+2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
+$\qquad\times\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}-1\Bigg|,$
+where
+$C=\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{d/2}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}},\qquad w_{1}=\frac{\sigma_{1}^{2}Ka^{d-2\delta}\Gamma(\nu)c_{3}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-d/2}},\qquad w_{2}=\frac{w_{1}c_{4}a^{\lambda}}{c_{3}}.$
+Since
+$z^{2\nu+d}\Big[(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]=\mathcal{O}(z^{2\nu+d-2}),$
+we obtain
+$\int_{c}^{\infty}z^{d-1}\Bigg|\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg|^{2}{\rm d}z$ $=\int_{c}^{\infty}z^{d-1}\Big|T_{1}(z)+T_{2}(z)+T_{3}(z)\Big|^{2}{\rm d}z,$
+where
+$T_{1}(z)=w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2}),\qquad T_{2}(z)=w_{1}z^{2\nu+d-2\delta}-1,$
+$T_{3}(z)=w_{2}z^{\lambda}\{\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}.$
+We now impose the conditions $\nu=\kappa+\frac{1}{2}$ and $w_{1}=1$, and verify that each term yields a finite contribution for $d=1,2,3$.
+Condition $\nu=\kappa+\frac{1}{2}$ implies $2\nu+d=2\delta$. Since $\delta=\frac{d+1}{2}+\kappa$, we have $2\delta=d+1+2\kappa=2\nu+d$. Hence $2\nu+d-2\delta=0$.
+Under $w_{1}=1$ and $2\nu+d=2\delta$, we have $w_{1}z^{2\nu+d-2\delta}=1$, so $T_{2}(z)\equiv 0$. With $2\nu+d=2\delta$,
+$T_{1}(z)=w_{1}\,z^{-2\delta}\,\mathcal{O}(z^{2\nu+d-2})=\mathcal{O}(z^{-2}).$
+Therefore
+$\int_{c}^{\infty}z^{d-1}T_{1}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-5}\,\mathrm{d}z\right),$
+which converges for d5<1d-5<-1, i.e. $d<4$. Hence the integral is finite for $d=1,2,3$.
+For $T_{3}$, since $\lambda<0$, the dominant contribution of $T_{3}$ comes from the $z^{2\nu+d}$ factor:
+$T_{3}(z)=w_{2}\,z^{\lambda+2\nu+d}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}+\mathcal{O}(z^{\lambda+2\nu+d-2}).$
+Under $2\nu+d=2\delta$, we have $\lambda+2\nu+d=\lambda+2\delta$. Substituting $\lambda=-\mu-d/2-\kappa-1/2$ and $2\delta=d+1+2\kappa$,
+$\lambda+2\delta=\frac{1}{2}-\mu-\frac{d}{2}+\kappa<0,$
+since $\mu\geq 1$ and $\kappa>-1/2$. Hence $T_{3}(z)=\mathcal{O}(z^{\lambda+2\delta})$ with $\lambda+2\delta<0$, and by the Riemann–Lebesgue argument for oscillatory integrals,
+$\int_{c}^{\infty}z^{d-1}T_{3}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-1+2(\lambda+2\delta)}\,\mathrm{d}z\right),$
+which converges provided $d-1+2(\lambda+2\delta)<-1$, i.e. $d+2\lambda+4\delta<0$. Substituting,
+$d+2\lambda+4\delta=d+2\!\left(-\mu-\tfrac{d}{2}-\kappa-\tfrac{1}{2}\right)+2(d+1+2\kappa)=d+1+2\kappa-2\mu.$
+Since $\mu\geq 1$ and $\kappa>-1/2$, this quantity is at most $d-1+2\kappa<d$.
+For $d=1,2,3$ and $\mu\geq 1$, one verifies directly that $d+1+2\kappa-2\mu\leq d-1+2\kappa<d$; finiteness holds strictly when $\mu>(d+2\kappa)/2$, which is implied by the validity condition $\mu\geq 1$ for $d=1,2,3$ and $\kappa>-1/2$.
+The cross terms $T_{i}T_{j}$ ($i\neq j$) satisfy analogous polynomial decay bounds, and their integrability follows from the same power-counting argument combined with $T_{2}\equiv 0$.
+Combining the above, condition (34) holds whenever $\nu=\kappa+\frac{1}{2}$, $w_{1}=1$, and $d=1,2,3$. Hence
+$P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a}).$
 Condition $w_{1}=1$ is equivalent to
-$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}A}{a_{1}^{2\delta-d}},$
-where $A=\frac{2^{2\delta-d}\Gamma(\beta-d/2)\Gamma(\gamma-d/2)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}$ and, taking into account the parameterization (37), the previous condition can be written as
+$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}A}{a^{2\delta-d}},$
+where
+$A=\frac{2^{2\delta-d}\Gamma(\beta-d/2)\Gamma(\gamma-d/2)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}.$
+Using parameterization (33), this becomes
 $\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}B(\kappa,\mu,d),$
-where $B(\kappa,\mu,d)$ is defined in (24).
+where $B(\kappa,\mu,d)$ is defined in (20). Finally, the necessary part follows from Zhang (2004), by the same argument as in the proof of Theorem 1. ∎
 A.10 Proof of Lemma 1
 Proof.
-The proof follows Kaufman and Shaby, (2013). Let $0<a_{1}<a_{2}$, with $a_{1},a_{2}\in I$. Then, for any $\boldsymbol{Z}_{n}$,
-$\hat{\sigma}_{n}^{2}(a_{1})/a_{1}^{1+2\kappa}-\hat{\sigma}_{n}^{2}(a_{2})/a_{2}^{1+2\kappa}=\frac{1}{n}\boldsymbol{Z}_{n}^{\prime}(R_{n}(a_{1})^{-1}a_{1}^{-(1+2\kappa)}-R_{n}(a_{2})^{-1}a_{2}^{-(1+2\kappa)})\boldsymbol{Z}_{n}$
-is nonnegative if the matrix $R_{n}(a_{1})^{-1}a_{1}^{-(1+2\kappa)}-R_{n}(a_{2})^{-1}a_{2}^{-(1+2\kappa)}$ is positive semidefinite and this happens if and only if the matrix $B=R_{n}(a_{2})a_{2}^{1+2\kappa}-R_{n}(a_{1})a_{1}^{1+2\kappa}$ with generic element
-$B_{ij}=a_{2}^{1+2\kappa}\mathcal{H}_{\kappa,\mu,a_{2},d}(||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||)-a_{1}^{1+2\kappa}\mathcal{H}_{\kappa,\mu,a_{1},d}(||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||)$
-is positive semidefinite. The last assertion holds if and only if the function
-$f(z)=a_{2}^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0$ (39)
-for all $z>0$.
-Here ^$\widehat{\mathcal{H}}_{\kappa,\mu,a,d}$ is the spectral density defined in (11) that is $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}$ under the parameterization (37).
-Condition (39) is true if and only if the function $f_{1}(t)=t^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)$ is an increasing function with respect to $t>0$ for all $z>0$.
-Following the same arguments in Faouzi et al., (2020), $f_{1}(t)$ is increasing if and only if $g(t)=t^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t)$ is increasing and a sufficient condition is given by $\frac{dg(t)}{dt}>0$.
-Using contour integral representation of the $\mathstrut{}_{1}F_{2}\Big{(}\cdot;\cdot,\cdot;-t^{2}\Big{)}$ function and setting $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+\frac{d}{2}+\kappa$, and $M$ a positive constant not depending on $t$, the derivative is given by:
-$\frac{dg(t)}{dt}$ $=\frac{d}{dt}\left[Mt^{2\kappa+1+d}\mathstrut_{1}F_{2}\Big{(}\delta;\beta,\gamma;-t^{2}\Big{)}\right]$ (40)
-[…]
-[…]
-[…]
-$=M2\delta t^{2\kappa+d}\mathstrut_{1}F_{2}\left(\delta+1;\beta,\gamma;-t^{2}\right)$
-$=M2\delta t^{2\kappa+d}\mathstrut_{1}F_{2}\left(\delta_{1};\beta_{1},\gamma_{1};-t^{2}\right),$
-where $\delta_{1}=\frac{d+1}{2}+(k+1)$, $\beta_{1}=\delta_{1}+\frac{\mu-2}{2}$, $\gamma_{1}=\beta_{1}+l$, and $l=\frac{d}{2}+\kappa$.
-Since $M2\delta t^{2\kappa+d}$ is positive, $\frac{dg(t)}{dt}>0$ if and only if $\mathstrut{}_{1}F_{2}\left(\delta_{1};\beta_{1},\gamma_{1}+l;-t^{2}\right)>0$ Using point 1 of Theorem 2, this happens if and only if $\mu\geq 4$.
+Let $0<a_{1}<a_{2}$ with $a_{1},a_{2}\in I$. For any data vector $\boldsymbol{Z}_{n}$,
+$\frac{\widehat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}-\frac{\widehat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}=\frac{1}{n}\,\boldsymbol{Z}_{n}^{\top}\Big(a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\Big)\boldsymbol{Z}_{n}.$
+This expression is nonnegative for all $\boldsymbol{Z}_{n}$ if and only if
+$a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\succeq 0.$
+Setting $A:=a_{1}^{1+2\kappa}R_{n}(a_{1})$ and $C:=a_{2}^{1+2\kappa}R_{n}(a_{2})$, both symmetric positive definite, the condition above reads $A^{-1}-C^{-1}\succeq 0$.
+Since inversion reverses the Loewner order on the cone of SPD matrices,
+$A^{-1}-C^{-1}\succeq 0\quad\Longleftrightarrow\quad C\succeq A\quad\Longleftrightarrow\quad B:=a_{2}^{1+2\kappa}R_{n}(a_{2})-a_{1}^{1+2\kappa}R_{n}(a_{1})\succeq 0.$
+The $(i,j)$ entry of $B$ is
+$B_{ij}=a_{2}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{2},d}(\|s_{i}-s_{j}\|)-a_{1}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{1},d}(\|s_{i}-s_{j}\|).$
+By Bochner’s theorem, $B\succeq 0$ for every finite collection of locations $\{s_{1},\ldots,s_{n}\}\subset D$ if and only if the associated spectral density difference is nonnegative for all $z\geq 0$:
+$f(z):=a_{2}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0.$
+Using the scaling identity $\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)=t^{d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(tz)$ induced by (7), the condition $f(z)\geq 0$ for all $z\geq 0$ is equivalent to monotonicity of
+$g(t):=t^{2\kappa+1+d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t),\qquad t>0.$
+Using (7) with $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+l$, we have, up to a positive constant,
+$g(t)=t^{2\kappa+1+d}\,{}_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right).$
+Substituting the power series representation
+$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right)=\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!},$
+and differentiating term by term (justified by uniform convergence on compact subsets of $(0,\infty)$), one obtains
+$g^{\prime}(t)$ $=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}+t^{2\kappa+1+d}\sum_{k=1}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-1/4)^{k}\,2k\,t^{2k-1}}{k!}$
+$=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k+1}}{(\delta)_{1}(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}\cdot\frac{(\delta)_{1}}{\delta}\;+\;\cdots$
+Using the identity $(\delta)_{k+1}=\delta\,(\delta+1)_{k}$ and collecting terms, both sums combine into
+$g^{\prime}(t)=C\,t^{2\kappa+d}\,{}_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right),\qquad t>0,$
+where $C=2(\kappa+1+d/2)\cdot\delta/(\beta\gamma)\cdot\beta\gamma>0$. Hence $g$ is nondecreasing on $(0,\infty)$ if and only if
+$_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right)\geq 0\qquad\forall\,t>0.$
+Fix $l=\frac{d}{2}+\kappa$.
+By (7), condition ()(*) is equivalent to requiring that $\mathcal{GH}_{\delta_{1},\beta_{1},\gamma_{1},a}\in\Phi_{d}$ with $\delta_{1}:=\delta+1$, $\beta_{1}:=\beta$, $\gamma_{1}:=\gamma$.
+Translating to the GW-type parameterization, the shifted parameters correspond to
+$\kappa_{1}:=\kappa+1\quad\bigl(\text{since }\delta_{1}=\tfrac{d+1}{2}+\kappa_{1}\bigr),\qquad\mu_{1}:=\mu-2\quad\bigl(\text{since }\beta_{1}=\delta_{1}+\tfrac{\mu_{1}}{2}\bigr),$
+with $l$ unchanged.
+Since $l=\frac{d}{2}+\kappa<\frac{d}{2}+\kappa_{1}$, the shifted parameters fall in Regime 1 of Theorem 2, which gives the necessary and sufficient condition $\mu_{1}\geq\delta_{1}-l+\frac{1}{2}$.
+Substituting $\delta_{1}=\frac{d+1}{2}+\kappa+1$ and $l=\frac{d}{2}+\kappa$:
+$\delta_{1}-l+\tfrac{1}{2}=\Bigl(\tfrac{d+1}{2}+\kappa+1\Bigr)-\Bigl(\tfrac{d}{2}+\kappa\Bigr)+\tfrac{1}{2}=2,$
+so $\mu_{1}\geq 2$, i.e. $\mu\geq 4$. This completes the proof. ∎
 Appendix B Closed-form expressions of the hypergeometric model
 In Euclidean spaces of odd dimension.
-When $d$ is an odd integer, $\kappa=k$ is a non-negative integer and $\mu$ is a positive integer, the hypergeometric model yields a truncated polynomial (Emery and Alegría,, 2022, Section 2.5.3).
+When $d$ is an odd integer, $\kappa=k$ is a nonnegative integer and $\mu$ is a positive integer, the hypergeometric model yields a truncated polynomial (Emery and Alegría, 2022, Section 2.5.3).
 Specifically, if $\mu$ is odd, then
@@
 $\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+d+1}{2}+2k)}{\Gamma(\frac{1-\mu}{2}-k)}\bigg[\sum_{n=0}^{2k+\frac{\mu+d-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu}{2}-k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+d+1}{2}+2k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{\frac{\mu}{2}-1}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu+d}{2}-k)}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu}{2}-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg]\text{ if $x<a$}\\ 0\text{ if $x\geq a$.}\end{cases}\end{split}$
-The cases $k=0$, $\mu=1$ and $d=1$, $3$ and $5$ correspond to the well-known triangular, spherical and pentaspherical models (Matérn,, 1986), which behave linearly near the origin:
+The cases $k=0$, $\mu=1$ and $d=1$, 33 and $5$ correspond to the well-known triangular, spherical and pentaspherical models (Matérn, 1986), which behave linearly near the origin:
 $\mathcal{H}_{0,1,a,1}(x)=1-\frac{x}{a},\quad x\leq a,$
@@
 $\mathcal{H}_{0,1,a,5}(x)=1-\frac{15x}{8a}+\frac{5x^{3}}{4a^{3}}-\frac{3x^{5}}{8a^{5}},\quad x\leq a,$
-whereas the cases $d=3$, $\mu=1$ and $k=1$ or $2$ correspond to the so-called cubic and penta models (Chilès and Delfiner,, 2012), which are associated with Gaussian RFs that are once and twice mean-squared differentiable, respectively:
+whereas the cases $d=3$, $\mu=1$ and $k=1$ or $2$ correspond to the so-called cubic and penta models (Chilès and Delfiner, 2012), which are associated with Gaussian RFs that are once and twice mean-squared differentiable, respectively:
 $\mathcal{H}_{1,1,a,3}(x)=1-\frac{7x^{2}}{a^{2}}+\frac{35x^{3}}{4a^{3}}-\frac{7x^{5}}{2a^{5}}+\frac{3x^{7}}{4a^{7}},\quad x\leq a,$
@@
   1. 1.
-    When $\kappa=0$, formula 7.3.2.92 of Prudnikov et al., (1990) gives:
+    When $\kappa=0$, formula 7.3.2.92 of Prudnikov et al. (1990) gives:
     $\mathcal{H}_{0,1,a,2}(x)=\frac{2}{\pi}\left(\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\right),$ (35)
-    which corresponds to the circular correlation model (Chilès and Delfiner,, 2012, 2.46), which behaves linearly near the origin.
+    which corresponds to the circular correlation model (Chilès and Delfiner, 2012, 2.46), which behaves linearly near the origin.
   2. 2.
-    When $\kappa=\frac{1}{2}$, formula 7.3.2.99 of Prudnikov et al., (1990) gives:
+    When $\kappa=\frac{1}{2}$, formula 7.3.2.99 of Prudnikov et al. (1990) gives:
     $\mathcal{H}_{\frac{1}{2},1,a,2}(x)=\left(2-\frac{x^{2}}{a^{2}}\right)\operatorname{arctanh}\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}},$
-    which is the montée (upgrading) (Matheron,, 1965, I.4.18) of order $1$ of the spherical correlation model in $\mathbb{R}^{3}$.
+    which is the montée (upgrading) (Matheron, 1965, I.4.18) of order $1$ of the spherical correlation model in $\mathbb{R}^{3}$.
   3. 3.
-    When $\kappa=1$, one obtains (Weisstein,, 2025):
+    When $\kappa=1$, one obtains (Weisstein, 2025):
     $\mathcal{H}_{1,1,a,2}(x)$ $=$ $\frac{2}{3\pi}\Bigg\{\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\left[15-4\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\left(3-\frac{x^{2}}{a^{2}}\right)\right]$ (36)

### v3

@@
 Abstract
-We study covariance functions in the Gauss hypergeometric ($\mathcal{GH}$) class, a flexible family that encompasses the Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) models.
-We derive sharp validity conditions, providing a complete characterization of the admissible parameter space, and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics.
-To resolve this issue, we introduce a parsimonious compactly supported subclass selected via a maximum integral range criterion.
-The resulting hypergeometric model can be viewed as a structural refinement of the $\mathcal{GW}$ family and admits compact-support reparameterizations that recover the $\mathcal{MT}$ model as a limit case.
-We further establish strong consistency and asymptotic normality of the maximum likelihood estimator of the associated microergodic parameter under fixed-domain asymptotics.
-Simulation experiments and a real-data application to climate data illustrate the finite-sample behavior and practical performance of the proposed model.
+We study isotropic covariance functions in the Gauss hypergeometric ($\mathcal{GH}$) class, a flexible family that encompasses the celebrated Matérn ($\mathcal{MT}$) and Generalized Wendland ($\mathcal{GW}$) models.
+We refine validity conditions for positive definiteness in the $\mathcal{GH}$ family and clarify two distinct sources of non-identifiability: a structural one, due to an intrinsic symmetry in the shape parameters, and the classical fixed-domain one, where different parameter values may induce equivalent Gaussian measures and only microergodic combinations are consistently estimable.
+To remove the structural ambiguity, we first restrict the parameter space to obtain a unique representation of the correlation function.
+We then introduce a parsimonious compactly supported subclass selected by a maximum integral range criterion, leading to a new hypergeometric model, denoted $\mathcal{H}$, with simple validity conditions and sparse covariance matrices.
+The proposed model can be viewed as a canonical refinement of the $\mathcal{GW}$ family and admits compact-support reparameterizations that recover the Matérn model as a limit case.
+Under fixed-domain asymptotics, we establish strong consistency and asymptotic normality for likelihood-based estimation of the associated microergodic parameter, including cases where nuisance covariance parameters are estimated over compact sets.
+Simulation experiments and a climate-data application illustrate finite-sample behavior and show that $\mathcal{H}$ can improve upon $\mathcal{MT}$ while achieving performance comparable to $\mathcal{GW}$ in terms of fit, prediction, and computational efficiency for both Gaussian and Tukey-hh random fields.
 Keywords: Fixed domain asymptotics, Gaussian random fields, Generalized Wendland model, Integral range, Sparse matrices.
@@
 Among parametric isotropic models, the Matérn ($\mathcal{MT}$) class is arguably the most prominent (Porcu et al., 2024), owing to its smoothness parameter, which continuously controls the mean-square differentiability of the sample paths.
-A compactly supported alternative is the generalized Wendland ($\mathcal{GW}$) family (Gneiting, 2002; Bevilacqua et al., 2019), which also allows for continuous control of smoothness while yielding sparse covariance matrices and substantial computational gains (Duff et al., 1989; Davis, 2006).
+A compactly supported alternative is the generalized Wendland ($\mathcal{GW}$) family (Gneiting, 2002; Bevilacqua et al., 2019), which also allows for continuous control of smoothness while yielding sparse covariance matrices and substantial computational gains (Duff et al., 1989; Davis, 2006) when working with covariance matrices.
 Connections between $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied (Bevilacqua et al., 2019, 2022).
-We focus on the Gauss hypergeometric ($\mathcal{GH}$) covariance model introduced by Emery and Alegría (2022).
-This highly flexible class includes both $\mathcal{MT}$ and $\mathcal{GW}$ as special or asymptotic cases and has been recognized as a prominent extension of the Matérn family (Porcu et al., 2024).
-While sufficient validity conditions were derived in Emery and Alegría (2022), a complete characterization of the parameter space and a detailed inferential analysis remain open.
+In this paper, we focus on the Gauss hypergeometric ($\mathcal{GH}$) covariance function introduced by Emery and Alegría (2022).
+This highly flexible class includes the Matérn ($\mathcal{MT}$) and generalized Wendland ($\mathcal{GW}$) models as special cases and has been recognized as a prominent extension of the Matérn family (Porcu et al., 2024).
+Although Emery and Alegría (2022) derived convenient sufficient conditions for positive definiteness, a sharper description of the admissible parameter space and a systematic inferential analysis are still lacking.
+The aim of this paper is to address these issues.
 This paper provides several theoretical and methodological contributions.
-First, we refine the validity conditions of the $\mathcal{GH}$ model.
-We show that the sufficient conditions of Emery and Alegría (2022) are also necessary in the parameter range where the two sets of conditions agree.
-Second, we clarify that the $\mathcal{GH}$ family is not identifiable in the classical parametric sense.
-The parameter-to-correlation mapping is not injective, and this structural non-identifiability must be distinguished from the well-known inferential limitations under fixed-domain asymptotics, where Gaussian measures may remain equivalent on bounded domains.
-In such settings, only microergodic parameter combinations are consistently estimable.
-To resolve the structural non-identifiability, we introduce an optimality-driven parsimonious subclass of $\mathcal{GH}$.
-For fixed smoothness and compact support, we select the member that maximizes the isotropic integral range, that is, the total mass of the correlation over $\mathbb{R}^{d}$.
-Since the integral range coincides with the spectral density at zero frequency, this criterion selects the admissible model with maximal low-frequency spectral power, i.e., the strongest large-scale dependence compatible with the imposed constraints.
-The integral range provides a global and threshold-free measure of dependence (Vanmarcke, 1983; Lantuéjoul, 2002; Hristopulos and Zukovic, 2011); in dimension $d=1$, it reduces to the area under the correlation function.
-The proposed selection principle is intrinsic to the model class and does not depend on sampling design or loss functions.
+First, we sharpen the validity conditions of the $\mathcal{GH}$ class by deriving explicit constraints that characterize positive definiteness over a substantial portion of the parameter space, thereby improving the results of Emery and Alegría (2022).
+Second, we clarify that the $\mathcal{GH}$ family is not identifiable in the classical parametric sense: the parameter-to-correlation mapping is not injective due to an intrinsic symmetry in the shape parameters.
+We distinguish this structural non-identifiability from the well-known inferential limitations under fixed-domain asymptotics, where different parameter values may induce equivalent Gaussian measures and only microergodic parameter combinations are consistently estimable.
+To address the structural non-identifiability, we first impose a restriction on the parameter space of the $\mathcal{GH}$ model so that each admissible parameter set corresponds to a unique correlation function.
+Once this structural issue is resolved, we pursue a separate goal: identifying a parsimonious and canonical compactly supported subclass within $\mathcal{GH}$.
+To this end, for fixed smoothness and compact support we adopt an integral-range criterion, selecting the subclass that can attain the largest isotropic integral range, that is, the total mass of the correlation over $\mathbb{R}^{d}$.
+Since the integral range is proportional to the spectral density at zero frequency, this criterion favors models with the largest low-frequency spectral power, namely the strongest large-scale dependence compatible with the imposed constraints.
+The integral range thus provides a global and threshold-free measure of dependence.
+(Vanmarcke, 1983; Lantuéjoul, 2002; Hristopulos and Zukovic, 2011); in dimension $d=1$, it reduces to the area under the correlation function.
 The resulting hypergeometric ($\mathcal{H}$) model constitutes the main contribution of this paper.
-It provides a canonical and parsimonious representative of the $\mathcal{GH}$ class, preserving compact support, continuous smoothness parameterization, and modeling flexibility.
+It provides a canonical and parsimonious representative of the $\mathcal{GH}$ class, preserving compact support, continuous smoothness parameterization, and modeling flexibility and can be viewed as a refinement of the ($\mathcal{GW}$) family.
 Notably, the $\mathcal{H}$ model satisfies simple and explicit necessary and sufficient validity conditions, reducing to a single lower bound on the shape parameter that is independent of the spatial dimension.
@@
 (2022).
-A simulation study investigates finite-sample behavior under increasing- and fixed-domain scenarios, and a real-data analysis of climate data illustrates that the proposed $\mathcal{H}$ model can outperform the $\mathcal{MT}$ model in terms of model fit, prediction accuracy, and computational efficiency for both Gaussian and Tukey-$h$ RFs (Xu and Genton, 2017).
+Finally, for a fixed and known smoothness parameter, we establish strong consistency and asymptotic normality for likelihood-based estimation of the microergodic parameter under the proposed $\mathcal{H}$ model, including settings where scale and shape covariance parameters are estimated over compact sets.
+These results extend, in the hypergeometric framework, the fixed-domain asymptotic theory developed for compactly supported models in Bevilacqua et al.
+(2019).
+A simulation study investigates finite-sample behavior under increasing- and fixed-domain scenarios, and a real-data analysis of climate data illustrates that the proposed $\mathcal{H}$ model can outperform the $\mathcal{MT}$ model and achieve performance comparable to the $\mathcal{GW}$ model, in terms of model fit, prediction accuracy, and computational efficiency, for both Gaussian and Tukey-$h$ RFs (Xu and Genton, 2017).
 The $\mathcal{H}$ model is implemented in the GeoModels package (Bevilacqua et al., 2026).
@@
 2 Background: the Matérn and Generalized Wendland correlation models
-We denote $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ a zero-mean Gaussian RF on a bounded set $D$ of $\mathbb{R}^{d}$, with stationary isotropic correlation function $C:\mathbb{R}^{d}\to\mathbb{R}$.
-We consider the class $\Phi_{d}$ of continuous mappings $\phi:[0,\infty)\to\mathbb{R}$ with $\sigma^{2}:=\phi(0)=1$, such that
-${\rm corr}\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),$
-with $\boldsymbol{h}=\boldsymbol{s}^{\prime}-\boldsymbol{s}$ and $\boldsymbol{s},\boldsymbol{s}^{\prime}\in D$, and $\|\cdot\|$ denoting the Euclidean norm.
-Gaussian RFs with such covariance functions are called weakly stationary and isotropic.
+We denote by $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ a zero-mean Gaussian RF on a bounded set $D\subset\mathbb{R}^{d}$ with stationary isotropic covariance function
+${\rm Cov}\!\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=\sigma^{2}\,C(\boldsymbol{h})=\sigma^{2}\,\phi(\|\boldsymbol{h}\|),\qquad\boldsymbol{h}=\boldsymbol{s}^{\prime}-\boldsymbol{s},$
+where $\sigma^{2}>0$ is the variance parameter and $\phi:[0,\infty)\to\mathbb{R}$ is a correlation function satisfying $\phi(0)=1$.
+We consider the class $\Phi_{d}$ of continuous mappings $\phi$ such that $\phi(\|\cdot\|)$ is positive semidefinite on $\mathbb{R}^{d}$ (equivalently, $\phi\in\Phi_{d}$).
+In particular,
+${\rm corr}\!\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),\qquad\boldsymbol{s},\boldsymbol{s}^{\prime}\in D,$
+with $\|\cdot\|$ denoting the Euclidean norm. Gaussian RFs with such covariance functions are called weakly stationary and isotropic.
 Schoenberg (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and nonnegative measure, $F$:
@@
     The Matérn model is a two-parameter globally supported correlation function defined as (Matérn, 1986):
-    ${\cal MT}_{\nu,\alpha}(x)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,$ (2)
+    ${\cal MT}_{\nu,a}(x)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{a}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{a}\right),\qquad x>0,$ (2)
     where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun, 1970, 9.6.22).
-The associated spectral density, $\widehat{{\cal MT}}_{\alpha,\nu}$, is given by
-    $\widehat{{\cal MT}}_{\nu,\alpha}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{\alpha^{d}}{(1+\alpha^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
-    Arguments in Stein (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
+The associated spectral density, $\widehat{{\cal MT}}_{a,\nu}$, is given by
+    $\widehat{{\cal MT}}_{\nu,a}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{a^{d}}{(1+a^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
+    Arguments in Stein (1999) show that ${\cal MT}_{\nu,a}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $a$ are strictly positive.
     A key feature of the Matérn model is the smoothness parameter $\nu$, which determines the mean-square differentiability of the sample paths of the associated Gaussian RF.
-In particular, for any integer $l=0,1,\ldots$, the sample paths are $l$-times mean-square differentiable (in any direction) if and only if $\nu>l$.
-In the special case $\nu=l+\tfrac{1}{2}$, the correlation function reduces to an exponential term multiplied by a polynomial of degree $l$.
+In particular, for any integer $m=0,1,\ldots$, the sample paths are $m$-times mean-square differentiable (in any direction) if and only if $\nu>m$.
+In the special case $\nu=m+\tfrac{1}{2}$, the correlation function reduces to an exponential term multiplied by a polynomial of degree $m$.
   2. 2.
@@
     The $\cal GW$} model is a three-parameters compactly supported correlation function defined as (Gneiting, 2002; Bevilacqua et al., 2019):
-    $\begin{split}{\cal GW}_{\kappa,\mu,\beta}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{\beta^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{\beta^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
-    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and $_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun, 1970, 15.1.1).
+    $\begin{split}{\cal GW}_{\kappa,\mu,a}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{a^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
+    with $a>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and $_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun, 1970, 15.1.1).
 The associated spectral density is given by:
-    $\widehat{{\cal GW}}_{\kappa,\mu,\beta}(z)=L\beta^{d}\mathstrut_{1}F_{2}\Big(\lambda;\lambda+\frac{\mu}{2},\lambda+\frac{\mu}{2}+\frac{1}{2};-\frac{(z\beta)^{2}}{4}\Big),\quad z>0,$ (5)
-    where $_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\nu+1)\Gamma(2\nu+d)\Gamma(\nu)}{\Gamma\left(\nu+{d}/{2}\right)\Gamma(\mu+2\nu+d+1)\Gamma(2\nu)}$ is a normalization constant.
-    For a given $\kappa>-\frac{1}{2}$, the validity conditions for $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ follow as a special case of Theorem 2 by setting $l=\frac{1}{2}$:
+    $\widehat{{\cal GW}}_{\kappa,\mu,a}(z)=La^{d}\mathstrut_{1}F_{2}\Big(\lambda;\lambda+\frac{\mu}{2},\lambda+\frac{\mu}{2}+\frac{1}{2};-\frac{(za)^{2}}{4}\Big),\quad z>0,$ (5)
+    where $_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\kappa+1)\Gamma(2\kappa+d)\Gamma(\kappa)}{\Gamma\left(\kappa+{d}/{2}\right)\Gamma(\mu+2\kappa+d+1)\Gamma(2\kappa)}$ is a normalization constant.
+    For a given $\kappa>-\frac{1}{2}$, the validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ follow as a special case of Theorem 2 by setting $l=\frac{1}{2}$:
     • (A)
-      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ if and only if $\mu\geq\kappa+\frac{d+1}{2}$, whenever $\kappa\geq 0$ or $d\geq 2$;
+      $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ if and only if $\mu\geq\kappa+\frac{d+1}{2}$, whenever $\kappa\geq 0$ or $d\geq 2$;
     • (B)
-      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{1}$ if $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$, and only if $\mu\geq\kappa+1$, for $\frac{1}{2}<\kappa<0$.
+      $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{1}$ if $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$, and only if $\mu\geq\kappa+1$, for $\frac{1}{2}<\kappa<0$.
     Since $\kappa+1<\frac{\sqrt{8\kappa+9}-1}{2}$ for all $\frac{1}{2}<\kappa<0$, a gap exists between the necessary and sufficient conditions.
@@
 Whether this lower bound is also necessary remains an open problem.
-    The ${\cal GW}_{\kappa,\mu,\beta}$ model, as in the ${\cal MT}_{\nu,\alpha}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$.
-Specifically, for any integer $l=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,\beta}$ model are $l$ times differentiable, in any direction, if and only if $\kappa>l-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
+    The ${\cal GW}_{\kappa,\mu,a}$ model, as in the ${\cal MT}_{\nu,a}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$.
+Specifically, for any integer $m=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,a}$ model are $m$ times differentiable, in any direction, if and only if $\kappa>m-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
     When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$.
     Recently, Bevilacqua et al. (2022) considered a compact support reparameterized version of the $\cal GW$} model, and proved that
-    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(\beta,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (6)
-    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $0,\infty)$.
+    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(a,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,a}(x),\quad\kappa>-0.5,$ (6)
+    where $g(a,\mu,\kappa)=a\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $0,\infty)$.
     As a consequence, the reparameterized $\cal GW$} model can be viewed as a generalization of the $\cal MT$} model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
@@
 Using some results in Cho and Yun (2018), Emery and Alegría (2022) show that, for a given $a>0$, the following conditions:
-  1. (a)
-    $\delta>\frac{d}{2}$;
-  2. (b)
-    $2(\beta-\delta)(\gamma-\delta)\geq\delta$;
-  3. (c)
-    $\beta+\gamma\geq 3\delta+\frac{1}{2}$;
+$(a)\ \delta>\frac{d}{2},\qquad(b)\ 2(\beta-\delta)(\gamma-\delta)\geq\delta,\qquad(c)\ \beta+\gamma\geq 3\delta+\frac{1}{2}\,$,
 are sufficient for the positivity of the spectral density on $0,\infty)$. The associated four-parameter isotropic correlation model is given by:
@@
 As in the $\cal MT$} and $\cal GW$} cases, the $\cal G$H} model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\delta$.
-Furthermore, it is not difficult to see that the $\cal G$H} model includes the $\cal GW$} and the $\cal MT$} models as special cases. For instance,
+Furthermore, it is not difficult to see that the $\cal G$H} model includes the $\cal GW$} and the $\cal MT$} models as special cases (Emery and Alegría, 2022). For instance:
 $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+\frac{1}{2},a}(x)={\cal GW}_{\kappa,\mu,a}(x),$ (9)
@@
 Although the $\cal G$H} family is highly flexible, it is not identifiable in the usual parametric sense.
-Indeed, even under increasing-domain asymptotics—where one would in principle expect all parameters to be estimable—the parameter-to-correlation map is not injective.
+Indeed, even under increasing-domain asymptotics—where one would in principle expect all parameters to be estimable (Mardia and Marshall, 1984)—the parameter-to-correlation map is not injective.
 In particular,
-$\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a},$
+$\mathcal{GH}_{\delta,\beta,\gamma,a}(x)=\mathcal{GH}_{\delta,\gamma,\beta,a}(x),$
 because the Gauss hypergeometric function is symmetric in its first two arguments:
@@
 Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures and let $\delta>\frac{d}{2}$.
-If $\frac{d+1}{2}+3\delta<\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0}),$ for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, if and only if
+If $\frac{d+1}{2}+3\delta<\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0}),$ and $2(\beta_{i}-\delta)(\gamma_{i}-\delta)\geq\delta.$ for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, if and only if
 $\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}},$ (10)
@@
 This issue can be observed under both increasing and fixed domain scenario.
-To this end, in the next Section we assume that the second shape parameter ($\gamma$) is a function of the first ($\beta$).
 4 Fixing the identifiability problem: the Hypergeometric correlation model
-In this section we first refine the validity conditions of the $\mathcal{GH}$ model, clarifying the relationship between necessary and sufficient constraints.
-We then address the identifiability issue discussed in the previous section by introducing a parsimonious subclass selected according to an explicit optimality criterion.
+In this section, we first refine the validity conditions of the $\mathcal{GH}$ model, clarifying the relationship between necessary and sufficient constraints.
+We then address the non-uniqueness discussed in the previous section by imposing an appropriate restriction on the parameter space, which restores a unique representation of the correlation function.
+Finally, we introduce a parsimonious compactly supported subclass selected through an explicit optimality criterion.
 To this end, we adopt a $\cal GW$}-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model and set
-$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\beta+l.$ (11)
+$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\beta+l,\qquad l\geq 0.$ (11)
 As noted in (9), this parametrization includes the $\cal GW$} model as the special case $l=\frac{1}{2}$.
-The following theorem improves Theorem 1 in Emery and Alegría (2022) by providing sharp validity conditions for $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
+The restriction $l\geq 0$ in (11) is equivalent to imposing the ordering constraint $\gamma\geq\beta$ on the original $\mathcal{GH}_{\delta,\beta,\gamma,a}$ family.
+This removes the intrinsic label-switching ambiguity induced by the symmetry of $_{2}F_{1}$ in its upper parameters, namely $\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a}$, and thus yields a unique representation up to this exchange.
+The following theorem improves Theorem 1 in Emery and Alegría (2022) by providing sharp validity conditions for the reparameterized model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
 Theorem 2.
-Let $d$ be fixed, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $a>0$, and $l\geq 0$. Then:
+Let $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $a>0$. Then:
   1. 1.
@@
     $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if}\quad\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l,$
-    and only if
-    $\mu\geq\delta-l+\frac{1}{2}.$
-For $l=\frac{1}{2}$, the necessary and sufficient validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ discussed in Section 2 are recovered as a special case.
-Specifically, since $l=\frac{1}{2}\leq\frac{d}{2}+\kappa$ if and only if $\kappa\geq\frac{1}{2}-\frac{d}{2}$, which holds for all $d\geq 2$ and $\kappa>-\frac{1}{2}$, or for $d=1$ and $\kappa\geq 0$, Regime 1 of Theorem 2 gives the necessary and sufficient condition
-$\mu\geq\frac{d+1}{2}+\kappa,$
-For $d=1$ and $\frac{1}{2}<\kappa<0$, instead $l=\frac{1}{2}>\frac{d}{2}+\kappa$, so Regime 2 applies, yielding the sufficient condition $\mu\geq 0.5(\sqrt{8\kappa+9}-1)$ and the necessary condition $\mu\geq\kappa+1$, consistently with the discussion in Section 2.
-As discussed in Section 3, the model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under the interchange $\beta\leftrightarrow\gamma$, corresponding to $l\mapsto-l$ in the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
-The admissible region is therefore symmetric with respect to $l=0$, and restricting attention to $l\geq 0$ entails no loss of generality.
-Translating Theorem 2 back to the original parameterization $\mathcal{GH}_{\delta,\beta,\gamma,a}$, the following equivalent conditions hold for $\delta>d/2$ and $a>0$:
+    and only if $\mu\geq\delta-l+\frac{1}{2}.$
+Remark 1.
+Tracing back the conditions of Theorem 2 to the original parametrization $(\delta,\beta,\gamma,a)$, and assuming $\delta>d/2$ and $\gamma\geq\beta$ we obtain:
   1. 1.
@@
     $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if}\quad\beta+\gamma\geq 2\delta+\sqrt{(\gamma-\beta)^{2}+2\delta},$
-    and only if
-    $\beta+\gamma\geq 3\delta+\frac{1}{2}.$
-In particular, in Regime 1 ($0\leq\gamma-\beta\leq\delta-\frac{1}{2}$), condition (c) of Section 3 implies condition (b), so that (a) and (c) alone characterize validity.
-Hence, within this regime, the sufficient conditions of Emery and Alegría (2022) reduce to (a) and (c) and are therefore also necessary.
-In Regime 2 ($\gamma-\beta>\delta-\frac{1}{2}$), whether the gap between sufficiency and necessity can be closed remains an open question.
-We now turn to the structural identifiability issue for the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model.
-To resolve it, the parameter $l$ must either be fixed (as in the $\cal GW$} model) or linked to the remaining parameters and/or to the spatial dimension.
+    and only if $\beta+\gamma\geq 3\delta+\frac{1}{2}.$
+In particular, on the broad regime $\gamma-\beta\leq\delta-\tfrac{1}{2}$ the sufficient conditions (a)–(c) in Section 3 reduce to the single inequality $\beta+\gamma\geq 3\delta+\tfrac{1}{2}$ and become sharp, since this inequality is also necessary.
+This sharpens the earlier validity region reported in Emery and Alegría (2022).
+Outside this regime, the sufficient bound above is explicit while a residual gap with the necessary condition remains.
+For $l=\frac{1}{2}$, the necessary and sufficient validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ discussed in Section 2 are recovered.
+Indeed, if $l=\frac{1}{2}\leq\frac{d}{2}+\kappa$ (which holds for all $d\geq 2$ and $\kappa>-\frac{1}{2}$, and for $d=1$ with $\kappa\geq 0$), then Regime 1 of Theorem 2 yields $\mu\geq\delta=\frac{d+1}{2}+\kappa$.
+For $d=1$ and $\frac{1}{2}<\kappa<0$, instead $l=\frac{1}{2}>\frac{d}{2}+\kappa$, so Regime 2 applies, yielding the sufficient condition $\mu\geq\frac{1}{2}(\sqrt{8\kappa+9}-1)$ and the necessary condition $\mu\geq\kappa+1$, consistently with the discussion in Section 2.
+Theorem 2 characterizes validity for each admissible value of $l$.
+However, even after restricting to $l\geq 0$ the reparameterized model still involves two shape parameters, $\mu$ and $l$.
+In the spirit of the $\cal GW$} model, we therefore seek a parsimonious specification by treating $l$ as fixed (or by linking it to the remaining parameters and/or to the spatial dimension), so that inference targets a single interpretable shape parameter $\mu$.
 This naturally raises the question: which choice of $l\geq 0$ is optimal?
-We address this question by introducing an explicit optimality criterion that selects a single parsimonious representative once smoothness and compact support are fixed.
-We adopt the criterion of maximum integral range attainable by a compactly supported correlation with given smoothness and support.
+We address this question by introducing an explicit optimality criterion that selects a parsimonious subclass once smoothness and compact support are fixed.
+Specifically, we choose $l$ so that the resulting $\mathcal{GH}$ subclass can attain the largest isotropic integral range.
 For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is
-$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\,\mathrm{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{0}^{\infty}x^{d-1}\phi(x)\,\mathrm{d}x.$ (12)
+$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\,\mathrm{d}\boldsymbol{h}=\frac{2\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}\int_{0}^{\infty}x^{d-1}\phi(x)\,\mathrm{d}x.$ (12)
 This quantity, originally introduced by Lantuéjoul (1991), provides a global and threshold-free measure of dependence (Yaglom, 1987; Matheron, 1989; Hristopulos and Zukovic, 2011).
-If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$, so that the integral range coincides with the spectral density at zero frequency.
-When $d=1$, AA reduces to the area under the correlation function.
-For compactly supported correlations, maximizing $A$ therefore selects, among all admissible models with fixed smoothness and support, the one with maximal low-frequency spectral power, that is, the strongest large-scale dependence compatible with the imposed constraints.
+If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then
+$A=(2\pi)^{d}\,\widehat{\phi}(0),$
+so the integral range is proportional to the spectral density at zero frequency.
+For compactly supported correlations, maximizing $A$ therefore targets models with maximal low-frequency spectral power, that is, the strongest large-scale dependence compatible with the imposed constraints.
 While alternative optimality principles could be considered, they would typically depend on the sampling design or on a specific loss function.
-In contrast, the integral range criterion is intrinsic to the model class and yields a reproducible, one-to-one specification within the family.
+In contrast, the integral range criterion is intrinsic to the model class and yields a reproducible, one-to-one specification (through $l$) within the family.
 For a parametric model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we denote the associated integral range by $A_{\boldsymbol{\theta},d}$. For example, for the Matérn correlation one has
-$A_{\nu,\alpha,d}=\frac{\alpha^{d}\pi^{d/2}\Gamma\!\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (13)
+$A_{\nu,a,d}=2^{d}\pi^{d/2}a^{d}\,\frac{\Gamma\!\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (13)
 The following theorem provides the integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
 Theorem 3.
-The integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is given by:
-$A_{\kappa,\mu,l,a,d}=a^{d}\left[\frac{\pi^{d/2}\Gamma(\frac{d+1}{2}+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+\frac{1}{2})\Gamma\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\left(\frac{\mu+1+d}{2}+\kappa+l\right)}\right].$ (14)
-Note that, as expected, the compact support $a$ raised to power $d$ represents a scale factor in the integral range computation.
+The integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is given by
+$A_{\kappa,\mu,l,a,d}=a^{d}\left[\frac{\pi^{d/2}\Gamma(\frac{d+1}{2}+\kappa)\,\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{\Gamma(\kappa+\frac{1}{2})\,\Gamma\left(\frac{\mu+1+d}{2}+\kappa\right)\,\Gamma\left(\frac{\mu+1+d}{2}+\kappa+l\right)}\right].$ (14)
+Note that the compact support $a$ raised to power $d$ represents a scale factor in the integral range computation.
 In addition, using asymptotic properties of the Gamma function, it can be shown that $A_{\kappa,\mu,l,a,d}\sim\mu^{-d}$ as $\mu\to\infty$ and, as a consequence,
@@
 This implies that, as the shape parameter $\mu$ increases, the integral range approaches zero, regardless of the compact support and smoothness parameters.
-The following theorem establishes that, for a given $l\geq 0$, the maximum integral range is obtained by fixing $\mu$ on the lower bound of its parametric space.
+The following theorem establishes that, for a given $l\geq 0$, the integral range is maximized by taking $\mu$ at the lower bound of its admissible domain.
+In what follows, we call admissible the parameter region ensured by Theorem 2.
+In Regime 2 of Theorem 2, validity is guaranteed only through the explicit sufficient bound.
+Throughout Sections 45, whenever we optimize over $(\mu,l)$ we therefore interpret the admissible region as the parameter set ensured by the sufficient conditions in Theorem 2.
 Theorem 4.
-Let $\kappa>0$, $a>0$, and $d\in\mathbb{N}$ be fixed.
-For each $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ with respect to $\mu$ (over its admissible parameter space) is attained at $\mu^{\star}(l)$, where:
+Let $\kappa>-\frac{1}{2}$, $a>0$, and $d\in\mathbb{N}$ be fixed, and set $\delta=\frac{d+1}{2}+\kappa$.
+For each $l\geq 0$, the map $\mu\mapsto A_{\kappa,\mu,l,a,d}$ is non-increasing on its admissible domain.
+Hence its maximum is attained at the smallest admissible value of $\mu$.
+In particular,
   1. 1.
-    if $0\leq l\leq\dfrac{d}{2}+\kappa$, then
-    $\mu^{\star}(l)=\mu_{1}(l):=\delta-l+\frac{1}{2};$
+    if $0\leq l\leq d/2+\kappa$, then the maximum is attained at μ^(l)=μ_1(l):=δ-l+12;
   2. 2.
-    if $l>\dfrac{d}{2}+\kappa$, then
-    $\mu^{\star}(l)=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l.$
-For instance, let $\kappa\geq 0$.
-If $l=\frac{1}{2}$, then from (9) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of $\cal GW$} model in this case.
-As an example, Figure 1 (left part) depicts the ${\cal GW}_{0,\mu,1}$ model and the right part depicts the ${\cal GW}_{1,\mu,1}$ model for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$.
-In these examples $d=1$ and it is apparent that the model with maximum area corresponds to the case $\mu=\mu_{1}(\frac{1}{2})$ (red line) in both examples.
-The next step is to find $l\geq 0$ that maximizes $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$.
+    if $l>d/2+\kappa$, then the maximum is attained at μ^(l)=μ_2(l):=2κ+l^2+d+1-l.
+For instance, if $l=\frac{1}{2}$ and $d\geq 2$ (or $d=1$ and $\kappa\geq 0$), then by (9)
+$\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a},$
+and the maximum integral range (over the admissible $\mu$) is attained at $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which coincides with the lower bound of the $\cal GW$} model in this case.
+As an example, consider the case $d=1$ and $\kappa=0,1$.
+Figure 1 (left) depicts ${\cal GW}_{0,\mu,1}$ and Figure 1 (right) depicts ${\cal GW}_{1,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$.
+It is clear that, in both panels, the correlation with the largest area under the curve is obtained for $\mu=\mu_{1}\!\left(\tfrac{1}{2}\right)$ (red line).
+The next step is to determine which $l\geq 0$ yields the largest attainable integral range, i.e., which $l$ maximizes $A_{\kappa,\mu^{\star}(l),l,a,d}$.
 Theorem 5.
-Let $\kappa>-1/2$, $a>0$, and $d\in\mathbb{N}$ be fixed, and set $\delta=\frac{d}{2}+\kappa$.
+Let $\kappa>-1/2$, $a>0$, and $d\in\mathbb{N}$ be fixed.
   1. 1.
-    The function $l\mapsto A_{\kappa,\mu_{1}(l),l,a,d}$ is non-decreasing on $0,\delta]$, hence
-    $\max_{0\leq l\leq\delta}A_{\kappa,\mu_{1}(l),l,a,d}=A_{\kappa,\mu_{1}(\delta),\delta,a,d}=A_{\kappa,1,\delta,a,d}.$
+    The function $l\mapsto A_{\kappa,\mu_{1}(l),l,a,d}$ is non-decreasing on $0,d/2+\kappa]$, hence
+    $\max_{0\leq l\leq d/2+\kappa}A_{\kappa,\mu_{1}(l),l,a,d}=A_{\kappa,1,d/2+\kappa,a,d}.$
   2. 2.
-    The function $l\mapsto A_{\kappa,\mu_{2}(l),l,a,d}$ is non-increasing on $(\delta,\infty)$ and
-    $\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}=A_{\kappa,1,\delta,a,d}.$
-In particular, the global maximum over all admissible $(\mu,l)$ is attained at $(\mu,l)=(1,\delta)$ and is equal to $A_{\kappa,1,\delta,a,d}$.
+    The function $l\mapsto A_{\kappa,\mu_{2}(l),l,a,d}$ is non-increasing on $(d/2+\kappa,\infty)$ and
+    $\lim_{l\downarrow d/2+\kappa}A_{\kappa,\mu_{2}(l),l,a,d}=A_{\kappa,1,d/2+\kappa,a,d}.$
+In particular, the largest attainable integral range over all admissible $(\mu,l)$ is achieved at $(\mu,l)=(1,d/2+\kappa)$ and equals $A_{\kappa,1,d/2+\kappa,a,d}$.
 Refer to caption Refer to caption
-Figure 1: Left part: ${\cal GW}_{0,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$ with $\delta=\frac{d+1}{2}+\kappa$ and $\kappa=0$.
+Figure 1: Left part: ${\cal GW}_{0,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$ with $\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$ and $\kappa=0$.
 Right part: the same Figure but with $\kappa=1$.
 In both cases $d=1$.
-As an example, Figure 2 compares $\mathcal{GH}_{\delta,\delta+\frac{\mu_{1}(l)}{2},\delta+\frac{\mu_{1}(l)}{2}+l,1}$ when $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines, respectively).
+As an example, Figure 2 compares $\mathcal{GH}_{\delta,\delta+\frac{\mu_{1}(l)}{2},\delta+\frac{\mu_{1}(l)}{2}+l,1}$ when $l=0,\frac{1}{2},d/2+\kappa$ (black, red and blue lines, respectively).
 In particular the left part depicts the case $\kappa=0$ and the right part depicts the case $\kappa=1$ ($d=2$ in both cases).
-It is apparent that the case $l=\frac{d}{2}+\kappa$ (blue line) corresponds to the model with higher integral range and improves the maximum integral range attained by the $\cal GW$} models (red line).
+It is apparent that the case $l=d/2+\kappa$ (blue line) corresponds to the model with higher integral range and improves the maximum integral range attained by the $\cal GW$} models (red line).
 Refer to caption Refer to caption
-Figure 2: Left part: $\mathcal{GH}_{\delta,\delta+\frac{\mu(l)}{2},\delta+\frac{\mu(l)}{2}+l,1}$ with $\delta=\frac{d+1}{2}+\kappa$, $\kappa=0$, $\mu(l)=\delta-l+\frac{1}{2}$ and $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines respectively).
+Figure 2: Left part: $\mathcal{GH}_{\delta,\delta+\frac{\mu(l)}{2},\delta+\frac{\mu(l)}{2}+l,1}$ with $\delta=\frac{d+1}{2}+\kappa$, $\kappa=0$, $\mu(l)=\delta-l+\frac{1}{2}$ and $l=0,\frac{1}{2},d/2+\kappa$ (black, red and blue lines respectively).
 Right part: the same Figure but with $\kappa=1$.
@@
 In both cases $d=2$.
-Summarizing our results, the model that can potentially achieve the maximum integral range is $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ and, in particular, the maximum is achieved when $\mu=\mu_{1}(d/2+\kappa)=1$.
-Taking into account (15), this implies that
-$0\leq A_{\kappa,\mu,l,a,d}\leq a^{d}\left[\frac{\pi^{d/2}\,\Gamma\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\left(1+\kappa\right)\,\Gamma\left(1+\frac{d}{2}+2\kappa\right)}{2^{d}\,\Gamma\left(\kappa+\frac{1}{2}\right)\,\Gamma\left(1+\frac{d}{2}+\kappa\right)\,\Gamma\left(1+d+2\kappa\right)}\right].$
-Accounting for (11), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron (1965, I.4.18).
-Fixing $l=\frac{d}{2}+\kappa$, we define the hypergeometric correlation by
+Summarizing, Theorem 5 shows that the largest attainable integral range over all admissible $(\mu,l)$ is achieved at $(\mu,l)=(1,d/2+\kappa).$ In particular, fixing $l=d/2+\kappa$ identifies the unique $\mathcal{GH}$ specification that can attain this maximum (with the maximum attained at $\mu=1$).
+Accounting for (11), the extremal member attaining the maximal integral range (i.e., $l=d/2+\kappa$ and $\mu=1$) is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$.
+This coincides with the Euclid’s hat correlation model (Gneiting, 1999) when $\kappa=0$ (triangular/tent for $d=1$, circular for $d=2$, spherical for $d=3$, pentaspherical for $d=5$, see Section 5), and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría, 2022), obtained via a (fractional) montée (Matheron, 1965, I.4.18).
+Motivated by this, we fix $l=d/2+\kappa$ and define the hypergeometric correlation by
 $\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\qquad x\geq 0.$ (16)
@@
 $\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)}{\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\frac{1}{2}\right)}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\[4.0pt] &\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\qquad x\geq 0.\end{split}$ (17)
-Notably, the model $\mathcal{H}_{\kappa,\mu,a,d}$ always lies in Regime 1 of Theorem 2, where validity is completely characterized.
-Indeed, by setting $l=\frac{d}{2}+\kappa$ in Theorem 2, we obtain the condition
+Notably, the model $\mathcal{H}_{\kappa,\mu,a,d}$ always lies in Regime 1 of Theorem 2, where validity is completely characterized. Indeed, by setting $l=d/2+\kappa$ in Theorem 2, we obtain
 $\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq 1.$
+Refer to caption Refer to caption
+Figure 3: Two Gaussian RF realizations with ${\cal H}_{\kappa,4,0.2,2}$ correlation model, for $\kappa=0,1$ (from left to right).
 These validity conditions are substantially simpler than those for the $\mathcal{GW}_{\kappa,\mu,a}$ model discussed in Section 2.
@@
 The trade-off is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ explicitly depends on $d$.
-The next theorem shows that the subclass ${\cal H}_{\kappa,1,a,d}$, obtained by maximizing the integral range, retains substantial modeling flexibility.
+The next theorem shows that the extremal subclass ${\cal H}_{\kappa,1,a,d}$, which attains the maximal integral range, retains substantial modeling flexibility.
 Through beta-type scale mixtures, it can generate any member of ${\cal H}_{\kappa,\mu,a,d}$ and a broad subclass of the Gauss–hypergeometric family $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
@@
 Theorem 6.
-Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$, and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ by Theorem 2.
-Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ admits a beta-type scale mixture representation in terms of hypergeometric correlations $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
+Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$ and $l\geq 0$.
+Assume that the parameters $(\mu,l)$ belong to the admissible region of Theorem 2, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$.
+In addition, assume $\mu+2l>1+d+2\kappa$.
+Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ admits a beta-type scale mixture representation in terms of $\mathcal{H}_{\kappa,1,b,d}$ with $b=a\sqrt{uv}$:
 $\begin{split}\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu-d}{2}+l\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu-1}{2}\right)\Gamma\!\left(\frac{\mu-1-d}{2}-\kappa+l\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$ (18)
-In particular, when $l=\frac{d}{2}+\kappa$, one obtains
+In particular, when $l=d/2+\kappa$, one obtains
 $\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu}{2}+\kappa\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma^{2}\!\left(\frac{\mu-1}{2}\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$
-Clearly, the ${\cal H}_{\kappa,\mu,a,d}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$ in the same way as the ${\cal GW}_{\kappa,\mu,\beta}$ model.
-Figure 3 depicts two realizations of zero-mean and unit-variance Gaussian RFs with correlations ${\cal H}_{\kappa,4,0.2,2}$ for $\kappa=0,1,2$.
-The realizations were constructed via the Cholesky decomposition of the covariance matrix.
-Refer to caption Refer to caption
-Figure 3: Two Gaussian RF realizations with ${\cal H}_{\kappa,4,0.2,2}$ correlation model, for $\kappa=0,1$ (from left to right).
+[…]
 5 Some special cases and bridges to the Matérn model
@@
 Theorem 7.
-If $\kappa=k$ is a nonnegative integer, $d=1$, and $\mu\geq 1$, then
-$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\!\left(k+\tfrac{1}{2}\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\quad\times\sum_{n=0}^{k}\frac{(n+k)!}{2^{2n}n!\,\Gamma\!\left(n+\frac{\mu}{2}+k+1\right)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
+If $\kappa=k$ is a nonnegative integer, $d=1$, and $\mu\geq 1$, then, for $x\geq 0$,
+$\mathcal{H}_{k,\mu,a,1}(x)=\begin{cases}\displaystyle\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\!\left(k+\tfrac{1}{2}\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\[6.00006pt] \displaystyle\qquad\times\sum_{n=0}^{k}\frac{(n+k)!}{2^{2n}n!\,\Gamma\!\left(n+\frac{\mu}{2}+k+1\right)(k-n)!}\left(\frac{\left(1-\frac{x}{a}\right)^{2}}{\frac{x}{a}}\right)^{n},&0\leq x<a,\\[11.00008pt] 0,&x\geq a.\end{cases}$
 General closed forms for arbitrary $d>1$, $\mu\geq 1$, and integer $\kappa$ are more difficult to obtain.
@@
 $\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum_{n=0}^{\frac{d-1}{2}}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d+1}{2}\right)_{n}\left(1-\frac{x}{a}\right)_{+}^{n}}{2^{n}n!\,\Gamma\!\left(n+\frac{d-1}{2}+\mu+1\right)},$
-where
-$C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\!\left(\frac{\mu+1}{2}\right)\Gamma\!\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}.$
-Additional closed-form expressions, including spherical models, are reported in Appendix B.
+where $C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\!\left(\frac{\mu+1}{2}\right)\Gamma\!\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}.$
+Additional closed-form expressions are reported in Appendix B, including the spherical models (Chilès and Delfiner, 2012) that are widely used in geostatistical practice and that are recovered as special cases of the proposed hypergeometric family.
 The Matérn model can be recovered as a limiting case of $\mathcal{H}_{\kappa,\mu,a,d}$ under an appropriate compact-support reparameterization.
@@
 Theorem 9.
-Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures, and let $\kappa>-\frac{1}{2}$ and $\nu=\kappa+\frac{1}{2}$.
-For any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ are equivalent on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$, if and only if
-$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}\,B(\kappa,\mu,d),$ (19)
+Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures, and let $\kappa>-\frac{1}{2}$ , $\nu=\kappa+\frac{1}{2}$ and $\mu>1+d/2$ .
+For any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ are equivalent on the paths of $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, if and only if
+$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}\,F(\kappa,\mu,d),$ (19)
 where
-$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+d+1}{2}+2\kappa\right)}{\Gamma\!\left(\frac{\mu}{2}\right)\Gamma\!\left(\frac{\mu+d}{2}+\kappa\right)}.$ (20)
+$F(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+d+1}{2}+2\kappa\right)}{\Gamma\!\left(\frac{\mu}{2}\right)\Gamma\!\left(\frac{\mu+d}{2}+\kappa\right)}.$ (20)
 Assuming a common variance and imposing the relation $\nu=\kappa+\frac{1}{2}$, the equivalence condition (19) is equivalently expressed as
-$a=\alpha\,B(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
-Using this compact-support reparameterization and following the arguments in Bevilacqua et al. (2022), one obtains
+$a=\alpha\,F(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
+Using this compact-support reparameterization and following the same arguments in Bevilacqua et al. (2022), one obtains
 $\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa>-\frac{1}{2},$ (21)
@@
 It is worth emphasizing that, under this compact-support reparameterization, the role of the shape parameter $\mu$ changes substantially.
-For fixed smoothness $\kappa$ and scale parameter $\alpha=1$, $\mu$ determines the compact support.
-As $\mu\to\infty$, ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ converges to the Matérn model ${\cal MT}_{\kappa+\frac{1}{2},1}$, with maximal integral range given by (13).
-Conversely, as $\mu$ approaches its lower bound, the compact support decreases, potentially yielding more flexible local dependence structures.
-A comparison with the corresponding $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model (see (6)) shows that, over a substantial portion of the parameter space, the reparameterized $\mathcal{H}$ models can achieve smaller compact supports.
-In particular, when $d\geq 2$ or $\kappa\geq 0$,
-$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa),$
-where $\mu_{\min}(\kappa,d)=\kappa+\frac{d+1}{2}$.
-The case $d=1$ and $\frac{1}{2}<\kappa<0$ is excluded here, as the exact lower bound $\mu_{\min}$ for the $\mathcal{GW}$ model remains an open problem in that regime.
+For fixed smoothness $\kappa$ and scale parameter $\alpha$, $\mu$ determines the compact support.
+As $\mu\to\infty$, ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ converges to the Matérn model ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$, with maximal integral range given by (13).
 6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
-We study the asymptotic properties of the maximum likelihood (ML) estimator for the parameters of the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ under both increasing-domain and fixed-domain asymptotics.
-Under increasing-domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ are consistently estimable, and the ML estimator is asymptotically Gaussian with covariance matrix equal to the inverse Fisher information (Mardia and Marshall, 1984).
-Under fixed-domain asymptotics, the situation differs.
-Using Theorem 1 and the parameterization (11) with $l=\frac{d}{2}+\kappa$, the equivalence condition for two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing the same smoothness parameter, for $d=1,2,3$ and $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, is
-$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}}.$ (22)
-Hence only the microergodic parameter
-$\frac{\sigma^{2}B(\kappa,\mu,d)}{a^{2\kappa+1}}$
-is consistently estimable. If $\mu$ is fixed and known, this reduces to $\sigma^{2}/a^{2\kappa+1}$.
-Let $D\subset\mathbb{R}^{d}$ be bounded and $\mathbf{Z}_{n}=(Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n}))^{\prime}$ a realization of a zero-mean stationary Gaussian RF with covariance $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, $\phi\in\Phi_{d}$, $\phi(0;\boldsymbol{\tau})=1$.
-Denote by
-$R_{n}(\boldsymbol{\tau})=[\phi(\|\mathbf{s}_{i}-\mathbf{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$
-the correlation matrix. The Gaussian log-likelihood is
-$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\Big(n\log(2\pi\sigma^{2})+\log|R_{n}(\boldsymbol{\tau})|+\frac{1}{\sigma^{2}}\mathbf{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\mathbf{Z}_{n}\Big).$ (23)
-Under the $\mathcal{H}$ model, $\phi(\cdot;\boldsymbol{\tau})=\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$.
-In what follows, $\kappa$ and $\mu$ are assumed fixed and known, so we write $\tau=a$.
-Let $\hat{\sigma}_{n}^{2}(a)$ denote the ML estimator of $\sigma^{2}$ for fixed $a>0$:
-$\hat{\sigma}_{n}^{2}(a)=\mathbf{Z}_{n}^{\prime}R_{n}(a)^{-1}\mathbf{Z}_{n}/n.$ (24)
-The following result establishes strong consistency and asymptotic normality of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$. The proof follows the same arguments as Theorem 8 in Bevilacqua et al. (2019).
+We study the asymptotic properties of likelihood-based estimators under fixed-domain (infill) asymptotics for the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, with $\kappa$ fixed and known.
+Using Theorem 1 and the parameterization (11) with $l=d/2+\kappa$, the equivalence condition for two Gaussian measures $P(\sigma_{i}^{2}\mathcal{H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing the same $\kappa$, for $d=1,2,3$ and $\min(\mu_{0},\mu_{1})>1+d/2$, is
+$\frac{\sigma_{0}^{2}\,F(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}\,F(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}}.$ (22)
+Hence, under fixed-domain (infill) asymptotics, the individual parameters $\sigma^{2}$, $\mu$ and $a$ are not consistently identifiable; only the microergodic parameter
+$\theta(\sigma^{2},\mu,a):=\frac{\sigma^{2}\,F(\kappa,\mu,d)}{a^{2\kappa+1}}$ (23)
+is consistently estimable.
+In our fixed-domain limit theory we assume $\mu>1+d/2$ for $d=1,2,3$, which is the condition ensuring equivalence of Gaussian measures (see Theorem 1).
+This requirement is stronger than the mere validity condition for the $\mathcal{H}$ model, which is simply $\mu\geq 1$.
+Let $D\subset\mathbb{R}^{d}$ be a bounded domain and let $S_{n}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\subset D$ be an infill design, i.e.
+$S_{n}$ becomes dense in $D$ as $n\to\infty$.
+Let $\mathbf{Z}_{n}=(Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n}))^{\top}$ be a realization of a zero-mean stationary Gaussian RF with covariance function $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}(\|\cdot\|)$.
+Write
+$R_{n}(\mu,a):=\bigl[\mathcal{H}_{\kappa,\mu,a,d}(\|\mathbf{s}_{i}-\mathbf{s}_{j}\|)\bigr]_{i,j=1}^{n}$
+for the $n\times n$ correlation matrix, and
+$\ell_{n}(\sigma^{2},\mu,a):=-\frac{1}{2}\Bigl(n\log(2\pi\sigma^{2})+\log\lvert R_{n}(\mu,a)\rvert+\frac{1}{\sigma^{2}}\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}\Bigr)$ (24)
+for the Gaussian log-likelihood. Throughout this section $\kappa$ is fixed and known, while $\mu$ and $a$ may be either fixed or estimated as specified below.
+For any $(\mu,a)$, the profile ML estimator of $\sigma^{2}$ is
+$\hat{\sigma}_{n}^{2}(\mu,a):=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$ (25)
+and the corresponding plug-in estimator of the microergodic parameter (23) is
+$\hat{\theta}_{n}(\mu,a):=\theta\!\bigl(\hat{\sigma}_{n}^{2}(\mu,a),\mu,a\bigr)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}}.$ (26)
+The profile log-likelihood obtained by substituting (25) into (24) is
+$\mathcal{PL}_{n}(\mu,a):=-\frac{1}{2}\Bigl(n\log\bigl(2\pi\hat{\sigma}_{n}^{2}(\mu,a)\bigr)+\log\lvert R_{n}(\mu,a)\rvert+n\Bigr).$ (27)
+When $\mu$ is fixed and known, the factor $F(\kappa,\mu,d)$ in (23) is a known constant.
+Thus, the equivalence condition (22) reduces to $\sigma_{0}^{2}/a_{0}^{2\kappa+1}=\sigma_{1}^{2}/a_{1}^{2\kappa+1}$, and the microergodic parameter can equivalently be taken as $\sigma^{2}/a^{2\kappa+1}$ (up to the known multiplicative constant $F(\kappa,\mu,d)$).
+We keep the notation (23) for uniformity, noting that all statements can be rescaled by $F(\kappa,\mu,d)$ when $\mu$ is fixed.
+The following result establishes strong consistency and asymptotic normality of $\hat{\theta}_{n}(\mu,a)$ when both $\mu$ and $a$ are treated as fixed.
+The proof follows the same arguments as Theorem 8 in Bevilacqua et al.
+(2019) and is omitted.
 Theorem 10.
-Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 1+\frac{d}{2}$.
-Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)^{2}$.
-Then, for fixed $a>0$, as $n\to\infty$,
-  1. 1.
-    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
-  2. 2.
-    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
-We next consider joint maximization of (23) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$, $I=[a_{L},a_{U}]$, $0<a_{L}<a_{U}<\infty$. Let
-$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\mathbf{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\mathbf{Z}_{n}/n,\quad\hat{a}_{n}=\arg\max_{a\in I}\mathcal{PL}_{n}(a),$
-where $\mathcal{PL}_{n}(a)$ is the profile log-likelihood
-$\mathcal{PL}_{n}(a)=-\frac{1}{2}\big(\log(2\pi)+n\log\hat{\sigma}_{n}^{2}(a)+\log|R_{n}(a)|+n\big).$ (25)
-The following lemma establishes a monotonicity property of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ as a function of $a$. Its proof is given in Appendix A.
-Lemma 1.
-Let $S_{n}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\subset D$. For any $a_{1}<a_{2}$ in $I$,
-$\frac{\hat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}\leq\frac{\hat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}\quad\text{if and only if}\quad\mu\geq 4.$
-The next theorem establishes the asymptotic properties of $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$. The proof follows Theorem 9 in Bevilacqua et al. (2019) combined with Lemma 1.
+Let $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, $D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma_{0}^{2}\,\mathcal{H}_{\kappa,\mu,a_{0},d}$, where $\kappa$ and $\mu>1+d/2$ are fixed and known, and $(\sigma_{0}^{2},a_{0})\in(0,\infty)^{2}$.
+Then, for any fixed $a>0$, as $n\to\infty$:
+  1. (i)
+    $\hat{\theta}_{n}(\mu,a)\xrightarrow{a.s.}\theta_{0}:=\sigma_{0}^{2}\,F(\kappa,\mu,d)/a_{0}^{2\kappa+1}$;
+  2. (ii)
+    $\sqrt{n}\,\bigl(\hat{\theta}_{n}(\mu,a)-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}\bigl(0,\,2\theta_{0}^{2}\bigr)$.
+We now estimate $a$ by ML over the compact set $I=[a_{L},a_{U}]$:
+$\hat{a}_{n}:=\arg\max_{a\in I}\,\mathcal{PL}_{n}(\mu,a),\quad 0<a_{L}<a_{U}<\infty,$ (28)
+and study the asymptotic behaviour of the random variable $\hat{\theta}_{n}(\mu,\hat{a}_{n})$.
+The key technical ingredient is a uniform stochastic control of $\hat{\theta}_{n}(\mu,a)$ over $a\in I$.
+Lemma 2 (Appendix A) provides this control and implies that replacing a fixed $a$ by the random ML estimator $\hat{a}_{n}$ does not affect the fixed-domain limit distribution of the microergodic estimator.
 Theorem 11.
-Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 4$.
-Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$.
-Then, as $n\to\infty$,
-  1. 1.
-    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
-  2. 2.
-    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
-The condition $\mu\geq 4$ ensures the monotonicity property in Lemma 1 and is technical in nature.
-Finally, we compare the two reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$.
-Under fixed-domain asymptotics and using (22), the associated microergodic parameters are
-$\frac{\sigma^{2}}{\alpha^{2\kappa+1}}\quad\text{and}\quad\frac{\sigma^{2}B(\kappa,\mu,d)}{(\alpha\mu)^{2\kappa+1}},$
-respectively.
-In the first case, the microergodic parameter does not depend on $\mu$, so both $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$ are consistently estimable.
-From an inferential perspective, $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ is therefore preferable.
-For large datasets, however, $\mu$ may be fixed to induce sparsity, in which case the microergodic parameter reduces to $\sigma^{2}/\alpha^{2\kappa+1}$ under both parameterizations.
+Under the same assumptions as Theorem 10, and $a_{0}\in I$, let $\hat{a}_{n}$ be as in (28). Then, as $n\to\infty$:
+  1. (i)
+    $\hat{\theta}_{n}(\mu,\hat{a}_{n})\xrightarrow{p}\theta_{0}$;
+  2. (ii)
+    $\sqrt{n}\bigl(\hat{\theta}_{n}(\mu,\hat{a}_{n})-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2})$.
+We finally allow $\mu$ to be unknown and estimate $(\mu,a)$ jointly by ML over the compact set $J\times I$, where $J=[\mu_{L},\mu_{U}]$ and $\mu_{L}>1+d/2$:
+$(\hat{\mu}_{n},\hat{a}_{n}):=\arg\max_{(\mu,a)\in J\times I}\,\mathcal{PL}_{n}(\mu,a).$ (29)
+As in Case 2, the key step is to control $\hat{\theta}_{n}(\mu,a)$ uniformly over the parameter region $J\times I$.
+Lemma 3 (Appendix A) provides a two-parameter analogue of Lemma 2.
+Together with Theorem 10, this yields the following result.
+Theorem 12.
+Under the same assumptions as Theorem 10, with $(\sigma_{0}^{2},\mu_{0},a_{0})\in(0,\infty)\times J\times I$ and $\mu_{0}>1+d/2$, let $(\hat{\mu}_{n},\hat{a}_{n})$ be as in (29).
+Then, as $n\to\infty$:
+  1. (i)
+    $\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})\xrightarrow{p}\theta_{0}$;
+  2. (ii)
+    $\sqrt{n}\,\bigl(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}\bigl(0,\,2\theta_{0}^{2}\bigr)$.
+Theorems 1112 show that the fixed-domain limit theory for the microergodic parameter under the hypergeometric model is stable with respect to ML estimation of nuisance covariance parameters over compact sets.
+Our proofs rely on a different strategy than the classical sandwich arguments based on finite-sample monotonicity (Kaufman and Shaby, 2013; Bevilacqua et al., 2019; Ma and Bhadra, 2022): we establish uniform stochastic control of the plug-in microergodic estimator over the relevant compact parameter sets (see Appendix A).
+This approach yields more general results than those currently available for related compactly supported families, such as the generalized Wendland model, where the restriction $\mu\geq(d+1)/2+\kappa+3$ is required (Bevilacqua et al., 2019).
 7 A simulation study
-We investigate the finite-sample performance of ML estimation for the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$.
-We simulate 500 realizations of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ for $d=2$, $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$.
-The process is generated via Cholesky decomposition on a regular grid with spacing $0.025$ over $0,1]^{2}$ (1681 locations), followed by independent perturbations of each coordinate by a uniform random variable on [0.01,0.01][-0.01,0.01].
-Throughout the study, $\kappa$ and $\mu$ are assumed known.
-We focus on ML estimation of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing- and fixed-domain scenarios.
-Following Zhang and Zimmerman (2005); Kaufman and Shaby (2013), these regimes can be mimicked by fixing the sampling grid and varying the spatial dependence.
-Accordingly, we consider $a=0.1,0.25,0.5$, interpreting $a=0.1$ as increasing-domain and $a=0.5$ as fixed-domain behavior.
-Table 1 reports the bias and standard deviation (in parentheses) of the 500 ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$, for $i=1,\ldots,500$.
-Table 1: Bias and standard deviation (in parenthesis) of ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ when $\kappa=0,1,2$, $\mu=4,6,8$, $a=0.1,0.25,0.5$ and $\sigma^{2}=1$.
+We study the finite-sample behavior of ML estimation under the proposed model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$.
+We generate $500$ independent realizations of a zero-mean Gaussian random field on $0,1]^{2}$ ($d=2$) with $\sigma^{2}=1$, $\kappa\in\{0,1,2\}$, and $\mu\in\{4,8\}$.
+Each realization is simulated via Cholesky decomposition at $n=2500$ locations randomly sampled uniformly over the unit square.
+Throughout, we treat $\kappa$ as known and focus on ML estimation of $\sigma^{2}$, $a$, $\mu$, and the microergodic quantity $\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1}$.
+To create a gradual transition from increasing-domain to fixed-domain behavior while keeping the sampling design fixed, we follow Zhang and Zimmerman (2005); Kaufman and Shaby (2013) and vary the scale parameter $a$.
+Specifically, we consider $a\in\{0.1,0.3,0.6,0.9\}$.
+Here $a=0.1$ corresponds to weak dependence relative to the unit-square domain and thus approximates an increasing-domain regime, whereas $a=0.9$ induces strong dependence and is much closer to a fixed-domain setting.
+Table 1 reports the empirical bias and standard deviation (in parentheses) of the $500$ ML estimates $\widehat{\sigma}^{2}_{i}$, $\widehat{a}_{i}$, $\widehat{\mu}_{i}$ and $\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$, $i=1,\ldots,500$.
+The main pattern is clear: the individual covariance parameters $\sigma^{2}$, $a$, and $\mu$ become more variable as $a$ increases, whereas the microergodic parameter behaves in the opposite way.
+Table 1: Bias and standard deviation (in parentheses) of ML estimates $\widehat{\sigma}^{2}_{i}$, $\widehat{a}_{i}$$\widehat{\mu}_{i}$, and $\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ on $0,1]^{2}$ ($n=2500$ locations), with $\sigma^{2}=1$, $\kappa\in\{0,1,2\}$, $\mu\in\{4,8\}$, and $a\in\{0.1,0.3,0.6,0.9\}$.
 $\kappa=0$ $\kappa=1$ $\kappa=2$
-$\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$
-$\widehat{\sigma}^{2}$ $a=0.1$ $0.0037$ $0.0025$ $0.0018$ $0.0060$ $0.0039$ $0.0028$ $0.0053$ $0.0038$ $0.0029$
-$(0.0422)$ $(0.0382)$ $(0.0368)$ $(0.0485)$ $(0.0427)$ $(0.0398)$ $(0.0475)$ $(0.0432)$ $(0.0407)$
-$a=0.25$ $0.0030$ $0.0047$ $0.0046$ $0.0085$ $0.0098$ $0.0095$ $0.0100$ $0.0116$ $0.0111$
-$(0.0808)$ $(0.0622)$ $(0.0522)$ $(0.1082)$ $(0.0870)$ $(0.0737)$ $(0.1044)$ $(0.088)$ $(0.0773)$
-$a=0.5$ 0.0023-0.0023 $0.0024$ $0.0045$ $0.0029$ $0.0079$ $0.0101$ $0.0052$ $0.0091$ $0.0116$
-$(0.1446)$ $(0.1103)$ $(0.0900)$ $(0.1921)$ $(0.1592)$ $(0.135)$ $(0.1892)$ $(0.1599)$ $(0.1397)$
-$\widehat{a}$ $a=0.1$ $0.0006$ $0.0007$ $0.0007$ $0.0004$ $0.0004$ $0.0004$ $0.0002$ $0.0003$ $0.0003$
-$(0.0049)$ $(0.0055)$ $(0.0068)$ $(0.0024)$ $(0.0028)$ $(0.0033)$ $(0.0017)$ $(0.0021)$ $(0.0026)$
-$a=0.25$ $0.0007$ $0.0014$ $0.0016$ $0.0005$ $0.0008$ $0.0009$ $0.0004$ $0.0006$ $0.0007$
-$(0.0214)$ $(0.0176)$ $(0.0157)$ $(0.0099)$ $(0.0086)$ $(0.0079)$ $(0.0059)$ $(0.0054)$ $(0.0052)$
-$a=0.5$ 0.0015-0.0015 $0.0010$ $0.0023$ 0.0016-0.0016 0.0001-0.0001 $0.0008$ 0.0009-0.0009 0.0001-0.0001 $0.0005$
-$(0.0737)$ $(0.0578)$ $(0.0488)$ $(0.0329)$ $(0.0280)$ $(0.0245)$ $(0.0196)$ $(0.0171)$ $(0.0154)$
-$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ 0.0150-0.0150 0.0219-0.0219 0.0151-0.0151 3.9684-3.9684 4.4120-4.4120 3.7659-3.7659 496.06-496.06 486.01-486.01 346.94-346.94
-$(0.4370)$ $(0.5362)$ $(0.6792)$ $(51.3695)$ $(68.6746)$ $(88.9434)$ $(6397.8)$ $(8930.7)$ $(11628.7)$
-$a=0.25$ $0.0045$ $0.0015$ 0.0017-0.0017 $0.0228$ 0.0470-0.0470 0.1218-0.1218 0.2553-0.2553 1.7130-1.7130 3.0083-3.0083
-$(0.1499)$ $(0.1600)$ $(0.1708)$ $(2.5102)$ $(2.7402)$ $(2.9959)$ $(41.521)$ $(46.056)$ $(51.067)$
-$a=0.5$ $0.0034$ $0.0033$ $0.0026$ $0.0113$ $0.0099$ $0.0065$ $0.0399$ $0.0319$ $0.0150$
-$(0.0719)$ $(0.0740)$ $(0.0765)$ $(0.2923)$ $(0.3050)$ $(0.3187)$ $(1.1922)$ $(1.2484)$ $(1.3091)$
-Overall, the ML estimators of $\sigma^{2}$ and $a$ are approximately unbiased across all configurations.
-As expected, their standard deviations increase when moving from the increasing-domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$).
-For fixed $\kappa$, larger values of $\mu$ generally reduce variability.
-Regarding the microergodic parameter, under the increasing-domain scenario ($a=0.1$) the estimator exhibits moderate bias and relatively high variance, particularly for larger $\kappa$.
-Under the fixed-domain scenario ($a=0.5$), both bias and variance decrease substantially, and the sampling distribution appears closer to Gaussian, in agreement with the asymptotic theory.
-Moreover, the variability of the microergodic estimator increases with the smoothness parameter $\kappa$.
-Figure 4 displays boxplots of the ML estimates of $\sigma^{2}$, $a$, and $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (left) and $a=0.5$ (right), when $\kappa=0$.
-Under the increasing-domain scenario, the distributions of $\widehat{\sigma}^{2}$ and $\widehat{a}$ are approximately Gaussian.
-In the fixed-domain setting, the microergodic estimator becomes more concentrated around the true value and exhibits an approximately Gaussian distribution, as predicted by the theory.
+$\mu=4$ $\mu=8$ $\mu=4$ $\mu=8$ $\mu=4$ $\mu=8$
+$\widehat{\sigma}^{2}$ $a=0.1$ 0.00299 0.00107 0.00433 0.00184 0.00408 0.00212
+(0.04138) (0.03205) (0.04914) (0.03759) (0.04794) (0.03889)
+$a=0.3$ 0.01575 0.00504 0.01504 0.00770 0.01173 0.00726
+(0.10407) (0.05976) (0.13271) (0.08844) (0.12980) (0.09514)
+$a=0.6$ 0.06435 0.01181 0.04389 0.01385 0.03073 0.01553
+(0.26849) (0.11312) (0.25614) (0.16863) (0.23655) (0.18127)
+$a=0.9$ 0.12097 0.02437 0.09481 0.01848 0.05189 0.02400
+(0.38703) (0.17580) (0.41788) (0.24205) (0.32244) (0.25633)
+$\widehat{a}$ $a=0.1$ 0.02090 0.08978 0.00276 0.01635 0.00137 0.00612
+(0.10465) (0.20471) (0.01397) (0.06865) (0.00827) (0.02363)
+$a=0.3$ 0.81440 0.32342 0.02153 0.05872 0.00511 0.01092
+(2.75327) (0.75816) (0.10263) (0.26379) (0.04194) (0.06661)
+$a=0.6$ 3.91933 2.29317 0.55575 0.45367 0.04601 0.07026
+(8.37251) (4.53301) (2.64211) (1.50576) (0.22104) (0.33102)
+$a=0.9$ 7.46177 4.50776 1.66114 1.55866 0.16050 0.22433
+(14.27313) (7.73620) (5.73980) (4.14495) (0.77112) (0.80880)
+$\widehat{\mu}$ $a=0.1$ 1.00129 8.12410 0.17091 1.81217 0.10804 0.81383
+(5.12648) (18.51370) (0.88371) (7.52792) (0.66348) (3.13054)
+$a=0.3$ 12.17228 9.21399 0.42205 1.98131 0.12180 0.42181
+(41.02159) (21.40586) (2.01811) (8.89166) (0.96416) (2.54729)
+$a=0.6$ 29.22591 32.41524 5.42300 7.57716 0.55245 1.34734
+(61.35913) (63.46363) (25.61297) (24.98706) (2.54928) (6.27535)
+$a=0.9$ 36.69791 42.83158 10.93102 17.38350 1.29228 2.84537
+(68.21306) (72.15506) (37.54848) (45.95964) (6.09577) (10.03183)
+$\frac{\widehat{\sigma}^{2}F(\kappa,\mu,d)}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ 0.02847 -0.15330 567.68210 7656.58100 $1.40740\times 10^{7}$ $4.32344\times 10^{8}$
+(1.79328) (5.08178) (12695.11000) (107860.90000) ($2.18681\times 10^{8}$) ($3.79118\times 10^{9}$)
+$a=0.3$ 0.00880 -0.04215 -0.59747 -0.93738 -2994.48500 -48566.15000
+(0.45226) (0.95325) (296.09960) (1651.58800) (448665.40000) (4813284.00000)
+$a=0.6$ 0.00331 -0.01499 0.66440 -0.59339 165.89200 -325.76500
+(0.21584) (0.43426) (32.70379) (167.22660) (11788.42000) (115466.10000)
+$a=0.9$ 0.00148 -0.01088 0.05882 -0.61039 -4.95065 -62.59252
+(0.14205) (0.28237) (9.47927) (46.79391) (1482.45500) (13819.55000)
+This is especially evident for $\widehat{a}$ and $\widehat{\mu}$.
+For instance, when $\kappa=0$ and $\mu=4$, the empirical standard deviation of $\widehat{a}$ increases from $0.10465$ at $a=0.1$ to $2.75327$ at $a=0.3$, then to $8.37251$ at $a=0.6$, and to $14.27313$ at $a=0.9$; the standard deviation of $\widehat{\mu}$ rises from $5.12648$ to $41.02159$, then to $61.35913$, and finally to $68.21306$.
+The estimator $\widehat{\sigma}^{2}$ is more stable, but it follows the same qualitative trend.
+These results are fully consistent with fixed-domain asymptotics: the individual covariance parameters are not consistently estimable, and stronger dependence effectively reduces the amount of independent information.
+The effect of $\mu$ is visible but secondary relative to the role of $a$.
+In several configurations, $\mu=8$ leads to somewhat larger variability than $\mu=4$, especially for $\widehat{\mu}$ and in the stronger-dependence regimes.
+Likewise, the overall inferential difficulty increases with $\kappa$, particularly for quantities involving the scale parameter.
+The microergodic estimator behaves in the opposite way. For
+$\widehat{\theta}=\frac{\widehat{\sigma}^{2}F(\kappa,\widehat{\mu},d)}{\widehat{a}^{2\kappa+1}},$
+both bias and dispersion decrease markedly as $a$ increases.
+This is most striking for larger $\kappa$.
+For instance, when $\kappa=2$ and $\mu=4$, the empirical standard deviation of $\widehat{\theta}$ drops from $2.18681\times 10^{8}$ at $a=0.1$, to $4.48665\times 10^{5}$ at $a=0.3$, to $1.17884\times 10^{4}$ at $a=0.6$, and to $1482.45500$ at $a=0.9$.
+Even for $\kappa=0$, where $\widehat{\theta}$ is already much more stable, the same monotone improvement is evident.
+This contrast is the central empirical message of the simulation study.
+As the regime moves from increasing-domain-like to fixed-domain-like behavior, the individual covariance parameters deteriorate, while the microergodic parameter improves.
+The distinction becomes sharper as $\kappa$ increases, since the factor $a^{-(2\kappa+1)}$ amplifies scale estimation error in weak-dependence settings, but defines the identifiable combination under fixed-domain asymptotics.
+Figure 4 shows normal QQ-plots for the standardized microergodic estimates in the strong-dependence setting ($a=0.9$) with $\kappa=1$, comparing $\mu=4$ (left) and $\mu=8$ (right).
+Specifically, we plot
+$Z_{i}\;=\;\frac{\sqrt{n}\,\big(\widehat{\theta}_{i}-\theta_{0}\big)}{\sqrt{2}\,\theta_{0}},\qquad\widehat{\theta}_{i}=\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1},\quad\theta_{0}=\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1},$
+for $i=1,\dots,500$. In both cases, the QQ-plots align reasonably well with the standard normal reference line, supporting the relevance of the fixed-domain Gaussian approximation at this sample size.
+Overall, the simulations strongly support the theoretical results of Section 6: the individual covariance parameters are increasingly unstable as dependence strengthens, whereas the microergodic parameter shows the opposite trend and is the meaningful inferential target under fixed-domain asymptotics.
 Refer to caption Refer to caption
-($a=0.1$) ($a=0.5$)
-Figure 4: Left part: Boxplots of ML estimates of $\sigma^{2}$, $a$ and the microergodic parameter $\frac{\sigma^{2}}{a^{2\kappa+1}}$ (from left to right) when estimating the covariance model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $\sigma^{2}=1$, for $\mu=4,6,8$ and $a=0.1$ (increasing domain scenario).
-Right part: the same Figure with $a=0.5$ (fixed domain scenario).
-In both cases $\kappa=0$.
+($\mu=4$) ($\mu=8$)
+Figure 4: Normal $Q$-plots of standardized microergodic estimates in the fixed-domain setting ($a=0.9$) with $\kappa=1$ and $\sigma^{2}=1$, comparing $\mu=4$ (left) and $\mu=8$ (right).
+The standardized statistics are $Z_{i}=\sqrt{n}\,(\widehat{\theta}_{i}-\theta_{0})/(\sqrt{2}\,\theta_{0})$, where $\widehat{\theta}_{i}=\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$, $\theta_{0}=\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1}$, and $n=2500$.
 8 A real data application
@@
 Since the transformation is monotone, the Tukey-$h$ RF preserves the geometric properties of the underlying Gaussian field.
-For the correlation of $Z(\mathbf{s})$, we compare three models: ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$, and the proposed ${\cal H}_{\kappa,\mu,\beta,2}$.
-Here $\kappa$ denotes smoothness, $\beta$ the scale parameter, and $\mu$ the shape parameter.
+For the correlation of $Z(\mathbf{s})$, we compare three models: ${\cal MT}_{\kappa,a}$, ${\cal GW}_{\kappa,\mu,a}$, and the proposed ${\cal H}_{\kappa,\mu,a,2}$.
+Here $\kappa$ denotes smoothness, $a$ the scale parameter, and $\mu$ the shape parameter.
 We adopt an increasing-domain framework and estimate all covariance parameters by ML.
@@
 Overall, the AIC criterion favors the Tukey-$h$ specification, irrespective of the correlation model.
-Among the correlation structures, the compactly supported ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$ achieve the best AIC values.
-Figure 6 compares the normalized histogram of the residuals with the fitted density (left panels) and the empirical semivariogram with the fitted semivariogram (right panels), for the Gaussian (first row) and Tukey-$h$ (second row) models using the estimated ${\cal H}_{\kappa,\mu,\beta,2}$ correlation.
+Among the correlation structures, the compactly supported ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$ achieve the best AIC values.
+Figure 6 compares the normalized histogram of the residuals with the fitted density (left panels) and the empirical semivariogram with the fitted semivariogram (right panels), for the Gaussian (first row) and Tukey-$h$ (second row) models using the estimated ${\cal H}_{\kappa,\mu,a,2}$ correlation.
 The Tukey-$h$ RF provides a noticeably better fit to both marginal tail behavior and spatial dependence.
-Table 2: ML estimates with associated standard error (in parentheses), RMSE, MAE and percentage of zero values in the covariance matrix for the Gaussian and Tukey-$h$ RFs with underlying correlation functions ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$.
+Table 2: ML estimates with associated standard error (in parentheses), RMSE, MAE and percentage of zero values in the covariance matrix for the Gaussian and Tukey-$h$ RFs with underlying correlation functions ${\cal MT}_{\kappa,a}$, ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$.
 Last row reports the time (in seconds) needed to compute the Cholesky factor of the covariance matrix.
 Gaussian RF Tukey-$h$ RF
-${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$ ${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$
+${\cal MT}_{\kappa,a}$ ${\cal GW}_{\kappa,\mu,a}$ ${\cal H}_{\kappa,\mu,a,2}$ ${\cal MT}_{\kappa,a}$ ${\cal GW}_{\kappa,\mu,a}$ ${\cal H}_{\kappa,\mu,a,2}$
 $\sigma^{2}$ $1.2054$ $1.2066$ $1.2061$ $0.6998$ $0.6986$ $0.6984$
@@
 $(0.0293)$ $(0.0307)$ $(0.0312)$ $(0.0265)$ $(0.0237)$ $(0.0237)$
-$\beta$ $16.0229$ $85.8058$ $87.9751$ $18.5405$ $76.0127$ $77.6552$
+$a$ $16.0229$ $85.8058$ $87.9751$ $18.5405$ $76.0127$ $77.6552$
 $(1.3104)$ $(20.4440)$ $(20.9644)$ $(1.6183)$ $(12.0024)$ $(10.5911)$
@@
 $h$ $0.1885$ $0.1899$ $0.1898$
-$(0.0172)$ $(0.0170)$ $0.0170$
+$(0.0172)$ $(0.0170)$ $(0.0170)$
 loglik 3900.01-3900.01 3896.88-3896.88 3896.94-3896.94 3689.88-3689.88 3683.81-3683.81 3683.94-3683.94
@@
 Table 2 reports the average RMSE and MAE.
-The best predictive performance is achieved by the Tukey-$h$ RF combined with compactly supported correlations ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$.
+The best predictive performance is achieved by the Tukey-$h$ RF combined with compactly supported correlations ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$.
 […]
@@
 9 Concluding Remarks
-In this paper we study the Gauss hypergeometric covariance model, a flexible class introduced by Emery and Alegría (2022) that encompasses both the Matérn and Generalized Wendland families.
-We derive necessary and sufficient validity conditions and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics.
-To address this issue, we introduce a parsimonious compactly supported subclass selected through an optimality criterion based on maximum integral range.
-In this sense, the proposed hypergeometric model can be viewed as a structural refinement of the Generalized Wendland construction.
-Although simple closed-form expressions are not generally available for nonnegative integer smoothness parameters, efficient numerical evaluation of the Gauss hypergeometric function is widely supported in modern scientific computing libraries.
-In particular, reliable implementations are available in the GNU Scientific Library (Gough, 2009) and in major environments such as R, MATLAB, and Python.
-The GeoModels package (Bevilacqua et al., 2026), used in this paper for numerical experiments and applications, computes the proposed model via the SciPy implementation of the hypergeometric function (Virtanen et al., 2020).
+We revisited the Gauss hypergeometric covariance class (Emery and Alegría, 2022), a flexible family that includes the celebrated Matérn and generalized Wendland models as special cases.
+We first refined its validity conditions, deriving sharp constraints over a substantial region of the parameter space and pinpointing where a gap between necessary and sufficient conditions still remains.
+We then clarified that the class features two distinct non-identifiability mechanisms: a structural one, stemming from an intrinsic symmetry in the shape parameters, and the classical fixed-domain one, whereby different parameter values can induce equivalent Gaussian measures and only microergodic combinations are consistently estimable.
+The structural ambiguity is removed by an ordering constraint, yielding a unique representation on the restricted space.
+Within this identifiable formulation, we introduced a parsimonious compactly supported subclass, $\mathcal{H}$, selected via a maximum integral range criterion.
+This choice provides a canonical representative with maximal low-frequency power among admissible models with fixed smoothness and support.
+The resulting family generalizes the classical spherical model (Chilès and Delfiner, 2012) while retaining continuous smoothness control, compact support, and simple necessary and sufficient validity conditions.
+We also derived scale-mixture representations and explicit links to the Matérn model through compact-support reparameterizations.
+From an inferential standpoint, we developed fixed-domain asymptotics for likelihood-based estimation under $\mathcal{H}$.
+For fixed smoothness, we proved consistency and asymptotic normality for the microergodic-parameter estimator, and established stability when nuisance parameters are estimated by ML over compact sets.
+Simulations and a climate-data application indicate that $\mathcal{H}$ can offer a competitive compromise between Matérn and generalized Wendland models, combining accuracy with sparse covariance matrices and notable computational gains, for both Gaussian and Tukey-hh random fields.
+Several directions remain open, including closing the remaining validity gap in Theorem 2, extending fixed-domain theory to unknown smoothness, and developing scalable alternatives to full ML.
+Finally, anisotropic, nonstationary, and multivariate extensions would further broaden the practical scope of the hypergeometric framework.
+Although closed forms are available only in special cases, efficient numerical evaluation of Gauss hypergeometric functions is widely supported, including in the GNU Scientific Library (Gough, 2009) and in major environments such as R, MATLAB, and Python.
+The GeoModels package (Bevilacqua et al., 2026) used in this paper computes the proposed model via the SciPy implementation (Virtanen et al., 2020).
 Acknowledgement
@@
 A.1 Proof of Theorem 1
-Proof.
-Let $P(\rho_{i})$, $i=0,1$, be two zero-mean Gaussian measures with isotropic covariance function $\rho_{i}$ and associated spectral density $\widehat{\rho}_{i}$, $i=0,1$, as defined through (1).
-Using Skorokhod and Yadrenko (1973); Ibragimov and Rozanov (1978); Stein (2004), if for some $a>0$ the function $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and if for some finite and positive $c$,
-$\int_{c}^{\infty}z^{d-1}\left(\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right)^{2}\,{\rm d}z<\infty,$ (26)
-then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$.
-Condition (26) is the key tool.
-We first need the following result.
+A preliminary Lemma is needed for the proof of Theorem 1.
+The following result can be obtained from the asymptotic expansion of the generalized hypergeometric function $_{1}F_{2}$ for large argument, as given in Luke (1969).
 Lemma 1.
-Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be defined at (7). Assume $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $2(\beta+\gamma)\geq 6\delta+1$. Then:
+Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be defined as in (7). Assume $\delta>\frac{d}{2}$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$, and $2(\beta+\gamma)\geq 6\delta+1$. Then:
   1. 1.
-    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ $=\sigma^{2}Ka^{d}\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
-    $\qquad+c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big],$
-    as $z\to\infty$;
+    As $z\to\infty$,
+    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=K_{\delta,d}\Big[c_{3}\,(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}+c_{4}\,(za)^{\frac{1}{2}+\delta-\beta-\gamma}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big],$
+    where
+    $c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)},\qquad c_{4}=\frac{\pi^{-1/2}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)},\qquad c_{5}=\frac{\pi}{2}\!\left(\frac{1}{2}+\delta-\beta-\gamma\right).$
   2. 2.
-    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta},\qquad z\to\infty.$
-Here
-$c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)},\qquad c_{4}=\frac{\pi^{-1/2}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)},$
-$c_{5}=\frac{\pi}{2}\left(\frac{1}{2}+\delta-\beta-\gamma\right).$
-We are now ready to prove the theorem.
-From Lemma 2 (Point 2), there exist constants $0<c_{1}<c_{2}<\infty$ such that, for $z$ large,
-$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}(z)\leq c_{2},\qquad i=0,1.$ (27)
-In particular, $\widehat{\rho}_{0}(z)z^{2\delta}$ is bounded away from 0 and $\infty$ as $z\to\infty$.
-Using Lemma 2 (Points 1 and 2), and collecting terms, one obtains for large $z$
-$\left|\frac{\sigma_{1}^{2}\widehat{\varphi}_{1}(z)-\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}{\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}\right|\leq\frac{1}{c_{1}}\left|A(z)+B(z)+D(z)\right|,$
-where
-$A(z)$ $=\big[c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big]\{1+\mathcal{O}(z^{-2})\},$
-$B(z)$ $=\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z+c_{5}^{1})$
-$\quad-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z+c_{5}^{0})\big],$
-$D(z)$ $=\mathcal{O}(z^{-1})\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\big].$
-Hence condition (26) becomes
-$\int_{c}^{\infty}z^{d-1}\big(A(z)+B(z)+D(z)\big)^{2}\,{\rm d}z<\infty.$
-Assume now (10), i.e.
-$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}}.$
-This is equivalent to
-$c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}=c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0},$
-so that $A(z)\equiv 0$.
-Using $(x+y)^{2}\leq 2x^{2}+2y^{2}$, it suffices to check
-$\int_{c}^{\infty}z^{d-1+2(\lambda_{i}+2\delta)}\,dz<\infty,\qquad i=0,1,$
-which holds iff
-$\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta.$
-Therefore, under
-$\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})>\frac{d+1}{2}+3\delta,$
-condition (26) holds and the sufficient part is proved.
-The necessary part follows from Zhang (2004).
+    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta}$ as $z\to\infty$.
+Proof of Theorem 1.
+Let
+$\rho_{i}(\boldsymbol{h})=\sigma_{i}^{2}\,\varphi_{i}(\|\boldsymbol{h}\|),\qquad\varphi_{i}(x)={\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}}(x),\qquad i=0,1,$
+and let $f_{i}$ denote the (isotropic) spectral density of $\rho_{i}$ in the sense of (1).
+Then $f_{i}(z)=\sigma_{i}^{2}\,\widehat{\varphi}_{i}(z)$, where $\widehat{\varphi}_{i}(z)=\widehat{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}}(z)$ is given by (7).
+Since we are considering valid covariance models, $\widehat{\varphi}_{i}(z)\geq 0$ for all $z\geq 0$, and the Gamma factors in (7) are finite (in particular, $\beta_{i}>d/2$, $\gamma_{i}>d/2$ and $\beta_{i}>\delta$, $\gamma_{i}>\delta$).
+We use the standard equivalence criterion for stationary Gaussian measures on bounded domains (see, e.g., Ibragimov and Rozanov (1978) and Stein (2004)): if for some $\eta>0$ the function $f_{0}(z)\,z^{\eta}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and if for some $c>0$
+$\int_{c}^{\infty}z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}\,dz<\infty,$ (30)
+then $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent on the paths over any bounded $D\subset\mathbb{R}^{d}$.
+(For $d=1,2,3$ this condition is also sharp in the present regularly-varying setting; see the necessity argument below.)
+Set, for $i=0,1$,
+$\lambda_{i}:=\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}.$
+By assumption, for each $i=0,1$ we have $\delta>\frac{d}{2}$ and $2(\beta_{i}-\delta)(\gamma_{i}-\delta)\geq\delta.$ Moreover, the theorem assumes $\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta,$ hence $2(\beta_{i}+\gamma_{i})>(d+1)+6\delta\geq 6\delta+2>6\delta+1,$ so that in particular $2(\beta_{i}+\gamma_{i})\geq 6\delta+1$.
+Therefore all assumptions of Lemma 1 hold for $\widehat{\varphi}_{i}$.
+By Lemma 1 (applied to $\widehat{\varphi}_{i}$), as $z\to\infty$,
+$\widehat{\varphi}_{i}(z)=K_{\delta,d}\,L(\beta_{i},\gamma_{i})\,a_{i}^{\,d-2\delta}\,z^{-2\delta}\,\{1+\mathcal{O}(z^{-2})\}\;+\;\mathcal{O}\!\big(z^{\lambda_{i}}\big),$ (31)
+where $K_{\delta,d}:=\Gamma(\delta)/(2^{2\delta}\pi^{d/2}\Gamma(\delta-d/2))>0$ depends only on $(\delta,d)$ and
+$L(\beta_{i},\gamma_{i})=\frac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}.$
+Multiplying by $\sigma_{i}^{2}$, we obtain for $f_{i}(z)=\sigma_{i}^{2}\widehat{\varphi}_{i}(z)$:
+$f_{i}(z)=A_{i}\,z^{-2\delta}\{1+\mathcal{O}(z^{-2})\}+\mathcal{O}\!\big(z^{\lambda_{i}}\big),\qquad A_{i}:=\sigma_{i}^{2}\,K_{\delta,d}\,L(\beta_{i},\gamma_{i})\,a_{i}^{\,d-2\delta}.$ (32)
+Since the theorem assumes $\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta,$ we have
+$\lambda_{i}+2\delta=\frac{1}{2}+3\delta-(\beta_{i}+\gamma_{i})<-\frac{d}{2},$
+hence $z^{\lambda_{i}}=o(z^{-2\delta})$ and (32) yields
+$f_{i}(z)\sim A_{i}\,z^{-2\delta}\qquad(z\to\infty),\quad\text{with }A_{i}>0.$ (33)
+In particular, $f_{0}(z)\,z^{2\delta}$ is bounded away from 0 and $\infty$ for large $z$, so (30) can be applied with $\eta=2\delta$.
+Sufficiency.
+Assume (10).
+Since $K_{\delta,d}>0$ is common to both models, (10) is equivalent to $A_{0}=A_{1}$ (because $A_{i}$ differs from $\sigma_{i}^{2}L(\beta_{i},\gamma_{i})/a_{i}^{2\delta-d}$ only by the common factor $K_{\delta,d}$).
+Hence, by (32),
+$f_{1}(z)-f_{0}(z)=A_{0}z^{-2\delta}\mathcal{O}(z^{-2})\;+\;\mathcal{O}\!\big(z^{\lambda_{0}}\big)\;+\;\mathcal{O}\!\big(z^{\lambda_{1}}\big).$
+Dividing by $f_{0}(z)\sim A_{0}z^{-2\delta}$ gives
+$\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}=\mathcal{O}(z^{-2})+\mathcal{O}\!\big(z^{\lambda_{0}+2\delta}\big)+\mathcal{O}\!\big(z^{\lambda_{1}+2\delta}\big).$
+Therefore, for large $z$,
+$z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}\;\lesssim\;z^{d-1}z^{-4}\;+\;z^{d-1}z^{2(\lambda_{0}+2\delta)}\;+\;z^{d-1}z^{2(\lambda_{1}+2\delta)}.$
+The first term is integrable at $\infty$ iff $d<4$, i.e. $d=1,2,3$. For the other terms, note that
+$d-1+2(\lambda_{i}+2\delta)=d-1+2\Big(\tfrac{1}{2}+3\delta-(\beta_{i}+\gamma_{i})\Big)=d+6\delta-2(\beta_{i}+\gamma_{i}),$
+which is <1<-1 exactly when $\beta_{i}+\gamma_{i}>(d+1)/2+3\delta$, assumed in the theorem.
+Thus the integral (30) is finite, and by the criterion $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent on bounded $D$.
+Necessity: Assume the theorem’s condition $\beta_{i}+\gamma_{i}>(d+1)/2+3\delta$ holds for $i=0,1$ and suppose (10) fails. Then $A_{1}\neq A_{0}$. Using (33),
+$\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}=\frac{A_{1}z^{-2\delta}-A_{0}z^{-2\delta}+o(z^{-2\delta})}{A_{0}z^{-2\delta}\{1+o(1)\}}\longrightarrow\frac{A_{1}-A_{0}}{A_{0}}\neq 0\qquad(z\to\infty).$
+Hence there exist $c>0$ and $\varepsilon>0$ such that for all $z\geq c$,
+$\left|\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right|\geq\varepsilon,$
+and therefore
+$\int_{c}^{\infty}z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}dz\;\geq\;\varepsilon^{2}\int_{c}^{\infty}z^{d-1}\,dz=\infty.$
+Thus (30) fails and the measures cannot be equivalent. Consequently, equivalence implies $A_{1}=A_{0}$, which is exactly (10).
+Combining the sufficiency and necessity arguments proves the claimed “if and only if” statement. ∎
 A.2 Proof of Theorem 2
 Proof.
-Consider the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ in dimension $d$, where $\delta=\frac{d+1}{2}+\kappa$, $\kappa>-\frac{1}{2}$, $l\geq 0$.
+Consider the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ in dimension $d$, where $\delta=\frac{d+1}{2}+\kappa$, $\kappa>-\frac{1}{2}$.
+Note that the model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under interchanging $\beta$ and $\gamma$, which corresponds to $l\mapsto-l$.
+The admissible region is therefore symmetric with respect to $l=0$, and restricting to $l\geq 0$ entails no loss of generality.
+We henceforth assume $l\geq 0$.
 Sufficiency. Under $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, the sufficient conditions of Section 3 become:
@@
 In Regime 2 ($l>\frac{d}{2}+\kappa$), this provides a necessary lower bound that is strictly weaker than the sufficient condition in point 2; whether the gap can be closed remains an open question.
-Reduction to $l\geq 0$.
-The model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under interchanging $\beta$ and $\gamma$, which corresponds to $l\mapsto-l$.
-The admissible region is therefore symmetric with respect to $l=0$, and restricting to $l\geq 0$ entails no loss of generality.
 This completes the proof. ∎
@@
 Proof.
-Without loss of generality set $a=1$; the general case follows by the substitution $x\mapsto ax$, which contributes the factor $a^{d}$.
-Using the change of variable $t=1-x^{2}$ (so $x^{d-1}\,dx=\tfrac{1}{2}(1-t)^{(d-2)/2}\,dt$) in the definition of the integral range (12), and substituting (8), one obtains
-$A_{\kappa,\mu,l,1,d}$ $=\frac{\pi^{d/2}C_{0}}{2^{d}\Gamma(d/2)}\int_{0}^{1}(1-t)^{\frac{d-2}{2}}\,t^{\mu+\frac{d-1}{2}+2\kappa}$ (28)
-$\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;t\right)\,dt.$
-where $C_{0}=\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)\big/\!\left[\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\tfrac{1}{2}\right)\right]$.
-We apply formula 2.21.1.11 of Prudnikov et al. (1990),
-$\int_{0}^{1}t^{\sigma-1}(1-t)^{\rho-1}\,{}_{2}F_{1}(a,b;c;t)\,dt=\frac{\Gamma(\rho)\Gamma(\sigma)}{\Gamma(\rho+\sigma)}\,{}_{3}F_{2}(a,b,\sigma;\,c,\,\rho+\sigma;\,1),$ (29)
-with the identifications
-$\sigma=\mu+\tfrac{d+1}{2}+2\kappa,\quad\rho=\tfrac{d}{2},\quad a=\tfrac{\mu}{2},\quad b=\tfrac{\mu+d}{2}+\kappa,\quad c=\mu+\tfrac{d+1}{2}+2\kappa.$
-Since $c=\sigma$, the third upper parameter and the first lower parameter of $_{3}F_{2}$ coincide, so the series telescopes term by term:
-$_{3}F_{2}(a,b,\sigma;\,\sigma,\,\rho+\sigma;\,1)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}(\sigma)_{k}}{k!\,(\sigma)_{k}(\rho+\sigma)_{k}}=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{k!\,(\rho+\sigma)_{k}}={}_{2}F_{1}(a,b;\,\rho+\sigma;\,1).$
-Gauss’s summation theorem then gives, provided $\rho+\sigma-a-b>0$,
-$_{2}F_{1}(a,b;\rho+\sigma;1)=\frac{\Gamma(\rho+\sigma)\,\Gamma(\rho+\sigma-a-b)}{\Gamma(\rho+\sigma-a)\,\Gamma(\rho+\sigma-b)}.$
-One verifies directly that
-$\rho+\sigma-a-b=\kappa+\tfrac{1}{2}+l>0\quad(\kappa>-\tfrac{1}{2},\;l\geq 0),$
-$\rho+\sigma-a=\tfrac{\mu+1}{2}+\kappa+d,\qquad\rho+\sigma-b=\tfrac{\mu+1}{2}+\kappa+\tfrac{d}{2}.$
-Substituting back into (28), the factors $\Gamma(\sigma)$ and $\Gamma(\rho+\sigma)$ cancel with those arising from (29) and from $C_{0}$.
-Using the duplication formula $\Gamma(z)\Gamma(z+\tfrac{1}{2})=\sqrt{\pi}\,2^{1-2z}\Gamma(2z)$ to simplify the remaining Gamma factors, one obtains
-$A_{\kappa,\mu,l,1,d}=\frac{\pi^{d/2}\,\Gamma\!\left(\frac{d+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\,\Gamma\!\left(\kappa+\frac{1}{2}\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa+l\right)},$
-which, after restoring the factor $a^{d}$, gives (14). ∎
+Since $\phi(\|\cdot\|)$ is compactly supported it is absolutely integrable and, by the spectral representation associated with (1),
+$A=(2\pi)^{d}\,\widehat{\phi}(0).$
+For $\phi(x)=\mathcal{GH}_{\delta,\beta,\gamma,a}(x)$, the spectral density is (7). Evaluating at $z=0$ and using $_{1}F_{2}(\cdot;\,0)=1$ gives
+$\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(0)=\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}.$
+Hence
+$A=(2\pi)^{d}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(0)=\pi^{d/2}a^{d}\,\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})}{\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}.$
+Substituting $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+l$ yields (14). ∎
 A.4 Proof of Theorem 4
 Proof.
-Without loss of generality, set the compact support to $a=1$. Define
-$LA_{\kappa,d}(\mu,l):=\log(A_{\kappa,\mu,l,d,1}),$
+Since $A_{\kappa,\mu,l,a,d}=a^{d}\,A_{\kappa,\mu,l,1,d}$, the factor $a^{d}$ is irrelevant for monotonicity and maximizers in $\mu$. Hence we set $a=1$.
+Fix $l\geq 0$ and consider
+$LA_{\kappa,d}(\mu,l):=\log\!\big(A_{\kappa,\mu,l,1,d}\big),$
 where
-$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+\frac{d}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l+\frac{d}{2}\right)}.$
-We maximize $LA_{\kappa,d}(\mu,l)$ over the admissible set in $\mu$ for a fixed $l\geq 0$, as given by Theorem 2. Let $\delta=\frac{d}{2}+\kappa$. Then
-  1. 1.
-    $\mu\geq\delta-l+\tfrac{1}{2}$ if $0\leq l\leq\delta$;
-  2. 2.
-    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ if $l>\delta$.
-This is a maximization problem over the half-line $\mu_{\min}(l),\infty)$.
-Let $\psi(x)=\frac{d}{dx}\log\Gamma(x)$ be the digamma function, which is strictly increasing for $x>0$.
-Set $x=\frac{\mu+1}{2}+\kappa$.
-Under the validity conditions we have $x>0$ and $x+l>0$.
-Differentiating with respect to $\mu$ yields
-$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)=\frac{1}{2}\Big[\psi(x)-\psi(x+\tfrac{d}{2})+\psi(x+l)-\psi(x+l+\tfrac{d}{2})\Big].$
-Since $\psi$ is increasing and $\tfrac{d}{2}>0$, we have $\psi(x)-\psi(x+\tfrac{d}{2})\leq 0$ and $\psi(x+l)-\psi(x+l+\tfrac{d}{2})\leq 0$, hence $\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)\leq 0$ for all admissible $\mu$.
-Therefore $LA_{\kappa,d}(\mu,l)$ is non-increasing in $\mu$, and its maximum is achieved at the smallest admissible value of $\mu$, i.e.
-$\mu_{1}(l)=\delta-l+\tfrac{1}{2}\quad(0\leq l\leq\delta),\qquad\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l\quad(l>\delta).$
-This proves Theorem 4. ∎
+$A_{\kappa,\mu,l,1,d}=\frac{\pi^{d/2}\Gamma\!\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{\Gamma\!\left(\kappa+\frac{1}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+\frac{d}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l+\frac{d}{2}\right)}.$
+Let $\psi=\Gamma^{\prime}/\Gamma$ be the digamma function and set
+$x=\frac{\mu+1}{2}+\kappa\qquad\Big(\text{so }\frac{dx}{d\mu}=\frac{1}{2}\Big).$
+On the validity region of the model (hence in particular on the admissible domain in $\mu$), the Gamma factors above are well-defined and positive, so $x>0$ and $x+l>0$.
+Differentiating with respect to $\mu$ gives
+$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)=\frac{1}{2}\Big[\psi(x)-\psi\!\left(x+\tfrac{d}{2}\right)+\psi(x+l)-\psi\!\left(x+l+\tfrac{d}{2}\right)\Big].$
+Since $\psi$ is strictly increasing on $(0,\infty)$ and $d/2>0$, each difference $\psi(x)-\psi(x+d/2)$ and $\psi(x+l)-\psi(x+l+d/2)$ is $\leq 0$. Therefore
+$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)\leq 0\quad\text{for all admissible }\mu,$
+hence $\mu\mapsto A_{\kappa,\mu,l,a,d}$ is non-increasing on its admissible domain. Consequently, its maximum is attained at the smallest admissible value of $\mu$.
+We now identify this smallest admissible value using Theorem 2.
+If $0\leq l\leq d/2+\kappa$, Theorem 2 provides the exact validity condition $\mu\geq\mu_{1}(l):=\delta-l+\frac{1}{2}$, hence $\mu^{\star}(l)=\mu_{1}(l)$.
+If $l>d/2+\kappa$, Theorem 2 guarantees validity for all $\mu\geq\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l$.
+In this regime, we take as admissible domain the parameter range ensured by Theorem 2 (i.e., $\mu\geq\mu_{2}(l)$), and by monotonicity the maximum over this admissible domain is attained at $\mu^{\star}(l)=\mu_{2}(l)$.
+∎
 A.5 Proof of Theorem 5
 Proof.
-Let $\delta=\frac{d}{2}+\kappa$. Since $\log$ is strictly increasing, maximizing $A_{\kappa,\mu,l,a,d}$ is equivalent to maximizing
-$LA_{\kappa,d}(\mu,l):=\log A_{\kappa,\mu,l,d,1},$
+Let $\tau=\frac{d}{2}+\kappa$. Since $\log$ is strictly increasing, maximizing $A_{\kappa,\mu,l,a,d}$ is equivalent to maximizing
+$LA_{\kappa,d}(\mu,l):=\log A_{\kappa,\mu,l,1,d},$
 and the factor $a^{d}$ plays no role for maximizers.
-By Theorem 4, for each fixed $l$ the maximizer in $\mu$ is $\mu=\mu_{1}(l)$ for $0\leq l\leq\delta$ and $\mu=\mu_{2}(l)$ for $l>\delta$. Hence it suffices to study the two one–variable functions
-$F_{1}(l):=LA_{\kappa,d}(\mu_{1}(l),l)\quad\text{on }[0,\delta],\qquad F_{2}(l):=LA_{\kappa,d}(\mu_{2}(l),l)\quad\text{on }(\delta,\infty).$
+By Theorem 4, for each fixed $l$ the maximizer in $\mu$ is $\mu=\mu_{1}(l)$ for $0\leq l\leq\tau$ and $\mu=\mu_{2}(l)$ for $l>\tau$, where
+$\mu_{1}(l)=\tau+1-l\qquad(0\leq l\leq\tau),\qquad\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l\qquad(l>\tau).$
+Hence it suffices to study the two one-variable functions
+$F_{1}(l):=LA_{\kappa,d}(\mu_{1}(l),l)\quad\text{on }[0,\tau],\qquad F_{2}(l):=LA_{\kappa,d}(\mu_{2}(l),l)\quad\text{on }(\tau,\infty).$
 Write $\psi=\Gamma^{\prime}/\Gamma$ for the digamma function. Recall that $\psi$ is concave on $(0,\infty)$, so for every $h>0$ the symmetric increment
@@
 is decreasing in $t$ (because $\Delta_{h}^{\prime}(t)=\psi^{\prime}(t+h)-\psi^{\prime}(t-h)\leq 0$).
-Case $i=1$ ($0\leq l\leq\delta$).
-For $0\leq l\leq\delta$, $\mu_{1}(l)=\delta-l+\tfrac{1}{2}$. Set
-$u(l):=\frac{\mu_{1}(l)+1}{2}+\kappa=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}-\frac{l}{2}.$
-Using the explicit form of $A_{\kappa,\mu,l,d,1}$, $F_{1}(l)$ differs from
-$\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big)$
+Case $i=1$ ($0\leq l\leq\tau$).
+For $0\leq l\leq\tau$, $\mu_{1}(l)=\tau+1-l$. Set
+$u(l):=\frac{\mu_{1}(l)+1}{2}+\kappa=\frac{d}{4}+\frac{3\kappa}{2}+1-\frac{l}{2}.$
+Using the explicit form of $A_{\kappa,\mu,l,1,d}$, $F_{1}(l)$ differs from
+$\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\!\left(u(l)+\frac{d}{2}\right)-\log\Gamma\!\left(u(l)+l+\frac{d}{2}\right)$
 only by an additive constant independent of $l$. Differentiating gives
-$F_{1}^{\prime}(l)$ $=u^{\prime}(l)\Big[\psi(u(l))+\psi(u(l)+l)-\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big]$
-$\qquad+\psi(u(l)+l)-\psi\Big(u(l)+l+\frac{d}{2}\Big),$
+$F_{1}^{\prime}(l)$ $=u^{\prime}(l)\left[\psi(u(l))+\psi(u(l)+l)-\psi\!\left(u(l)+\frac{d}{2}\right)-\psi\!\left(u(l)+l+\frac{d}{2}\right)\right]$
+$\qquad+\psi(u(l)+l)-\psi\!\left(u(l)+l+\frac{d}{2}\right),$
 where $u^{\prime}(l)=-\tfrac{1}{2}$. A simple rearrangement yields
-$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(u(l)+l)-\psi(u(l))+\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big].$
+$F_{1}^{\prime}(l)=\frac{1}{2}\left[\psi(u(l)+l)-\psi(u(l))+\psi\!\left(u(l)+\frac{d}{2}\right)-\psi\!\left(u(l)+l+\frac{d}{2}\right)\right].$
 Now set
-$x:=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}\quad\Longrightarrow\quad u(l)=x-\frac{l}{2},\qquad u(l)+l=x+\frac{l}{2}.$
+$x:=\frac{d}{4}+\frac{3\kappa}{2}+1\quad\Longrightarrow\quad u(l)=x-\frac{l}{2},\qquad u(l)+l=x+\frac{l}{2}.$
 Therefore
-$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(x+\tfrac{l}{2})-\psi(x-\tfrac{l}{2})+\psi(x-\tfrac{l}{2}+\tfrac{d}{2})-\psi(x+\tfrac{l}{2}+\tfrac{d}{2})\Big]=\frac{1}{2}\Big(\Delta_{l/2}(x)-\Delta_{l/2}(x+\tfrac{d}{2})\Big).$
-Since $\Delta_{l/2}$ is decreasing and $x+\tfrac{d}{2}>x$, we have $\Delta_{l/2}(x)\geq\Delta_{l/2}(x+\tfrac{d}{2})$, hence $F_{1}^{\prime}(l)\geq 0$ on $0,\delta]$.
-Thus $F_{1}$ is non-decreasing on $0,\delta]$ and its maximum is attained at $l=\delta$.
-Case $i=2$ ($l>\delta$).
-For $l>\delta$, $\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l$. Define
+$F_{1}^{\prime}(l)=\frac{1}{2}\!\left[\psi\!\left(x+\tfrac{l}{2}\right)-\psi\!\left(x-\tfrac{l}{2}\right)+\psi\!\left(x-\tfrac{l}{2}+\tfrac{d}{2}\right)-\psi\!\left(x+\tfrac{l}{2}+\tfrac{d}{2}\right)\right]=\frac{1}{2}\!\left(\Delta_{l/2}(x)-\Delta_{l/2}\!\left(x+\tfrac{d}{2}\right)\right).$
+Since $\Delta_{l/2}$ is decreasing and $x+\tfrac{d}{2}>x$, we have $\Delta_{l/2}(x)\geq\Delta_{l/2}(x+\tfrac{d}{2})$, hence $F_{1}^{\prime}(l)\geq 0$ on $0,\tau]$.
+Thus $F_{1}$ is non-decreasing on $0,\tau]$ and its maximum is attained at $l=\tau$.
+Case $i=2$ ($l>\tau$).
+For $l>\tau$, $\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l$. Define
 $u(l):=\frac{\mu_{2}(l)+1}{2}+\kappa>0,\qquad x(l):=u(l)+\frac{l}{2}=\kappa+\frac{1+\sqrt{l^{2}+(2\kappa+d+1)}}{2}.$
-Then $u(l)=x(l)-\tfrac{l}{2}$ and $u(l)+l=x(l)+\tfrac{l}{2}$. As above, up to an additive constant independent of $l$,
-$F_{2}(l)=\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big).$
-Differentiating (same computation as in Case $i=1$ but with general $u$) gives the identity
-$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big)-\frac{1}{2}\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big),$
-i.e.
-$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)-\frac{1}{2}\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big).$
-Now observe that
-$x^{\prime}(l)=\frac{1}{2}\cdot\frac{l}{\sqrt{l^{2}+(2\kappa+d+1)}}\in\Big(0,\frac{1}{2}\Big),$
-so $x^{\prime}(l)-\frac{1}{2}<0$ for all $l>\delta$.
-On the other hand, as in Case $i=1$, concavity of $\psi$ implies $\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\geq 0$.
-Therefore $F_{2}^{\prime}(l)\leq 0$ for all $l>\delta$.
-Hence $F_{2}$ is non-increasing on $(\delta,\infty)$ and
-$\sup_{l>\delta}LA_{\kappa,d}(\mu_{2}(l),l)=\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l),$
-equivalently
-$\sup_{l>\delta}A_{\kappa,\mu_{2}(l),l,a,d}=\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}.$
-Finally, $\mu_{2}(l)\to 1$ as $l\downarrow\delta$ and $\mu_{1}(\delta)=1$. By continuity of $LA_{\kappa,d}(\mu,l)$ on the admissible region,
-$\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l)=LA_{\kappa,d}(1,\delta)=LA_{\kappa,d}(\mu_{1}(\delta),\delta).$
-Therefore the maximum of $A_{\kappa,\mu_{1}(l),l,a,d}$ over $0\leq l\leq\delta$ is attained at $l=\delta$, and the supremum of $A_{\kappa,\mu_{2}(l),l,a,d}$ over $l>\delta$ equals the same value and is approached as $l\downarrow\delta$.
-Consequently, the global maximum over the admissible ranges is achieved at $(\mu,l)=(1,\delta)$, while in the regime $l>\delta$ it is not attained but only approached in the limit $l\downarrow\delta$.
-This proves Theorem 5.
+Then $u(l)=x(l)-\tfrac{l}{2}$, $u(l)+l=x(l)+\tfrac{l}{2}$, and
+$x^{\prime}(l)=\frac{1}{2}\cdot\frac{l}{\sqrt{l^{2}+(2\kappa+d+1)}}\in\left(0,\frac{1}{2}\right),\qquad u^{\prime}(l)=x^{\prime}(l)-\frac{1}{2},\qquad\frac{d}{dl}(u(l)+l)=x^{\prime}(l)+\frac{1}{2}.$
+As above, up to an additive constant independent of $l$,
+$F_{2}(l)=\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\!\left(u(l)+\frac{d}{2}\right)-\log\Gamma\!\left(u(l)+l+\frac{d}{2}\right).$
+A direct differentiation, using $u^{\prime}(l)=x^{\prime}(l)-\tfrac{1}{2}$ and $\tfrac{d}{dl}(u(l)+l)=x^{\prime}(l)+\tfrac{1}{2}$, gives
+$F_{2}^{\prime}(l)=\Bigl(x^{\prime}(l)-\tfrac{1}{2}\Bigr)\bigl[\psi(u)-\psi\!\left(u+\tfrac{d}{2}\right)\bigr]+\Bigl(x^{\prime}(l)+\tfrac{1}{2}\Bigr)\bigl[\psi(u+l)-\psi\!\left(u+l+\tfrac{d}{2}\right)\bigr],$ (34)
+where $u=u(l)$. To determine the sign of $F_{2}^{\prime}(l)$, set
+$h(t):=\psi(t)-\psi\!\left(t+\frac{d}{2}\right),\qquad t>0.$
+Since $\psi$ is concave, $\psi^{\prime}$ is decreasing, hence $h^{\prime}(t)=\psi^{\prime}(t)-\psi^{\prime}(t+\tfrac{d}{2})\geq 0$, so $h$ is non-decreasing.
+Moreover $h(t)<0$ for all $t>0$ since $\psi$ is increasing.
+Because $u<u+l$, it follows that
+$h(u)\leq h(u+l)<0.$
+Since $x^{\prime}(l)-\tfrac{1}{2}<0$, replacing $h(u)$ by the larger value $h(u+l)$ in the first term of (34) gives the upper bound
+$F_{2}^{\prime}(l)\leq\Bigl(x^{\prime}(l)-\tfrac{1}{2}\Bigr)h(u+l)+\Bigl(x^{\prime}(l)+\tfrac{1}{2}\Bigr)h(u+l)=2x^{\prime}(l)\,h(u+l)\leq 0.$
+Therefore $F_{2}$ is non-increasing on $(\tau,\infty)$ and $\sup_{l>\tau}A_{\kappa,\mu_{2}(l),l,a,d}=\lim_{l\downarrow\tau}A_{\kappa,\mu_{2}(l),l,a,d}.$ Since $\mu_{2}(l)\to 1$ as $l\downarrow\tau$ and $\mu_{1}(\tau)=1$, by continuity of $LA_{\kappa,d}(\mu,l)$ on the admissible region,
+$\lim_{l\downarrow\tau}LA_{\kappa,d}(\mu_{2}(l),l)=LA_{\kappa,d}(1,\tau)=LA_{\kappa,d}(\mu_{1}(\tau),\tau).$
+Therefore the maximum of $A_{\kappa,\mu_{1}(l),l,a,d}$ over $0\leq l\leq\tau$ is attained at $l=\tau$, and the supremum of $A_{\kappa,\mu_{2}(l),l,a,d}$ over $l>\tau$ equals the same value and is approached as $l\downarrow\tau$.
+Consequently, the global maximum over the admissible region is achieved at $(\mu,l)=(1,\tau)=(1,d/2+\kappa)$, while in the regime $l>\tau$ it is not attained but only approached in the limit $l\downarrow\tau$.
+Restoring the factor $a^{d}$ yields the statement.
 ∎
@@
 Proof.
-Using (7) and Equation (25) in Emery and Alegría (2022), the Fourier transform of the right-hand side member of (18) is found to be
-$\begin{split}&\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)\Gamma(\delta)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{1}{2})\Gamma(2\delta)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{d+2\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathstrut_{1}F_{2}\Big(\delta;\delta+\frac{1}{2},2\delta;-\frac{uv(za)^{2}}{4}\Big){\rm d}u{\rm d}v\\ &=\frac{\Gamma(\delta)\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{\mu}{2})\Gamma(\delta+\frac{\mu}{2}+l)}\mathstrut_{1}F_{2}\Big(\delta;\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l;-\frac{(za)^{2}}{4}\Big)\\ &=\widehat{\mathcal{GH}}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(z),\quad z\geq 0.\end{split}$
-The claim of the theorem is obtained by inverse Fourier transformation. ∎
+Since all models are compactly supported and hence absolutely integrable, it suffices to verify the identity at the level of spectral densities.
+Taking the Fourier transform of both sides of (18) and using the scaling identity $\widehat{\mathcal{H}}_{\kappa,1,a\sqrt{uv},d}(z)=(a\sqrt{uv})^{d}\,\widehat{\mathcal{H}}_{\kappa,1,1,d}(za\sqrt{uv})$, the right-hand side of (18) has spectral density
+$I(z)=C\,a^{d}\int_{0}^{1}\!\!\int_{0}^{1}u^{A}(1-u)^{B}\,v^{P}(1-v)^{Q}\,(uv)^{d/2}\,\widehat{\mathcal{H}}_{\kappa,1,1,d}(za\sqrt{uv})\,\mathrm{d}u\,\mathrm{d}v,$ (35)
+where, from the exponents in (18),
+$A=\delta-\tfrac{d+1}{2},\quad B=\tfrac{\mu-3}{2},\quad P=\delta-\tfrac{1}{2}+\kappa,\quad Q=\tfrac{\mu-d-3}{2}-\kappa+l.$
+From (7), the spectral density of $\mathcal{H}_{\kappa,1,1,d}$ is
+$\widehat{\mathcal{H}}_{\kappa,1,1,d}(w)=K\,{}_{1}F_{2}\!\left(\delta;\,\delta+\tfrac{1}{2},\,2\delta;\,-\tfrac{w^{2}}{4}\right)=K\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\delta+\frac{1}{2})_{k}\,(2\delta)_{k}}\frac{(-w^{2}/4)^{k}}{k!},$
+where $K>0$ is a constant depending only on $(\kappa,d)$. Substituting $w=za\sqrt{uv}$ and inserting into (35):
+$I(z)=C\,K\,a^{d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\delta+\frac{1}{2})_{k}\,(2\delta)_{k}}\frac{(-z^{2}a^{2}/4)^{k}}{k!}\;\mathcal{B}_{u}(k)\;\mathcal{B}_{v}(k),$
+where the two integrals separate as
+$\mathcal{B}_{u}(k):=\int_{0}^{1}u^{A+d/2+k}(1-u)^{B}\,\mathrm{d}u,\qquad\mathcal{B}_{v}(k):=\int_{0}^{1}v^{P+d/2+k}(1-v)^{Q}\,\mathrm{d}v.$
+By the standard identity $\int_{0}^{1}t^{\alpha+k-1}(1-t)^{\beta-1}\,\mathrm{d}t=B(\alpha,\beta)\,(\alpha)_{k}/(\alpha+\beta)_{k}$, one obtains
+$\mathcal{B}_{u}(k)$ $=B\!\left(A+\tfrac{d}{2}+1,\,B+1\right)\cdot\frac{(A+\frac{d}{2}+1)_{k}}{(A+B+\frac{d}{2}+2)_{k}},$ (36)
+$\mathcal{B}_{v}(k)$ $=B\!\left(P+\tfrac{d}{2}+1,\,Q+1\right)\cdot\frac{(P+\frac{d}{2}+1)_{k}}{(P+Q+\frac{d}{2}+2)_{k}}.$ (37)
+Substituting $A,B,P,Q$ and $\delta=\frac{d+1}{2}+\kappa$:
+$A+\tfrac{d}{2}+1=\delta-\tfrac{d+1}{2}+\tfrac{d}{2}+1=\delta+\tfrac{1}{2},\quad A+B+\tfrac{d}{2}+2=\delta-\tfrac{d+1}{2}+\tfrac{\mu-3}{2}+\tfrac{d}{2}+2=\delta+\tfrac{\mu}{2}=\beta,$
+$P+\tfrac{d}{2}+1=\delta-\tfrac{1}{2}+\kappa+\tfrac{d}{2}+1=2\delta,\quad P+Q+\tfrac{d}{2}+2=\delta-\tfrac{1}{2}+\kappa+\tfrac{\mu-d-3}{2}-\kappa+l+\tfrac{d}{2}+2=\delta+\tfrac{\mu}{2}+l=\gamma.$
+Hence (36) and (37) become
+$\mathcal{B}_{u}(k)=B\!\left(\delta+\tfrac{1}{2},\,\tfrac{\mu-1}{2}\right)\frac{(\delta+\frac{1}{2})_{k}}{(\beta)_{k}},\qquad\mathcal{B}_{v}(k)=B\!\left(2\delta,\,\gamma-2\delta\right)\frac{(2\delta)_{k}}{(\gamma)_{k}}.$
+Inserting back into the power series, the factors $(\delta+\tfrac{1}{2})_{k}$ and $(2\delta)_{k}$ cancel with those in the denominator of the $\mathcal{H}_{\kappa,1}$ expansion, leaving
+$I(z)=C^{\prime}\,a^{d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}\,(\gamma)_{k}}\frac{(-z^{2}a^{2}/4)^{k}}{k!}=C^{\prime}\,\widehat{\mathcal{GH}}_{\delta,\beta,\gamma,a}(z),$
+where $C^{\prime}=C\cdot K\cdot B(\delta+\frac{1}{2},\frac{\mu-1}{2})\cdot B(2\delta,\gamma-2\delta)$. Setting $C$ to be the reciprocal of this product of Beta constants establishes the identity.
+The integrability of the kernel (i.e., that the exponents $B=(\mu-3)/2>-1$ and $Q=(\mu-d-3)/2-\kappa+l>-1$) is guaranteed respectively by $\mu>1$ and $\mu+2l>1+d+2\kappa$, which are precisely the hypotheses of the theorem.
+∎
 A.7 Proof of Theorem 7
+Proof.
 If $d=1$ and $\kappa=k$ is a nonnegative integer, using formula 7.3.1.70 of Prudnikov et al. (1990) and applying the identity
@@
 so that $\tau=2\nu+d$.
-Let $\lambda=\frac{1}{2}+\delta-\beta-\gamma.$ Note that under parameterization (33),
-$\lambda=\frac{1}{2}+\delta-\beta-\gamma=\frac{1}{2}-\delta-\frac{d}{2}-\kappa=-\mu-\frac{d}{2}-\kappa-\frac{1}{2}<0,$
-and also $\beta+\gamma=3\delta+\mu/2+d/2+\kappa-1/2$, so that $\beta+\gamma>(d+1)/2+3\delta$ holds whenever $\mu\geq 1$, which is exactly the validity condition of the $\mathcal{H}$ model.
-Hence Lemma 2 applies.
+Let $\lambda=\frac{1}{2}+\delta-\beta-\gamma.$ Under parameterization (41),
+$\lambda$ $=\frac{1}{2}+\delta-\!\left(\delta+\frac{\mu}{2}\right)-\!\left(\delta+\frac{\mu}{2}+\frac{d}{2}+\kappa\right)$
+$=\frac{1}{2}-\delta-\mu-\frac{d}{2}-\kappa$
+$=-d-2\kappa-\mu\;<\;0,$
+and also
+$\beta+\gamma\;=\;2\delta+\mu+\frac{d}{2}+\kappa\;=\;3\delta+\mu-\frac{1}{2},$
+so that $\beta+\gamma\geq 3\delta+\frac{1}{2}$ holds whenever $\mu\geq 1$, which is exactly the validity condition of the $\mathcal{H}$ model. Hence Lemma 1 applies.
 Using Point 1 of Lemma 1, as $z\to\infty$,
@@
 $\qquad+w_{2}z^{\lambda+2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
-$\qquad\times\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}-1\Bigg|,$
+$\qquad\times\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}-1\Bigg|,$
 where
@@
 $T_{1}(z)=w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2}),\qquad T_{2}(z)=w_{1}z^{2\nu+d-2\delta}-1,$
-$T_{3}(z)=w_{2}z^{\lambda}\{\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}.$
+$T_{3}(z)=w_{2}z^{\lambda}\{\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\}\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}.$
 We now impose the conditions $\nu=\kappa+\frac{1}{2}$ and $w_{1}=1$, and verify that each term yields a finite contribution for $d=1,2,3$.
@@
 which converges for d5<1d-5<-1, i.e. $d<4$. Hence the integral is finite for $d=1,2,3$.
-For $T_{3}$, since $\lambda<0$, the dominant contribution of $T_{3}$ comes from the $z^{2\nu+d}$ factor:
-$T_{3}(z)=w_{2}\,z^{\lambda+2\nu+d}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}+\mathcal{O}(z^{\lambda+2\nu+d-2}).$
-Under $2\nu+d=2\delta$, we have $\lambda+2\nu+d=\lambda+2\delta$. Substituting $\lambda=-\mu-d/2-\kappa-1/2$ and $2\delta=d+1+2\kappa$,
-$\lambda+2\delta=\frac{1}{2}-\mu-\frac{d}{2}+\kappa<0,$
-since $\mu\geq 1$ and $\kappa>-1/2$. Hence $T_{3}(z)=\mathcal{O}(z^{\lambda+2\delta})$ with $\lambda+2\delta<0$, and by the Riemann–Lebesgue argument for oscillatory integrals,
-$\int_{c}^{\infty}z^{d-1}T_{3}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-1+2(\lambda+2\delta)}\,\mathrm{d}z\right),$
-which converges provided $d-1+2(\lambda+2\delta)<-1$, i.e. $d+2\lambda+4\delta<0$. Substituting,
-$d+2\lambda+4\delta=d+2\!\left(-\mu-\tfrac{d}{2}-\kappa-\tfrac{1}{2}\right)+2(d+1+2\kappa)=d+1+2\kappa-2\mu.$
-Since $\mu\geq 1$ and $\kappa>-1/2$, this quantity is at most $d-1+2\kappa<d$.
-For $d=1,2,3$ and $\mu\geq 1$, one verifies directly that $d+1+2\kappa-2\mu\leq d-1+2\kappa<d$; finiteness holds strictly when $\mu>(d+2\kappa)/2$, which is implied by the validity condition $\mu\geq 1$ for $d=1,2,3$ and $\kappa>-1/2$.
+For $T_{3}$, since $\lambda<0$, the dominant contribution comes from the $z^{2\nu+d}$ factor:
+$T_{3}(z)=w_{2}\,z^{\lambda+2\nu+d}\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}+\mathcal{O}(z^{\lambda+2\nu+d-2}).$
+Under $2\nu+d=2\delta$,
+$\lambda+2\nu+d\;=\;\lambda+2\delta\;=\;(-d-2\kappa-\mu)+(d+1+2\kappa)\;=\;1-\mu.$
+Since $\mu\geq 1+\frac{d}{2}$ (as required by Theorems 10), we have $\lambda+2\delta=1-\mu<0$. By the Riemann–Lebesgue argument for oscillatory integrals,
+$\int_{c}^{\infty}z^{d-1}T_{3}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-1+2(1-\mu)}\,\mathrm{d}z\right),$
+which converges provided $d-1+2(1-\mu)<-1$, i.e. $\mu>\frac{d}{2}+1$, which holds under the stated assumption.
 The cross terms $T_{i}T_{j}$ ($i\neq j$) satisfy analogous polynomial decay bounds, and their integrability follows from the same power-counting argument combined with $T_{2}\equiv 0$.
-Combining the above, condition (34) holds whenever $\nu=\kappa+\frac{1}{2}$, $w_{1}=1$, and $d=1,2,3$. Hence
+Combining the above, condition (42) holds whenever $\nu=\kappa+\frac{1}{2}$, $w_{1}=1$, $\mu\geq 1+\frac{d}{2}$, and $d=1,2,3$. Hence
 $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a}).$
@@
 Using parameterization (41), this becomes
-$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}B(\kappa,\mu,d),$
-where $B(\kappa,\mu,d)$ is defined in (20). Finally, the necessary part follows from Zhang (2004), by the same argument as in the proof of Theorem 1. ∎
-A.10 Proof of Lemma 1
+$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}F(\kappa,\mu,d),$
+where $F(\kappa,\mu,d)$ is defined in (20). Finally, the necessary part follows from Zhang (2004), by the same argument as in the proof of Theorem 1. ∎
+A.10 Proof of Theorem 11
+A preliminary lemma is required for the proof of Theorem 11.
+It establishes a uniform stochastic control of the plug-in microergodic estimator $\hat{\theta}_{n}(\mu,a)$ over $a\in I$, which is the key ingredient for showing that replacing a fixed $a$ by its ML estimator does not affect the fixed-domain limit distribution in Theorem 11.
+Lemma 2.
+Assume the setting of Theorem 10 with fixed $\mu>1+\frac{d}{2}$ and $a_{0}\in I$. Let
+$\hat{\theta}_{n}(\mu,a)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}},\qquad\hat{\sigma}_{n}^{2}(\mu,a)=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$
+and $\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu,d)/a_{0}^{2\kappa+1}$. Then, as $n\to\infty$,
+  1. (i)
+    $\sup_{a\in I}\bigl|\hat{\theta}_{n}(\mu,a)-\theta_{0}\bigr|\xrightarrow{p}0$;
+  2. (ii)
+    for any fixed $a^{\star}\in I$, $\sup_{a\in I}\sqrt{n}\,\bigl|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr|=o_{p}(1)$.
 Proof.
-Let $0<a_{1}<a_{2}$ with $a_{1},a_{2}\in I$. For any data vector $\boldsymbol{Z}_{n}$,
-$\frac{\widehat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}-\frac{\widehat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}=\frac{1}{n}\,\boldsymbol{Z}_{n}^{\top}\Big(a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\Big)\boldsymbol{Z}_{n}.$
-This expression is nonnegative for all $\boldsymbol{Z}_{n}$ if and only if
-$a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\succeq 0.$
-Setting $A:=a_{1}^{1+2\kappa}R_{n}(a_{1})$ and $C:=a_{2}^{1+2\kappa}R_{n}(a_{2})$, both symmetric positive definite, the condition above reads $A^{-1}-C^{-1}\succeq 0$.
-Since inversion reverses the Loewner order on the cone of SPD matrices,
-$A^{-1}-C^{-1}\succeq 0\quad\Longleftrightarrow\quad C\succeq A\quad\Longleftrightarrow\quad B:=a_{2}^{1+2\kappa}R_{n}(a_{2})-a_{1}^{1+2\kappa}R_{n}(a_{1})\succeq 0.$
-The $(i,j)$ entry of $B$ is
-$B_{ij}=a_{2}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{2},d}(\|s_{i}-s_{j}\|)-a_{1}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{1},d}(\|s_{i}-s_{j}\|).$
-By Bochner’s theorem, $B\succeq 0$ for every finite collection of locations $\{s_{1},\ldots,s_{n}\}\subset D$ if and only if the associated spectral density difference is nonnegative for all $z\geq 0$:
-$f(z):=a_{2}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0.$
-Using the scaling identity $\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)=t^{d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(tz)$ induced by (7), the condition $f(z)\geq 0$ for all $z\geq 0$ is equivalent to monotonicity of
-$g(t):=t^{2\kappa+1+d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t),\qquad t>0.$
-Using (7) with $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+l$, we have, up to a positive constant,
-$g(t)=t^{2\kappa+1+d}\,{}_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right).$
-Substituting the power series representation
-$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right)=\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!},$
-and differentiating term by term (justified by uniform convergence on compact subsets of $(0,\infty)$), one obtains
-$g^{\prime}(t)$ $=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}+t^{2\kappa+1+d}\sum_{k=1}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-1/4)^{k}\,2k\,t^{2k-1}}{k!}$
-$=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k+1}}{(\delta)_{1}(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}\cdot\frac{(\delta)_{1}}{\delta}\;+\;\cdots$
-Using the identity $(\delta)_{k+1}=\delta\,(\delta+1)_{k}$ and collecting terms, both sums combine into
-$g^{\prime}(t)=C\,t^{2\kappa+d}\,{}_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right),\qquad t>0,$
-where $C=2(\kappa+1+d/2)\cdot\delta/(\beta\gamma)\cdot\beta\gamma>0$. Hence $g$ is nondecreasing on $(0,\infty)$ if and only if
-$_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right)\geq 0\qquad\forall\,t>0.$
-Fix $l=\frac{d}{2}+\kappa$.
-By (7), condition ()(*) is equivalent to requiring that $\mathcal{GH}_{\delta_{1},\beta_{1},\gamma_{1},a}\in\Phi_{d}$ with $\delta_{1}:=\delta+1$, $\beta_{1}:=\beta$, $\gamma_{1}:=\gamma$.
-Translating to the GW-type parameterization, the shifted parameters correspond to
-$\kappa_{1}:=\kappa+1\quad\bigl(\text{since }\delta_{1}=\tfrac{d+1}{2}+\kappa_{1}\bigr),\qquad\mu_{1}:=\mu-2\quad\bigl(\text{since }\beta_{1}=\delta_{1}+\tfrac{\mu_{1}}{2}\bigr),$
-with $l$ unchanged.
-Since $l=\frac{d}{2}+\kappa<\frac{d}{2}+\kappa_{1}$, the shifted parameters fall in Regime 1 of Theorem 2, which gives the necessary and sufficient condition $\mu_{1}\geq\delta_{1}-l+\frac{1}{2}$.
-Substituting $\delta_{1}=\frac{d+1}{2}+\kappa+1$ and $l=\frac{d}{2}+\kappa$:
-$\delta_{1}-l+\tfrac{1}{2}=\Bigl(\tfrac{d+1}{2}+\kappa+1\Bigr)-\Bigl(\tfrac{d}{2}+\kappa\Bigr)+\tfrac{1}{2}=2,$
-so $\mu_{1}\geq 2$, i.e. $\mu\geq 4$. This completes the proof. ∎
+Fix $\mu>1+\frac{d}{2}$ and $I=[a_{L},a_{U}]$. For $a\in I$ set
+$G_{n}(a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\,R_{n}(\mu,a),\qquad\hat{\theta}_{n}(\mu,a)=\frac{1}{n}\,\mathbf{Z}_{n}^{\top}G_{n}(a)^{-1}\mathbf{Z}_{n}.$
+Let $G_{0n}:=G_{n}(a_{0})$ and define the vector
+$\mathbf{Y}_{n}:=\theta_{0}^{-1/2}\,G_{0n}^{-1/2}\mathbf{Z}_{n}\sim\mathcal{N}(0,I_{n}),\qquad\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu,d)/a_{0}^{2\kappa+1}.$
+Then
+$\frac{\hat{\theta}_{n}(\mu,a)}{\theta_{0}}=\frac{1}{n}\,\mathbf{Y}_{n}^{\top}M_{n}(a)\mathbf{Y}_{n},\qquad M_{n}(a):=G_{0n}^{1/2}G_{n}(a)^{-1}G_{0n}^{1/2}.$ (43)
+Fix $a^{\star}\in I$ and set $\Delta_{n}(a):=M_{n}(a)-M_{n}(a^{\star})$. The decomposition
+$\frac{\sqrt{n}\,\bigl(\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr)}{\theta_{0}}=\underbrace{\frac{1}{\sqrt{n}}\operatorname{tr}(\Delta_{n}(a))}_{\mathrm{(I)}}+\underbrace{\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\operatorname{tr}(\Delta_{n}(a))\bigr)}_{\mathrm{(II)}}$ (44)
+will be used to prove (ii), and then (i) follows from (ii) by taking $a^{\star}=a_{0}$.
+Hereafter, we use $\|A\|_{\mathrm{HS}}^{2}:=\mathrm{tr}(A^{*}A)$ for Hilbert–Schmidt operators $A$, where $A^{*}$ denotes the adjoint operator.
+For matrices $B$, we use $\|B\|_{F}^{2}:=\mathrm{tr}(B^{\top}B)$.
+Let $g_{a}$ denote the isotropic spectral density of the stationary covariance function $G(\cdot\,;a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\mathcal{H}_{\kappa,\mu,a,d}(\cdot)$.
+By Lemma 1 specialised to the $\mathcal{H}$ model, the high-frequency expansion of $g_{a}$ has the form
+$g_{a}(z)=C_{\kappa,\mu,d}\,z^{-2\delta}\bigl(1+\mathcal{O}(z^{-2})\bigr)+\mathcal{O}(z^{1-\mu-2\delta}),\qquad z\to\infty,$
+with $\delta=\frac{d+1}{2}+\kappa$ and with $\mathcal{O}$-constants uniform over $a\in I$ (the leading term is independent of $a$ because of the prefactor $a^{2\kappa+1}$). Hence, for any fixed $c>0$,
+$\sup_{a\in I}\int_{c}^{\infty}z^{d-1}\Bigl(\frac{g_{a}(z)-g_{a_{0}}(z)}{g_{a_{0}}(z)}\Bigr)^{2}\,dz<\infty,$ (45)
+since $\mu>1+d/2$ and $d\in\{1,2,3\}$ make the tail square-integrable.
+By the fixed-domain equivalence criterion of Stein (2004) and its connection to covariance-ratio operators (see Skorokhod and Yadrenko (1973, Assertion B)), (45) implies that the operator
+$\mathcal{M}(a):=\mathcal{G}(a_{0})^{1/2}\mathcal{G}(a)^{-1}\mathcal{G}(a_{0})^{1/2}-\mathcal{I}\quad\text{on }L^{2}(D)$
+is Hilbert–Schmidt, with $\sup_{a\in I}\|\mathcal{M}(a)\|_{\mathrm{HS}}<\infty$.
+Let $\Pi_{n}$ denote the orthogonal projector onto the finite-dimensional subspace generated by the observations $\{Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n})\}$ (identified with its image in $L^{2}(D)$ through the covariance-operator embedding).
+Then $M_{n}(a)-I_{n}$ represents the restriction of $\Pi_{n}\mathcal{M}(a)\Pi_{n}$ to that subspace, hence
+$\sup_{a\in I}\|M_{n}(a)-I_{n}\|_{F}^{2}=\sup_{a\in I}\|\Pi_{n}\mathcal{M}(a)\Pi_{n}\|_{\mathrm{HS}}^{2}\leq\sup_{a\in I}\|\mathcal{M}(a)\|_{\mathrm{HS}}^{2}=:C_{\mathrm{HS}}<\infty,\qquad\forall n.$ (46)
+In particular,
+$\sup_{a\in I}\|M_{n}(a)\|$ $\leq 1+\sup_{a\in I}\|M_{n}(a)-I_{n}\|\leq 1+\sup_{a\in I}\|M_{n}(a)-I_{n}\|_{F}$ (47)
+$\leq 1+\sqrt{C_{\mathrm{HS}}},\qquad\sup_{a\in I}\|\Delta_{n}(a)\|_{F}^{2}\leq 4C_{\mathrm{HS}},\qquad\forall n.$
+Let $p:=\mu+\frac{d-1}{2}+2\kappa$ and $\alpha:=\min(1,p)$.
+Since $\mu>1+\frac{d}{2}$ and $\kappa>-1/2$, one has $\alpha>1/2$.
+From the explicit form of $\mathcal{H}_{\kappa,\mu,a,d}$ and its compact support in $[0,a]$, the map $a\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is uniformly $\alpha$-Hölder on $I$, namely
+$\sup_{r\geq 0}\bigl|\mathcal{H}_{\kappa,\mu,a,d}(r)-\mathcal{H}_{\kappa,\mu,a^{\prime},d}(r)\bigr|\leq C_{\alpha}|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I,$ (48)
+where the exponent $\alpha=\min(1,p)$ reflects the boundary behaviour at $r=a$.
+Moreover, using (48), continuity in $(a,z)$ on compact sets, positivity of $g_{a_{0}}$, and the common high-frequency expansion above, one obtains that for any fixed $c>0$,
+$\int_{c}^{\infty}z^{d-1}\Bigl(\frac{g_{a}(z)-g_{a^{\prime}}(z)}{g_{a_{0}}(z)}\Bigr)^{2}dz\leq C|a-a^{\prime}|^{2\alpha},\qquad a,a^{\prime}\in I,$
+with $C$ independent of $(a,a^{\prime})$. Consequently, by the same covariance-ratio/operator correspondence (Stein, 2004, Appendix A) (see also Skorokhod and Yadrenko (1973, Assertion B)),
+$\|\mathcal{M}(a)-\mathcal{M}(a^{\prime})\|_{\mathrm{HS}}\leq C|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I.$
+Passing to the discretised matrices and using again that orthogonal projections do not increase HS norms yields the Frobenius increment bound
+$\|\Delta_{n}(a)-\Delta_{n}(a^{\prime})\|_{F}\leq C|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I,\ \forall n.$ (49)
+By Hanson–Wright (Hanson and Wright, 1971; Rudelson and Vershynin, 2013), for any symmetric $A$ and $\mathbf{Y}\sim\mathcal{N}(0,I_{n})$,
+$P\!\Bigl(\bigl|\mathbf{Y}^{\top}A\mathbf{Y}-\operatorname{tr}(A)\bigr|>t\Bigr)\leq 2\exp\!\Bigl(-c\min\!\Bigl\{\frac{t^{2}}{\|A\|_{F}^{2}},\frac{t}{\|A\|}\Bigr\}\Bigr).$
+With $A=\Delta_{n}(a)$ and $t=\varepsilon\sqrt{n}$, using (47) and $\|A\|\leq\|A\|_{F}$,
+$P\!\Bigl(\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|>\varepsilon\Bigr)\leq 2e^{-c^{\prime}n},$
+uniformly in $a\in I$. Let $\delta_{n}:=n^{-4}$ and let $\mathcal{G}_{n}\subset I$ be a regular grid of mesh $\delta_{n}$ (so $|\mathcal{G}_{n}|=\mathcal{O}(n^{4})$). A union bound gives
+$\sup_{a\in\mathcal{G}_{n}}\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|=o_{p}(1).$ (50)
+For arbitrary $a\in I$, choose $a^{(n)}\in\mathcal{G}_{n}$ with $|a-a^{(n)}|\leq\delta_{n}$. Then by (49) and $\|\mathbf{Y}_{n}\|^{2}=O_{p}(n)$,
+$\frac{1}{\sqrt{n}}\bigl|\mathbf{Y}_{n}^{\top}(\Delta_{n}(a)-\Delta_{n}(a^{(n)}))\mathbf{Y}_{n}\bigr|$ $\leq\frac{\|\Delta_{n}(a)-\Delta_{n}(a^{(n)})\|_{F}}{\sqrt{n}}\,\|\mathbf{Y}_{n}\|^{2}$
+$\leq C\delta_{n}^{\alpha}\,O_{p}(\sqrt{n})=O_{p}(n^{1/2-4\alpha})=o_{p}(1).$
+since $\alpha>1/2$. Combining with (50) yields
+$\sup_{a\in I}\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|=o_{p}(1).$ (51)
+Now, define
+$r_{n}(a):=\frac{1}{n}\mathrm{tr}(M_{n}(a))-1=\frac{E[\hat{\theta}_{n}(\mu,a)]}{\theta_{0}}-1.$
+Then $\frac{1}{\sqrt{n}}\mathrm{tr}(\Delta_{n}(a))=\sqrt{n}\bigl(r_{n}(a)-r_{n}(a^{\star})\bigr)$. First, by (49) and Cauchy–Schwarz,
+$\sqrt{n}\,|r_{n}(a)-r_{n}(a^{\prime})|=\frac{|\mathrm{tr}(M_{n}(a)-M_{n}(a^{\prime}))|}{\sqrt{n}}\leq\|M_{n}(a)-M_{n}(a^{\prime})\|_{F}\leq C|a-a^{\prime}|^{\alpha},$
+uniformly in $n$, so $\{\sqrt{n}\,r_{n}(\cdot)\}_{n\geq 1}$ is equicontinuous on $I$.
+Second, for each fixed $a$, Theorem 10(ii) gives $\sqrt{n}(\hat{\theta}_{n}(\mu,a)-\theta_{0})\Rightarrow\mathcal{N}(0,2\theta_{0}^{2})$.
+Moreover, for Gaussian quadratic forms,
+$\rm Var\!\Big(\mathbf{Y}_{n}^{\top}M_{n}(a)\mathbf{Y}_{n}-\operatorname{tr}(M_{n}(a))\Big)=2\,\operatorname{tr}(M_{n}(a)^{2})\leq 2\,\|M_{n}(a)\|_{F}^{2},$
+so (46) yields a uniform second-moment bound and hence uniform integrability. Therefore
+[…]
+By equicontinuity and compactness of $I$ (finite-cover argument),
+$\sup_{a\in I}\frac{1}{\sqrt{n}}|\mathrm{tr}(\Delta_{n}(a))|=\sup_{a\in I}\sqrt{n}\,|r_{n}(a)-r_{n}(a^{\star})|\leq 2\sup_{a\in I}\sqrt{n}\,|r_{n}(a)|\longrightarrow 0.$ (52)
+Finally, combining (44), (51), and (52) gives
+$\sup_{a\in I}\sqrt{n}\,\bigl|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr|=o_{p}(1),$
+which proves (ii).
+To show (i), take $a^{\star}=a_{0}$ in (ii) to get
+$\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\theta_{0}|\leq\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a_{0})|+|\hat{\theta}_{n}(\mu,a_{0})-\theta_{0}|.$
+The first term is $o_{p}(1)$ by (ii), and the second converges to 0 almost surely by Theorem 10(i). Hence $\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\theta_{0}|\to 0$ in probability, proving (i). ∎
+A.11 Proof of Theorem 12
+The proof of Theorem 12 requires a uniform stochastic control of the plug-in microergodic estimator over the compact parameter set $J\times I$, since the ML estimators $(\hat{\mu}_{n},\hat{a}_{n})$ are random.
+The next lemma extends Lemma 2 to the case where $\mu$ is unknown and varies in $J$.
+Lemma 3.
+Assume the setting of Theorem 10, with $(\sigma_{0}^{2},\mu_{0},a_{0})\in(0,\infty)\times J\times I$ and $\mu_{0}>1+\frac{d}{2}$, where $J=[\mu_{L},\mu_{U}]$ and $I=[a_{L},a_{U}]$ with $\mu_{L}>1+\frac{d}{2}$.
+Let
+$\hat{\theta}_{n}(\mu,a)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}},\qquad\hat{\sigma}_{n}^{2}(\mu,a)=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$
+and $\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu_{0},d)/a_{0}^{2\kappa+1}$. Then, as $n\to\infty$,
+  1. (i)
+    $\sup_{(\mu,a)\in J\times I}\big|\hat{\theta}_{n}(\mu,a)-\theta_{0}\big|\xrightarrow{p}0;$
+  2. (ii)
+    for any fixed $(\mu^{\star},a^{\star})\in J\times I$,
+    $\sup_{(\mu,a)\in J\times I}\sqrt{n}\,\big|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu^{\star},a^{\star})\big|=o_{p}(1).$
+Proof.
+The argument follows closely the proof of Lemma 2.
+We keep the same notation and decomposition, replacing $a\in I$ by $(\mu,a)\in J\times I$, and we record only the modifications needed to obtain uniformity over the two-parameter compact set.
+All bounds that in Lemma 2 were uniform over $a\in I$ remain uniform over $(\mu,a)\in J\times I$, because $J\times I$ is compact and $\mu_{L}>1+\frac{d}{2}$.
+In particular, the high-frequency expansion of the normalised spectral density and Stein’s integrability condition hold uniformly on $J\times I$, yielding uniform Hilbert–Schmidt bounds for the covariance-ratio operators and hence
+$\sup_{(\mu,a)\in J\times I}\|M_{n}(\mu,a)-I_{n}\|_{F}^{2}\leq C,\qquad\sup_{(\mu,a)\in J\times I}\|M_{n}(\mu,a)\|\leq 1+\sqrt{C},$
+for some finite constant $C$ and all $n$, where $M_{n}(\mu,a)$ is defined as in Lemma 2 with $G_{n}(\mu,a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\,R_{n}(\mu,a)$.
+Moreover, on $J\times I$ the map $\mu\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is $C^{1}$ and thus uniformly Lipschitz in $\mu$, while the map $a\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is uniformly $\alpha$–Hölder in $a$ with $\alpha>1/2$, implying the two-parameter increment bound
+$\|\Delta_{n}(\mu,a)-\Delta_{n}(\mu^{\prime},a^{\prime})\|_{F}\leq C\bigl(|\mu-\mu^{\prime}|+|a-a^{\prime}|^{\alpha}\bigr),\qquad(\mu,a),(\mu^{\prime},a^{\prime})\in J\times I,$
+where $\Delta_{n}(\mu,a)$ denotes the corresponding increment matrix.
+Finally, the Hanson–Wright and grid argument is applied on a two-dimensional grid of mesh $\delta_{n}=n^{-4}$ in $J\times I$ (of cardinality $\mathcal{O}(n^{8})$), and the remainder of the proof proceeds exactly as in Lemma 2, yielding (ii).
+Statement (i) then follows by taking $(\mu^{\star},a^{\star})=(\mu_{0},a_{0})$ and using Theorem 10(i).
+∎
+Proof of Theorem 12 .
+By definition, $(\hat{\mu}_{n},\hat{a}_{n})\in J\times I$ for all $n$.
+Proof of (i). Since $(\hat{\mu}_{n},\hat{a}_{n})\in J\times I$,
+$\big|\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big|\leq\sup_{(\mu,a)\in J\times I}\big|\hat{\theta}_{n}(\mu,a)-\theta_{0}\big|.$
+By Lemma 3(i), the right-hand side converges to 0 in probability, hence $\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})\xrightarrow{p}\theta_{0}$.
+Proof of (ii). Fix the nonrandom point $(\mu^{\star},a^{\star})=(\mu_{0},a_{0})\in J\times I$. Then
+$\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big)$ $=\sqrt{n}\big(\hat{\theta}_{n}(\mu_{0},a_{0})-\theta_{0}\big)$
+$\quad+\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\hat{\theta}_{n}(\mu_{0},a_{0})\big).$
+By Lemma 3(ii), the second term is $o_{p}(1)$. By Theorem 10(ii) (applied with $\mu=\mu_{0}$ and $a=a_{0}$ fixed),
+$\sqrt{n}\big(\hat{\theta}_{n}(\mu_{0},a_{0})-\theta_{0}\big)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2}).$
+Therefore, Slutsky’s theorem yields
+$\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2}),$
+which proves (ii). ∎
 Appendix B Closed-form expressions of the hypergeometric model